### v1

# Lightweight Transformer via Unrolling of Mixed Graph Algorithms for Traffic Forecast

Abstract
To forecast traffic with both spatial and temporal dimensions, we unroll a mixed-graph-based optimization algorithm into a lightweight and interpretable transformer-like neural net. Specifically, we construct two graphs: an undirected graph ${\mathcal{G}}^{u}$ capturing spatial correlations across geography, and a directed graph ${\mathcal{G}}^{d}$ capturing sequential relationships over time. We formulate a prediction problem for the future samples of signal ${\mathbf{x}}$, assuming it is “smooth” with respect to both ${\mathcal{G}}^{u}$ and ${\mathcal{G}}^{d}$, where we design new $\ell_{2}$ and $\ell_{1}$-norm variational terms to quantify and promote signal smoothness (low-frequency reconstruction) on a directed graph. We construct an iterative algorithm based on alternating direction method of multipliers (ADMM), and unroll it into a feed-forward network for data-driven parameter learning. We insert graph learning modules for ${\mathcal{G}}^{u}$ and ${\mathcal{G}}^{d}$, which are akin to the self-attention mechanism in classical transformers. Experiments show that our unrolled networks achieve competitive traffic forecast performance as state-of-the-art prediction schemes, while reducing parameter counts drastically. Our code is available in https://github.com/SingularityUndefined/Unrolling-GSP-STForecast.
1 Introduction
We study the problem of forecasting city traffic with spatial and temporal dimensions, e.g., predicting car density at a set of intersections 5 minutes into the future given measurements in the past. Optimizing prediction here means exploiting both spatial correlations across geography and sequential relationships in time of how the past measurements inform future observations. When modeling the pairwise relationships as a graph, the problem is also called time-vertex signal processing (TVSP) [1].
Among model-based methods, one approach to TVSP is to filter the signal along the spatial and temporal dimensions independently [2]e.g., employ graph spectral filters in the graph signal processing (GSP) literature [3, 4] along the irregular spatial dimension, and traditional Fourier filters along the regular time dimension. Another approach is to model the spatial correlations across time as a sequence of slowly time-varying graphs, and then process each spatial signal at time $t$ separately, where the graph changes over time are constrained by the Frobenius norm [5], $\ell_{1}$-norm [6], or low-rankness [7] of consecutive adjacency matrices. However, neither approach exploits spatial and temporal relations simultaneously for optimal performance.
Leveraging the prowess of modern deep learning (DL) models like transformers [8], recent data-driven efforts address TVSP by building elaborate transformer architectures to capture spatial and temporal relations for prediction [9, 10, 11]. However, these schemes are inherently uninterpretable “black boxes” requiring large training data to tune millions of parameters. They are also fragile to covariate shift [12]: statistical mismatch between training and testing data.
Graph attention networks (GAT) [13] and graph transformers [14] are adaptations of the basic self-attention mechanism in transformers for graph-structured data (e.g., computing output embeddings using only input embeddings in local graph neighborhoods). However, they are still intrinsically uninterpretable and maintain large parameter sizes.
In contrast, algorithm unrolling [15] offers an alternative hybrid model-based / data-driven paradigm; by “unrolling” iterations of a model-based algorithm minimizing a well-defined optimization objective into neural layers stacked together to form a feed-forward network for data-driven parameter learning via back-propagation, the resulting network can be both mathematically interpretable and competitive in performance to state-of-the-art (SOTA) DL models. As examples, Yang et al. [16], Yu et al. [17] design optimization algorithms that unroll into transformer-like neural nets. Inspired by Yu et al. [17] and extending from Thuc et al. [18] [^1] While Thuc et al. [18] focuses on unrolling of undirected-graph-based algorithms into transformer-like neural nets for image interpolation, we extend the study to mixed graphs with both undirected and directed edges to capture spatial and temporal pairwise relations, respectively, for traffic forecast., we unroll a dual-graph-based optimization algorithm into a lightweight interpretable transformer-like neural net for the traffic forecast problem.
Specifically, we propose a novel mixed-graph approach and learn undirected graph ${\mathcal{G}}^{u}$ and directed graph ${\mathcal{G}}^{d}$ from data, to capture spatial correlations across geography and sequential relationships over time, respectively. We show that our graph learning modules are akin to the basic self-attention mechanism [19], and thus our unrolled graph-based neural nets are transformers. Given the two learned graphs, we define a prediction objective for future samples in signal ${\mathbf{x}}$, assuming ${\mathbf{x}}$ is “smooth” with respect to (w.r.t.) both ${\mathcal{G}}^{u}$ and ${\mathcal{G}}^{d}$ in variational terms: graph Laplacian regularizer (GLR) [20] for undirected ${\mathcal{G}}^{u}$, and newly designed directed graph Laplacian regularizer (DGLR) and directed graph total variation (DGTV) for directed ${\mathcal{G}}^{d}$. We construct a corresponding iterative optimization algorithm based on alternating direction method of multipliers (ADMM) [21] that unrolls into neural layers for parameter learning. For the first time, our proposed DGLR and DGTV variational terms quantify and promote signal smoothness on a directed graph, with intuitive low-pass filter interpretations. Experiments show that our unrolled networks achieve competitive traffic forecast performance as SOTA prediction schemes, while employing drastically fewer parameters (6.4% of transformer-based PDFormer [22]).
2 Preliminaries
Undirected Graph Definitions:
Denote by ${\mathcal{G}}^{u}({\mathcal{V}},{\mathcal{E}}^{u},{\mathbf{W}}^{u})$ an undirected graph with node set ${\mathcal{V}}=\{1,\ldots,N\}$ and undirected edge set ${\mathcal{E}}^{u}$, where $(i,j)\in{\mathcal{E}}^{u}$ implies that an undirected edge exists connecting nodes $i$ and $j$ with weight $w_{i,j}^{u}=W_{i,j}^{u}$, and $(i,j)\notin{\mathcal{E}}$ implies $W_{i,j}^{u}=0$. ${\mathbf{W}}^{u}\in\mathbb{R}^{N\times N}$ is the symmetric adjacency matrix for ${\mathcal{G}}^{u}$. Denote by ${\mathbf{D}}^{u}\in\mathbb{R}^{N\times N}$ a diagonal degree matrix, where $D_{i,i}^{u}=\sum_{j}W_{j,i}^{u}$. The symmetric graph Laplacian matrix is ${\mathbf{L}}^{u}\triangleq{\mathbf{D}}^{u}-{\mathbf{W}}^{u}$ [3]. ${\mathbf{L}}^{u}$ is positive semi-definite (PSD) if all edge weights are non-negative $w^{u}_{i,j}\geq 0,\forall i,j$ [4]. The symmetric normalized graph Laplacian matrix is ${\mathbf{L}}^{u}_{n}\triangleq{\mathbf{I}}-({\mathbf{D}}^{u})^{-1/2}{\mathbf{W}}^{u}({\mathbf{D}}^{u})^{-1/2}$, while the asymmetric random-walk graph Laplacian matrix is ${\mathbf{L}}^{u}_{r}\triangleq{\mathbf{I}}-({\mathbf{D}}^{u})^{-1}{\mathbf{W}}^{u}$.
One can define a graph spectrum by eigen-decomposing the real and symmetric graph Laplacian ${\mathbf{L}}^{u}$ (or normalized graph Laplacian ${\mathbf{L}}_{n}^{u}$), where the $k$-th eigen-pair $(\lambda_{k},{\mathbf{v}}_{k})$ is interpreted as the $k$-th graph frequency and Fourier mode, respectively [3]. Given $\{{\mathbf{v}}_{k}\}$ are orthonormal vectors, one can decompose a signal ${\mathbf{x}}$ into its graph frequency components as ${\boldsymbol{\alpha}}={\mathbf{V}}^{\top}{\mathbf{x}}$, where ${\mathbf{L}}^{u}={\mathbf{V}}\mathrm{diag}(\{\lambda_{k}\}){\mathbf{V}}^{\top}$, and ${\mathbf{V}}^{\top}$ is the graph Fourier transform (GFT).
The $\ell_{2}$-norm graph Laplacian regularizer (GLR) [20], ${\mathbf{x}}^{\top}{\mathbf{L}}^{u}{\mathbf{x}}=\sum_{(i,j)\in{\mathcal{E}}^{u}}w^{u}_{i,j}(x_{i}-x_{j})^{2}$, is used to regularize an ill-posed signal restoration problem, like denoising, to bias low-frequency signal reconstruction (i.e., signals consistent with similarity graph ${\mathcal{G}}^{u}$) given observation ${\mathbf{y}}$:
${\mathbf{x}}^{*}$ $=\arg\min_{\mathbf{x}}\|{\mathbf{y}}-{\mathbf{x}}\|^{2}_{2}+\mu\,{\mathbf{x}}^{\top}{\mathbf{L}}^{u}{\mathbf{x}}$
$=({\mathbf{I}}+\mu{\mathbf{L}}^{u})^{-1}{\mathbf{y}}={\mathbf{V}}\mathrm{diag}\left(\frac{1}{1+\mu\lambda_{1}},\ldots,\frac{1}{1+\mu\lambda_{N}}\right){\mathbf{V}}^{\top}{\mathbf{y}}.$ (1)
$\mu$ is a bias-variance tradeoff parameter [23], and the low-pass filter response is $f(\lambda)=(1+\mu\lambda)^{-1}$.
Directed Graph Definitions:
We define analogous notations for a directed graph, denoted by ${\mathcal{G}}^{d}({\mathcal{V}},{\mathcal{E}}^{d},{\mathbf{W}}^{d})$. $[i,j]\in{\mathcal{E}}^{d}$ implies a directed edge exists from node $i$ to $j$ with weight $w_{j,i}^{d}=W_{j,i}^{d}$. ${\mathbf{W}}^{d}\in\mathbb{R}^{N\times N}$ is the asymmetric adjacency matrix. Denote by ${\mathbf{D}}^{d}\in\mathbb{R}^{N\times N}$ a diagonal in-degree matrix, where $D_{i,i}^{d}=\sum_{j}W_{i,j}^{d}$. The directed graph Laplacian matrix is defined as ${\mathbf{L}}^{d}\triangleq{\mathbf{D}}^{d}-{\mathbf{W}}^{d}$. The directed random-walk graph Laplacian matrix is ${\mathbf{L}}^{d}_{r}\triangleq{\mathbf{I}}-({\mathbf{D}}^{d})^{-1}{\mathbf{W}}^{d}$.
Similar frequency notion cannot be simply defined for directed graphs via eigen-decomposition of directed graph Laplacian ${\mathbf{L}}^{d}$, because ${\mathbf{L}}^{d}$ is asymmetric, and thus the Spectral Theorem [24] does not apply. We circumvent this problem by designing new variational terms in Section 3.2.
3 Optimization Formulation & Algorithm
3.1 Mixed Graph for Spatial/Temporal Data
Figure 1: Example of a mixed graph with undirected edges (blue) connecting nodes of the same time instants, and directed edges (red) connecting nodes at $t$ to nodes in window $\{t+1,t+2\}$.
We describe the construction of a mixed graph for spatial/temporal data. A measurement station $i$ observes samples $x_{i}^{t-T},\ldots,x_{i}^{t}$ at past and present instants $t-T,\ldots,t$ and samples $x_{i}^{t+1},\ldots,x_{i}^{t+S}$ at future instants $t+1,\ldots,t+S$. Denote by ${\mathbf{x}}=[{\mathbf{x}}^{t-T};\ldots;{\mathbf{x}}^{t+S}]\in\mathbb{R}^{N(T+S+1)}$ the target graph signal—a concatenation of samples from all $N$ nodes across all $T+S+1$ instants. A product graph ${\mathcal{G}}$ of $N\times(T+S+1)$ nodes represents samples at $T+S+1$ instants. ${\mathcal{G}}$ is a mixture of: i) undirected graph ${\mathcal{G}}^{u}$ connecting spatially node pairs of the same instants, and ii) directed graph ${\mathcal{G}}^{d}$ connecting temporally each node $i$ at instant $\tau$ to the same node at instants $\tau+1,\ldots,\tau+W$, where $W$ denotes the pre-defined time window. See Fig. 1 for an illustration of a mixed graph with undirected spatial edges (blue) and directed temporal edges (red) for $W=2$.
3.2 Variational Terms for Directed Graphs
Our constructed ${\mathcal{G}}^{d}$ is a directed acyclic graph (DAG); a directed edge always stems from a node at instant $\tau$ to a node at a future instant $\tau+s$, and thus no cycles exist. We first define a symmetric variational term for ${\mathcal{G}}^{d}$ called directed graph Laplacian regularizer (DGLR) to quantify variation of a signal ${\mathbf{x}}$ on ${\mathcal{G}}^{d}$. Denote by ${\mathcal{S}}\subset{\mathcal{N}}$ the set of source nodes with zero in-degrees, and $\bar{{\mathcal{S}}}\triangleq{\mathcal{N}}\setminus{\mathcal{S}}$ its complement. First, we add a self-loop of weight $1$ to each node in ${\mathcal{S}}$; this ensures that $({\mathbf{D}}^{d})^{-1}$ is well-defined. Next, we define the row-stochastic random-walk adjacency matrix ${\mathbf{W}}_{r}^{d}\triangleq({\mathbf{D}}^{d})^{-1}{\mathbf{W}}^{d}$.
3.2.1 $\ell_{2}$-norm Variational Term
Interpreting ${\mathbf{W}}_{r}^{d}$ as a graph shift operator (GSO) [25] on ${\mathcal{G}}^{d}$, we first define an $\ell_{2}$-norm variational term [^2] A similar directed graph variational term was defined in Li et al. [26] towards an objective for directed graph sampling. We focus instead on directed graph signal restoration. for ${\mathcal{G}}^{d}$ as the squared difference between ${\mathbf{x}}$ and its graph-shifted version ${\mathbf{W}}_{r}^{d}{\mathbf{x}}$:
$\|{\mathbf{x}}-{\mathbf{W}}_{r}^{d}{\mathbf{x}}\|^{2}_{2}=\|({\mathbf{I}}-{\mathbf{W}}_{r}^{d}){\mathbf{x}}\|^{2}_{2}={\mathbf{x}}^{\top}\underbrace{({\mathbf{L}}_{r}^{d})^{\top}{\mathbf{L}}_{r}^{d}}_{{\mathcal{L}}_{r}^{d}}{\mathbf{x}}$ (2)
where symmetrized directed graph Laplacian matrix ${\mathcal{L}}_{r}^{d}\triangleq({\mathbf{L}}^{d}_{r})^{\top}{\mathbf{L}}^{d}_{r}$ is symmetric and PSD. We call ${\mathbf{x}}^{\top}{\mathcal{L}}^{d}_{r}{\mathbf{x}}$ the DGLR, which computes the sum of squared differences between each child $j\in\bar{{\mathcal{S}}}$ and its parents $i$ for $[i,j]\in{\mathcal{E}}^{d}$. Note that DGLR computes to zero for the constant vector ${\mathbf{1}}$:
${\mathbf{1}}^{\top}{\mathcal{L}}^{d}_{r}{\mathbf{1}}$ $={\mathbf{1}}^{\top}({\mathbf{L}}^{d}_{r})^{\top}{\mathbf{L}}^{d}_{r}{\mathbf{1}}={\mathbf{1}}^{\top}({\mathbf{L}}^{d}_{r})^{\top}({\mathbf{I}}-{\mathbf{W}}^{d}_{r}){\mathbf{1}}\stackrel{{\scriptstyle(a)}}{{=}}{\mathbf{0}}$
where $(a)$ follows from ${\mathbf{W}}^{d}_{r}$ being row-stochastic. This makes sense, as constant ${\mathbf{1}}$ is the smoothest signal and has no variation across any graph kernels.
We interpret eigen-pairs $\{(\xi_{k},{\mathbf{u}}_{k})\}$ of ${\mathcal{L}}^{d}_{r}$ as graph frequencies and graph Fourier modes for directed graph ${\mathcal{G}}^{d}$, respectively, similar to eigen-pairs $\{(\lambda_{k},{\mathbf{v}}_{k})\}$ of graph Laplacian ${\mathbf{L}}^{u}$ for an undirected graph. Importantly, we prove that our frequency definition using DGLR defaults to pure sinusoids as frequencies in the unweighted directed line graph case. Thus, our algorithm minimizing DGLR includes TVSP schemes like Ramakrishna et al. [2] that analyze signals along the time dimension using classical Fourier filters as special cases.
Theorem 3.1.
Consider a directed line graph ${\mathcal{G}}^{d}$ of $N$ nodes with directed edge weights equal to $1$, where the first (source) node is augmented with a self-loop of weight $1$. The symmetrized directed graph Laplacian ${\mathcal{L}}^{d}_{r}=({\mathbf{L}}^{d}_{r})^{\top}{\mathbf{L}}^{d}_{r}$, where ${\mathbf{L}}^{d}_{r}$ is the random-walk graph Laplacian for ${\mathcal{G}}^{d}$, is the same as graph Laplacian ${\mathbf{L}}^{u}$ for an undirected line graph ${\mathcal{G}}^{u}$ of $N$ nodes with edge weights equal to $1$.
See Appendix A for a formal proof.
Figure 2: Example of a 4-node DAG ${\mathcal{G}}^{d}$, with an added self-loop at node $1$, specified by ${\mathbf{L}}^{d}_{r}$ (a), and corresponding undirected graph ${\mathcal{G}}^{u}$, specified by ${\mathcal{L}}^{d}_{r}=({\mathbf{L}}^{d}_{r})^{\top}{\mathbf{L}}^{d}_{r}={\mathbf{L}}^{u}$ (b). Example of a 3-node DAG (c), and a 3-node DAG with opposite directional edges (d).
Remark: Using our symmetrized ${\mathcal{L}}^{d}_{r}$ does not mean directionality has been lost. For the 3-node DAG in Fig. 2(c), the variational term is $\|{\mathbf{L}}^{d}_{r}{\mathbf{x}}\|^{2}_{2}=(x_{3}-\frac{1}{2}(x_{1}+x_{2}))^{2}$ and computes to 00 for ${\mathbf{x}}=[2~{}~{}0~{}~{}1]^{\top}$. On the other hand, for the DAG in Fig. 2(d) with opposite directions, the variational term is $\|{\mathbf{L}}^{d}_{r}{\mathbf{x}}\|^{2}_{2}=(x_{1}-x_{3})^{2}+(x_{2}-x_{3})^{2}$ and computes to 00 only if ${\mathbf{x}}$ is a constant vector.
3.2.2 $\ell_{1}$-Norm Variational Term
Using ${\mathbf{W}}^{d}_{r}$ again as a GSO, we define an $\ell_{1}$-norm variational term called directed graph total variation (DGTV) as
$\|{\mathbf{x}}-{\mathbf{W}}^{d}_{r}{\mathbf{x}}\|_{1}=\|{\mathbf{L}}^{d}_{r}{\mathbf{x}}\|_{1}=\sum_{j\in\bar{{\mathcal{S}}}}|x_{j}-\sum_{i}w_{j,i}x_{i}|.$ (3)
Unlike DGLR in (2), DGTV is not symmetric and has no obvious frequency interpretation. Nonetheless, we provide a two-channel filterbank interpretation in the sequel.
3.3 Optimization Formulation
Denote by ${\mathbf{y}}\in\mathbb{R}^{M}$ the observation, ${\mathbf{x}}\in\mathbb{R}^{N(T+S+1)}$ the target signal, and ${\mathbf{H}}\in\{0,1\}^{M\times N(T+S+1)}$ the sampling matrix that selects $M$ observed samples from $N(T+S+1)$ entries in ${\mathbf{x}}$. Given defined undirected and directed edges, we write an objective for ${\mathbf{x}}$ containing a squared-error fidelity term and graph smoothness terms for the two graphs ${\mathcal{G}}^{u}$ and ${\mathcal{G}}^{d}$—GLR, DGLR, and DGTV:
$\min_{\mathbf{x}}\|{\mathbf{y}}-{\mathbf{H}}{\mathbf{x}}\|^{2}_{2}+\mu_{u}{\mathbf{x}}^{\top}{\mathbf{L}}^{u}{\mathbf{x}}+\mu_{d,2}{\mathbf{x}}^{\top}{\mathcal{L}}_{r}^{d}{\mathbf{x}}+\mu_{d,1}\|{\mathbf{L}}^{d}_{r}{\mathbf{x}}\|_{1}$ (4)
where $\mu_{u},\mu_{d,2},\mu_{d,1}\in\mathbb{R}_{+}$ are weight parameters for the three regularization terms. Combination of $\ell_{2}$- and $\ell_{1}$-norm penalties—DGLR and DGTV for directed graph ${\mathcal{G}}^{d}$ in our case [^3] We can employ both $\ell_{2}$- and $\ell_{1}$-norm regularization terms for the undirected graph ${\mathcal{G}}^{u}$ as well. For simplicity, we focus on designing new regularization terms for directed graphs in this paper. —is called elastic net regularization in statistics [27] with demonstrable improved robustness. Further, employing both regularization terms means that, after algorithm unrolling, we can adapt an appropriate mixture by learning weights $\mu_{d,2}$ and $\mu_{d,1}$ per neural layer.
(4) is a convex minimization problem composed of convex smooth $\ell_{2}$-norm terms and one convex non-smooth $\ell_{1}$-norm term. We pursue a divide-and-conquer approach and solve (4) via an ADMM framework [21] by minimizing $\ell_{2}$- and $\ell_{1}$-norm terms in (4) alternately.
3.4 ADMM Optimization Algorithm
We first introduce auxiliary variable ${\boldsymbol{\phi}}$ and rewrite the optimization in (4) as
$\min_{{\mathbf{x}},{\boldsymbol{\phi}}}$ $\|{\mathbf{y}}-{\mathbf{H}}{\mathbf{x}}\|^{2}_{2}+\mu_{u}{\mathbf{x}}^{\top}{\mathbf{L}}^{u}{\mathbf{x}}+\mu_{d,2}{\mathbf{x}}^{\top}{\mathcal{L}}_{r}^{d}{\mathbf{x}}+\mu_{d,1}\|{\boldsymbol{\phi}}\|_{1},\qquad\mbox{s.t.}~{}~{}~{}{\boldsymbol{\phi}}={\mathbf{L}}^{d}_{r}{\mathbf{x}}.$ (5)
Using the augmented Lagrangian method [28], we rewrite (5) in an unconstrained form:
$\min_{{\mathbf{x}},{\boldsymbol{\phi}}}$ $\|{\mathbf{y}}-{\mathbf{H}}{\mathbf{x}}\|^{2}_{2}+\mu_{u}{\mathbf{x}}^{\top}{\mathbf{L}}^{u}{\mathbf{x}}+\mu_{d,2}{\mathbf{x}}^{\top}{\mathcal{L}}_{r}^{d}{\mathbf{x}}+\mu_{d,1}\|{\boldsymbol{\phi}}\|_{1}+{\boldsymbol{\gamma}}^{\top}({\boldsymbol{\phi}}-{\mathbf{L}}^{d}_{r}{\mathbf{x}})+\frac{\rho}{2}\|{\boldsymbol{\phi}}-{\mathbf{L}}^{d}_{r}{\mathbf{x}}\|^{2}_{2}$ (6)
where ${\boldsymbol{\gamma}}\in\mathbb{R}^{N(T+S+1)}$ is a Lagrange multiplier vector, and $\rho\in\mathbb{R}_{+}$ is a non-negative ADMM parameter. We solve (6) iteratively by minimizing variables ${\mathbf{x}}$ and ${\boldsymbol{\phi}}$ in alternating steps till convergence.
3.4.1 Minimizing ${\mathbf{x}}^{\tau+1}$
Fixing ${\boldsymbol{\phi}}^{\tau}$ and ${\boldsymbol{\gamma}}^{\tau}$ at iteration $\tau$, optimization in (6) for ${\mathbf{x}}^{\tau+1}$ becomes an unconstrained convex quadratic objective. The solution is a system of linear equations:
$\left({\mathbf{H}}^{\top}{\mathbf{H}}+\mu_{u}{\mathbf{L}}^{u}+(\mu_{d,2}+\frac{\rho}{2}){\mathcal{L}}^{d}_{r}\right){\mathbf{x}}^{\tau+1}$ $=({\mathbf{L}}_{r}^{d})^{\top}(\frac{\rho}{2}\,{\boldsymbol{\phi}}^{\tau}+\frac{{\boldsymbol{\gamma}}^{\tau}}{2})+{\mathbf{H}}^{\top}{\mathbf{y}}.$ (7)
Because coefficient matrix ${\mathbf{H}}^{\top}{\mathbf{H}}+\mu_{u}{\mathbf{L}}^{u}+(\mu_{d,2}+\frac{\rho}{2}){\mathcal{L}}^{d}_{r}$ is sparse, symmetric and PD, (7) can be solved in linear time via conjugate gradient (CG) [29] without matrix inversion.
Splitting $\ell_{2}$-norm Terms: Instead of solving (7) directly for ${\mathbf{x}}^{\tau+1}$ at iteration $\tau$, we introduce again auxiliary variables ${\mathbf{z}}_{u}$ and ${\mathbf{z}}_{d}$ in (5) for the $\ell_{2}$-norm GLR and DGLR terms respectively, and rewrite the optimization as
$\min_{{\mathbf{x}},{\mathbf{z}}_{u},{\mathbf{z}}_{d}}{}{}$ $\|{\mathbf{y}}-{\mathbf{H}}{\mathbf{x}}\|^{2}_{2}+\mu_{u}{\mathbf{z}}_{u}^{\top}{\mathbf{L}}^{u}{\mathbf{z}}_{u}+\mu_{d,2}{\mathbf{z}}_{d}^{\top}{\mathcal{L}}_{r}^{d}{\mathbf{z}}_{d}+({\boldsymbol{\gamma}}^{\tau})^{\top}({\boldsymbol{\phi}}^{\tau}-{\mathbf{L}}^{d}_{r}{\mathbf{x}})+\frac{\rho}{2}\|{\boldsymbol{\phi}}^{\tau}-{\mathbf{L}}^{d}_{r}{\mathbf{x}}\|^{2}_{2}$ (8)
s.t. $~{}{\mathbf{x}}={\mathbf{z}}_{u}={\mathbf{z}}_{d}.$
Using again the augmented Lagrangian method, we rewrite the unconstrained version as
$\min_{{\mathbf{x}},{\mathbf{z}}_{u},{\mathbf{z}}_{d}}$ $\|{\mathbf{y}}-{\mathbf{H}}{\mathbf{x}}\|^{2}_{2}+\mu_{u}{\mathbf{z}}_{u}^{\top}{\mathbf{L}}^{u}{\mathbf{z}}_{u}+\mu_{d,2}{\mathbf{z}}_{d}^{\top}{\mathcal{L}}_{r}^{d}{\mathbf{z}}_{d}+({\boldsymbol{\gamma}}^{\tau})^{\top}({\boldsymbol{\phi}}^{\tau}-{\mathbf{L}}^{d}_{r}{\mathbf{x}})+\frac{\rho}{2}\|{\boldsymbol{\phi}}^{\tau}-{\mathbf{L}}^{d}_{r}{\mathbf{x}}\|^{2}_{2}$ (9)
$+({\boldsymbol{\gamma}}_{u}^{\tau})^{\top}({\mathbf{x}}-{\mathbf{z}}_{u})+\frac{\rho_{u}}{2}\|{\mathbf{x}}-{\mathbf{z}}_{u}\|^{2}_{2}+({\boldsymbol{\gamma}}_{d}^{\tau})^{\top}({\mathbf{x}}-{\mathbf{z}}_{d})+\frac{\rho_{d}}{2}\|{\mathbf{x}}-{\mathbf{z}}_{d}\|^{2}_{2}$
where ${\boldsymbol{\gamma}}_{u},{\boldsymbol{\gamma}}_{d}\in\mathbb{R}^{N(T+S+1)}$ are multipliers, and $\rho_{u},\rho_{d}\in\mathbb{R}_{+}$ are ADMM parameters. Optimizing ${\mathbf{x}}$, ${\mathbf{z}}_{u}$ and ${\mathbf{z}}_{d}$ in (9) in turn at iteration $\tau$ lead to three linear systems:
${\footnotesize\left({\mathbf{H}}^{\top}{\mathbf{H}}+\frac{\rho}{2}{\mathcal{L}}^{d}_{r}+\frac{\rho_{u}+\rho_{d}}{2}{\mathbf{I}}\right)}{\mathbf{x}}^{\tau+1}$ $={\small({\mathbf{L}}^{d}_{r})^{\top}(\frac{{\boldsymbol{\gamma}}^{\tau}}{2}+\frac{\rho}{2}{\boldsymbol{\phi}}^{\tau})-\frac{{\boldsymbol{\gamma}}_{u}^{\tau}}{2}+\frac{\rho_{u}}{2}{\mathbf{z}}_{u}^{\tau}-\frac{{\boldsymbol{\gamma}}_{d}^{\tau}}{2}+\frac{\rho_{d}}{2}{\mathbf{z}}_{d}^{\tau}+{\mathbf{H}}^{\top}{\mathbf{y}}}$ (10)
${\small\left(\mu_{u}{\mathbf{L}}^{u}+\frac{\rho_{u}}{2}{\mathbf{I}}\right)}{\mathbf{z}}_{u}^{\tau+1}$ $={\small\frac{{\boldsymbol{\gamma}}_{u}^{\tau}}{2}+\frac{\rho_{u}}{2}{\mathbf{x}}^{\tau+1}}$ (11)
${\small\left(\mu_{d,2}{\mathcal{L}}^{d}_{r}+\frac{\rho_{d}}{2}{\mathbf{I}}\right)}{\mathbf{z}}_{d}^{\tau+1}$ $={\small\frac{{\boldsymbol{\gamma}}_{d}^{\tau}}{2}+\frac{\rho_{d}}{2}{\mathbf{x}}^{\tau+1}}.$ (12)
Thus, instead of solving (7) for ${\mathbf{x}}^{\tau+1}$ directly at iteration $\tau$, ${\mathbf{x}}^{\tau+1}$, ${\mathbf{z}}_{u}^{\tau+1}$ and ${\mathbf{z}}_{d}^{\tau+1}$ are optimized in turn via (10) through (12) using CG. This splitting induces more network parameters for data-driven learning after algorithm unrolling (see Section 4.1) towards better performance, and affords us spectral filter interpretations of the derived linear systems.
Interpretation: To interpret solution ${\mathbf{z}}_{u}^{\tau+1}$ spectrally, we rewrite (11) as
${\mathbf{z}}_{u}^{\tau+1}=\left(\frac{2\mu_{u}}{\rho_{u}}{\mathbf{L}}^{u}+{\mathbf{I}}\right)^{-1}\left(\frac{{\boldsymbol{\gamma}}_{u}}{\rho_{u}}+{\mathbf{x}}^{\tau+1}\right)={\mathbf{V}}\mathrm{diag}\left(\left\{\frac{1}{1+\frac{2\mu_{u}}{\rho_{u}}\lambda_{k}}\right\}\right){\mathbf{V}}^{\top}\left(\frac{{\boldsymbol{\gamma}}_{u}}{\rho_{u}}+{\mathbf{x}}^{\tau+1}\right)$ (13)
where ${\mathbf{L}}^{u}={\mathbf{V}}\mathrm{diag}(\{\lambda_{k}\}){\mathbf{V}}^{\top}$ is the eigen-decomposition. Thus, using frequencies defined by eigen-pairs $\{(\lambda_{k},{\mathbf{v}}_{k})\}$ of ${\mathbf{L}}^{u}$, ${\mathbf{z}}_{u}^{\tau+1}$ is the low-pass filter output of ${\mathbf{x}}^{\tau+1}$ offset by $\frac{{\boldsymbol{\gamma}}_{u}}{\rho_{u}}$, with frequency response $f(\lambda)=(1+\frac{2\mu_{u}}{\rho_{u}}\lambda)^{-1}$.
As done for ${\mathbf{z}}_{u}^{\tau+1}$, we can rewrite ${\mathbf{z}}_{d}^{\tau+1}$ in (12) as
${\mathbf{z}}_{d}^{\tau+1}$ $={\mathbf{U}}\mathrm{diag}\left(\left\{\frac{1}{1+\frac{2\mu_{d,2}}{\rho_{d}}\xi_{k}}\right\}\right){\mathbf{U}}^{\top}\left(\frac{{\boldsymbol{\gamma}}_{d}}{\rho_{d}}+{\mathbf{x}}^{\tau+1}\right)$ (14)
where ${\mathcal{L}}^{d}_{r}={\mathbf{U}}\mathrm{diag}(\{\xi_{k}\}){\mathbf{U}}^{\top}$ is the eigen-decomposition. Analogously, ${\mathbf{z}}_{d}^{\tau+1}$ is the low-pass filter output of ${\mathbf{x}}^{\tau+1}$ offset by $\frac{{\boldsymbol{\gamma}}_{d}}{\rho_{d}}$, with frequency response $f(\xi)=(1+\frac{2\mu_{d,2}}{\rho_{d}}\xi)^{-1}$.
To interpret solution ${\mathbf{x}}^{\tau+1}$ in (10) as a low-pass filter output, see Appendix B.
3.4.2 Minimizing ${\boldsymbol{\phi}}^{\tau+1}$
Fixing ${\mathbf{x}}^{\tau+1}$ at iteration $\tau$, optimizing ${\boldsymbol{\phi}}^{\tau+1}$ in (6) means
${\boldsymbol{\phi}}^{\tau+1}=$ $\arg\min_{{\boldsymbol{\phi}}}g({\boldsymbol{\phi}})=\mu_{d,1}\|{\boldsymbol{\phi}}\|_{1}+({\boldsymbol{\gamma}}^{\tau})^{\top}({\boldsymbol{\phi}}-{\mathbf{L}}^{d}_{r}{\mathbf{x}}^{\tau+1})+\frac{\rho}{2}\|{\boldsymbol{\phi}}-{\mathbf{L}}^{d}_{r}{\mathbf{x}}^{\tau+1}\|^{2}_{2}.$ (15)
(15) can be computed element-wise as follows (see Appendix C for a derivation):
$\delta$ $=({\mathbf{L}}^{d}_{r})_{i}{\mathbf{x}}^{\tau+1}-\rho^{-1}\gamma_{i}^{\tau}$ (16)
$\phi_{i}^{\tau+1}$ $=\text{sign}(\delta)\cdot\max(|\delta|-\rho^{-1}\mu_{d,1},\,0).$
Interpretation: Soft-thresholding in (16) to compute $\phi_{i}^{\tau+1}$ attenuates the $i$-th entry of ${\mathbf{L}}^{d}_{r}{\mathbf{x}}^{\tau+1}$. Given that graph Laplacians—second difference matrices on graphs—are high-pass filters [3], we interpret ${\mathbf{L}}^{d}_{r}$ as the high-pass channel of a two-channel filterbank [30], where the corresponding low-pass channel is ${\mathbf{I}}-{\mathbf{L}}^{d}_{r}={\mathbf{W}}^{d}_{r}$. Using (16) for processing means that the low-pass channel ${\mathbf{W}}^{d}_{r}$ is unused and thus preserved. For example, the constant (low-frequency) signal ${\mathbf{1}}$ passes through the low-pass channel undisturbed, i.e., ${\mathbf{1}}={\mathbf{W}}^{d}_{r}{\mathbf{1}}$, while it is filtered out completely by the high-pass channel, i.e., ${\mathbf{0}}={\mathbf{L}}^{d}_{r}{\mathbf{1}}$. Thus, attenuation of the high-pass channel means (16) is a low-pass filter.
3.4.3 Updating Lagrange Multipliers
For objective (6), we follow standard ADMM procedure to update Lagrange multiplier ${\boldsymbol{\gamma}}^{\tau+1}$ after each computation of ${\mathbf{x}}^{\tau+1}$ and ${\boldsymbol{\phi}}^{\tau+1}$:
${\boldsymbol{\gamma}}^{\tau+1}$ $={\boldsymbol{\gamma}}^{\tau}+\rho({\boldsymbol{\phi}}^{\tau+1}-{\mathbf{L}}^{d}_{r}{\mathbf{x}}^{\tau+1}).$ (17)
Splitting the $\ell_{2}$-norm terms means that additional multipliers ${\boldsymbol{\gamma}}_{u}^{\tau+1}$ and ${\boldsymbol{\gamma}}_{d}^{\tau+1}$ in objective (9) must also be updated after each computation of ${\mathbf{x}}^{\tau+1}$, ${\mathbf{z}}_{u}^{\tau+1}$, ${\mathbf{z}}_{d}^{\tau+1}$, and ${\boldsymbol{\phi}}^{\tau+1}$:
${\boldsymbol{\gamma}}_{u}^{\tau+1}={\boldsymbol{\gamma}}_{u}^{\tau}+\rho_{u}({\mathbf{x}}^{\tau+1}-{\mathbf{z}}_{u}^{\tau+1}),~{}~{}~{}~{}~{}~{}~{}~{}~{}{\boldsymbol{\gamma}}_{d}^{\tau+1}={\boldsymbol{\gamma}}_{d}^{\tau}+\rho_{d}({\mathbf{x}}^{\tau+1}-{\mathbf{z}}_{d}^{\tau+1}).$ (18)
The ADMM loop is repeated till ${\mathbf{x}}^{\tau+1}$ and ${\boldsymbol{\phi}}^{\tau+1}$ converge.
4 Unrolled Neural Network
Figure 3: Unrolling of proposed iterative ADMM algorithm into blocks and neural layers.
We unroll our iterative ADMM algorithm to an interpretable neural net. The crux to our unrolled neural net is two periodically inserted graph learning modules for undirected and directed edge weight computation that are akin to the self-attention mechanism in classical transformers [19].
4.1 ADMM Algorithm Unrolling
We implement each iteration of our algorithm in Section 3.4 as a neural layer. Specifically, each ADMM iteration is implemented as four sub-layers in sequence: solving linear systems (10) for ${\mathbf{x}}^{\tau+1}$, (11) for ${\mathbf{z}}_{u}^{\tau+1}$, and (12) for ${\mathbf{z}}_{d}^{\tau+1}$ via CG, and computing ${\boldsymbol{\phi}}^{\tau+1}$ via (16). CG is itself an iterative descent algorithm, where parameters $\alpha$ and $\beta$ corresponding to step size and momentum can be tuned end-to-end (see Appendix D for detailed implementation). Weight parameters $\mu_{u},\mu_{d,2},\mu_{d,1}$ for prior terms and ADMM parameters $\rho,\rho_{u},\rho_{d}$ are also tuned per layer. Multipliers ${\boldsymbol{\gamma}}^{\tau+1},{\boldsymbol{\gamma}}_{u}^{\tau+1},{\boldsymbol{\gamma}}_{d}^{\tau+1}$ are then updated via (17) and (18) to complete one layer. Parameters learned during back-propagation for each ADMM block $b$ are denoted by ${\boldsymbol{\Theta}}_{b}$. See Fig. 3 for an illustration.
4.2 Self-Attention Operator in Transformer
We first review the basic self-attention mechanism: a scaled dot product following by a softmax operation [19]. Denote by ${\mathbf{x}}_{i}\in\mathbb{R}^{E}$ an embedding for token $i$. The affinity between tokens $i$ and $j$, $e(i,j)$, is the dot product of linear-transformed ${\mathbf{K}}{\mathbf{x}}_{i}$ and ${\mathbf{Q}}{\mathbf{x}}_{j}$, where ${\mathbf{Q}},{\mathbf{K}}\in\mathbb{R}^{E\times E}$ are the query and key matrices, respectively. Using softmax, non-negative attention weight $a_{i,j}$ is computed from $e(i,j)$’s as
$a_{i,j}=\frac{\exp(e(i,j))}{\sum_{l=1}^{N}\exp(e(i,l))},~{}~{}~{}~{}~{}~{}e(i,j)=({\mathbf{Q}}{\mathbf{x}}_{j})^{\top}({\mathbf{K}}{\mathbf{x}}_{i}).$ (19)
Given attention weights $a_{i,j}$, output embedding ${\mathbf{y}}_{i}$ for token $i$ is computed as
${\mathbf{y}}_{i}=\sum_{l=1}^{N}a_{i,l}{\mathbf{x}}_{l}{\mathbf{V}}$ (20)
where ${\mathbf{V}}\in\mathbb{R}^{E\times E}$ is a value matrix. By “self-attention”, we mean that output embeddings are computed from weighted input embeddings. A transformer is thus a sequence of embedding-to-embedding mappings via self-attention operations defined by learned ${\mathbf{Q}}$, ${\mathbf{K}}$ and ${\mathbf{V}}$ matrices.
4.3 Graph Learning Modules
Undirected Graph Learning: We learn an undirected graph ${\mathcal{G}}^{u}$ in a module $\texttt{UGL}_{b}$ at block $b$. For each node $i$, we compute a low-dimensional feature vector ${\mathbf{f}}_{i}^{u}=F^{u}({\mathbf{e}}_{i})\in\mathbb{R}^{K}$, where ${\mathbf{e}}_{i}\in\mathbb{R}^{E}$ is a vector of signal values, position embeddings in time and Laplacian embeddings of the physical graph as done in Feng and Tassiulas [9], and $F^{u}(\cdot)$ is a chosen learned nonlinear feature function $F^{u}:\mathbb{R}^{E}\mapsto\mathbb{R}^{K}$, e.g., a shallow GNN (see Appendix E.3 for details). As done in Thuc et al. [18], the feature distance between nodes $i$ and $j$ is computed as the Mahalanobis distance $\text{d}^{u}(i,j)$:
$\text{d}^{u}(i,j)=({\mathbf{f}}_{i}^{u}-{\mathbf{f}}_{j}^{u})^{\top}{\mathbf{M}}({\mathbf{f}}_{i}^{u}-{\mathbf{f}}_{j}^{u})$ (21)
where ${\mathbf{M}}\succeq 0$ is a symmetric PSD metric matrix of dimension $\mathbb{R}^{K\times K}$. Distance in (21) is non-negative and symmetric, i.e., $\text{d}^{u}(i,j)=\text{d}^{u}(j,i)$. Weight $w^{u}_{i,j}$ of undirected edge $(i,j)\in{\mathcal{E}}^{u}$ is computed as
$w^{u}_{i,j}=\frac{\exp\left(-\text{d}^{u}(i,j)\right)}{\sqrt{\sum_{l\in{\mathcal{N}}_{i}}\exp(-\text{d}^{u}(i,l))}\sqrt{\sum_{k\in{\mathcal{N}}_{j}}\exp(-\text{d}^{u}(k,j))}}$ (22)
where the denominator summing over 1-hop neighborhood ${\mathcal{N}}_{i}$ is inserted for normalization.
Remark: Interpreting distance $\text{d}^{u}(i,j)$ as $-e(i,j)$, edge weights $w^{u}_{i,j}$’s in (22) are essentially attention weights $a_{i,j}$’s in (19), and thus the undirected graph learning module constitutes a self-attention mechanism, albeit requiring fewer parameters, since parameters for function $F^{u}(\cdot)$ and metric matrix ${\mathbf{M}}$ can be much smaller than large and dense query and key matrices, ${\mathbf{Q}}$ and ${\mathbf{K}}$. Further, there is no value matrix ${\mathbf{V}}$ required, since the output signal ${\mathbf{x}}^{\tau+1}$ is computed through a set of low-pass filters—(10), (11), (12) and (16)—derived from our optimization of objective (4).
Directed Graph Learning: Similarly, we learn a directed graph ${\mathcal{G}}^{d}$ in a module $\texttt{DGL}_{b}$ at block $b$. For each node $i$, we compute a feature vector ${\mathbf{f}}^{d}_{i}=F^{d}({\mathbf{e}}_{i})\in\mathbb{R}^{K}$ using feature function $F^{d}(\cdot)$. Like (21), we compute Mahalanobis distance $\text{d}^{d}(i,j)$ between nodes $i$ and $j$, given feature vectors ${\mathbf{f}}_{i}^{d}$ and ${\mathbf{f}}_{j}^{d}$, using a learned PSD metric matrix ${\mathbf{P}}$. Weight $w_{j,i}^{d}$ of directed edge $[i,j]\in{\mathcal{E}}^{d}$ is then computed using exponential functions and normalization, as done in (22) for an undirected edge.
Parameters learned via back-propagation by the two graph learning modules at block $b$—parameters of $F^{u}(\cdot)$, $F^{d}(\cdot)$, ${\mathbf{M}}$ and ${\mathbf{P}}$—are denoted by ${\boldsymbol{\Phi}}_{b}$.
5 Experiments
5.1 Experimental Setup
We evaluate our model’s performance on traffic speed datasets METR-LA, PEMS-BAY [31], and traffic flow dataset PEMS03 [32], each of which contains traffic data with a sampling interval of 5 minutes in a local region, together with the real connections and distances (or travel cost) between sensors. We reduce the size of the original dataset by sub-sampling 1/3 of all data uniformly, and split it into training, validation, and test sets by 6:2:2. We predict the traffic speed/flow in the following 30 / 60 minutes (6 / 12 steps) from the previously observed 60 minutes (12 steps) for all datasets.
We restrict our model to learn a sparse graph based on the real road connections and a limited time window to model only the local spatial/temporal influence. For undirected graphs ${\mathcal{G}}^{u}$, we connect each node with its $k$ nearest neighbors (NNs). We construct 5 ADMM blocks, each contains 25 ADMM layers, and insert a graph learning module before each ADMM block. We set feature dimensions $K=6$, and stack vertically 4 graph learning modules to learn in parallel as an implementation of the multi-head attention mechanism. We minimize the Huber loss with $\delta=1$ for the entire reconstructed sequence and the groundtruths. See Appendix E for more details.
5.2 Experimental Results
Table 1: Comparison of RMSE / MAE / MAPE(%) metrics of our lightweight transformer to baseline models on 30 / 60-minute forecasting in PeMS03, METR-LA and PEMS-BAY datasets. We use boldface for the smallest error and underline the next 3 smallest errors.
Dataset & PEMS03 (358 nodes, 547 edges) METR-LA (207 nodes, 1315 edges) PEMS-BAY (325 nodes, 2090 edges)
Horizon 30 minutes 60 minutes 30 minutes 60 minutes 30 minutes 60 minutes
VAR [33] 28.07 / 16.53 / 17.49 30.54 / 18.31 / 19.61 10.72 / 5.55 / 11.29 12.59 / 6.99 / 13.46 3.81 / 2.08 / 4.46 3.98 / 2.56 / 5.52
STGCN [34] 28.06 / 18.24 / 17.76 37.31 / 24.29 / 23.00 11.97 / 5.43 / 13.45 14.77 / 6.86 / 17.51 4.32 / 2.65 / 5.21 7.39 / 4.26 / 8.25
STSGCN [35] 26.41 / 16.75 / 16.86 30.62 / 19.35 / 19.15 10.25 / 4.05 / 9.18 12.65 / 5.18 / 11.48 3.43 / 1.54 / 3.28 4.50 / 1.93 / 4.42
ST-Wave [36] 25.57 / 15.11 / 15.04 28.65 / 16.81 / 19.24 10.81 / 4.11 / 9.16 13.24 / 5.33 / 11.32 3.53 / 1.49 / 3.18 4.30 / 1.82 / 4.33
ASTTN [9] 26.84 / 17.04 / 21.46 27.24 / 16.43 / 15.63 11.49 / 4.80 / 10.97 11.10 / 5.08 / 12.51 5.07 / 2.23 / 4.65 4.82 / 2.13 / 4.40
PDFormer [22] 23.71 / 15.05 / 18.16 27.16 / 17.26 / 21.21 10.21 / 3.89 / 8.50 12.17 / 4.81 / 10.92 3.18 / 1.41 / 2.99 4.55 / 2.00 / 4.88
STAEformer [37] 30.22 / 18.85 / 26.62 38.36 / 23.68 / 29.21 10.16 / 3.73 / 8.25 12.58 / 4.79 / 10.08 4.07 / 1.75 / 3.95 5.28 / 2.28 / 5.59
Ours 26.10 / 17.03 / 18.85 27.67 / 17.46 / 17.72 10.35 / 4.22 / 9.70 12.37 / 5.27 / 11.91 3.78 / 1.70 / 4.08 4.60 / 2.03 / 4.79
Table 2: Comparison of the number of parameters between our model and baseline deep learning models for 60-minute forecasting in PEMS03 dataset.
Model Ours STGCN STSGCN ST-Wave ASTTN PDFormer STAEformer
Params # 34,096 321,941 3,496,212 882,558 161,635 531,165 1,403,940
We train each run of our model on NVIDIA GeForce RTX 3090. We evaluate our unrolling model against model-based method VAR [33], GNN-based models STGCN [^4] STGCN predicts only one time step at a time. We train the model to predict only the next step, and predict the entire sequence recursively. [34] and STSGCN [35], GAT-based models ST-Wave [36] and ASTTN [9], and transformer-based models PDFormer [22] and STAEformer [37] with 70 epochs of training on the reduced dataset with the Root Mean Squared Error (RMSE), Mean Absolute Error (MAE), and Mean Absolute Percentage Error (MAPE) in the predicted 6 / 12 steps.
Table 1 shows the performance of our models and baselines in traffic forecasting, and Table 2 compares the number of parameters. We observe that our models reach comparable performance to most baseline models, while employ drastically fewer parameters (6.4% of transformer-based PDFormer). Among all the models, ST-Wave generally achieves the best performance, which is attributed to an application-specific usage of wavelet transforms, but requires even more parameters than transformers. The largest model STSGCN fails to achieve satisfying performance for PEMS03, where the real connections are highly sparse. STAEformer, the largest transformer, also overfits heavily in the reduced PEMS03 and PEMS-BAY datasets. We also observe significant oscillations in validation loss during training in ASTTN due to the adaptive attention maps. In summary, none of the DL-based baseline schemes unambiguously achieves the best performance across the three datasets, while all require significantly more parameters than our models.
5.3 Ablation Study
Next, we discuss the effectiveness of the new directed graph priors, DGLR and DGTV, as well as the benefit of directed-graph modeling of temporal relations. First, we separately test the effects of DGLR and DGTV by removing the prior and unrolling the ADMM algorithm. We denote these models as “w/o DGTV” and “w/o DGLR”. To test the benefit of the directed-graph modeling, we change the temporal graph ${\mathcal{G}}^{d}$ into a fully undirected graph ${\mathcal{G}}^{n}$ while maintaining the connections, and replace the directed graph priors with a GLR prior ${\mathbf{x}}^{\top}{\mathbf{L}}^{n}{\mathbf{x}}$ with the same coefficient $\mu_{n}=\mu_{d,2}$. We denote the undirected-time unrolling model as “UT”. Detailed derivations of the ADMM algorithm for all the above ablation studies are in Appendix F.
Table 3: Performance comparison in RMSE / MAE / MAPE(%) of the ablation studies of directed-graph modeling and directed graph priors in 60-minute forecasting with PEMS03 and METR-LA.
Dataset Mixed-Graph UT w/o DGTV w/o DGLR
PEMS03 27.67 / 17.46 / 17.72 27.52 / 17.87 / 18.82 27.78 / 17.85 / 17.90 30.89 / 20.02 / 21.10
METR-LA 12.37 / 5.27 / 11.91 12.51 / 5.42 /12.11 12.36 / 5.40 / 12.31 12.41 / 5.35 / 12.20
As is summarized in Table 3, removing the DGLR or DGTV term leads to noticeable drops in performance, which underscores their importance as directed graph priors. Moreover, our model outperforms the fully-undirected-graph-based model, demonstrating that sequential relationships—such as time—are better modeled as directed graphs.
For further ablation studies in the number of nearest neighbors $k$, size of time window $W$, number of ADMM blocks and layers, and the splitting-term strategy in (8), see Appendix G.
6 Conclusion
To forecast city traffic, we unroll a mixed-graph optimization algorithm into a lightweight transformer-like neural net, where an undirected graph models spatial correlations and a directed graph models temporal relationships. We show that the two graph learning modules are akin to the classical self-attention mechanism. We design new $\ell_{2}$ and $\ell_{1}$-norm regularizers to quantify and promote signal smoothness on directed graphs. We interpret the derived processing operations from a constructed ADMM algorithm as low-pass filters. Experiments demonstrate competitive prediction performance at drastically reduced parameter counts. One limitation of our approach is that our computed Mahalanobis distances $\text{d}^{u}(i,j)$ and $\text{d}^{d}(i,j)$ for respective undirected and directed graphs are non-negative (due to learned PSD metric matrices ${\mathbf{M}}$ and ${\mathbf{P}}$), whereas affinity $e(i,j)$ in traditional self-attention can be negative. We will study extensions to signed distances as future work.
Appendix A Proof of Theorem 3.1
We prove Theorem 3.1 by construction as follows. A directed line graph ${\mathcal{G}}^{d}$ of $N$ nodes with inter-node edge weights $1$ (and self-loop of weight $1$ at node 1) has adjacency matrix ${\mathbf{W}}^{d}_{r}$ and random-walk Laplacian matrix ${\mathbf{L}}^{d}_{r}={\mathbf{I}}_{N}-{\mathbf{W}}^{d}_{r}$ of the following forms:
${\mathbf{W}}^{d}_{r}$ $=\left[\begin{array}[]{ccccc}1&0&\ldots&&0\\ 1&0&\ldots&&0\\ 0&1&0&\ldots&0\\ \vdots&\ddots&\ddots&&0\\ 0&\ldots&0&1&0\end{array}\right],~{}~{}~{}{\mathbf{L}}^{d}_{r}=\left[\begin{array}[]{ccccc}0&0&\ldots&&0\\ -1&1&\ldots&&0\\ 0&-1&1&\ldots&0\\ \vdots&\ddots&\ddots&&0\\ 0&\ldots&0&-1&1\end{array}\right].$ (33)
Note that rows $2$ to $N$ of ${\mathbf{L}}^{d}_{r}$ actually compose the incidence matrix of an undirected line graph ${\mathcal{G}}^{u}$ with edge weights $1$. Thus, the symmetrized directed graph Laplacian ${\mathcal{L}}^{d}_{r}=({\mathbf{L}}^{d}_{r})^{\top}{\mathbf{L}}^{d}_{r}$ is also the combinatorial graph Laplacian ${\mathbf{L}}^{u}$:
${\mathcal{L}}^{d}_{r}={\mathbf{L}}^{u}$ $=\left[\begin{array}[]{cccccc}1&-1&0&\ldots&&0\\ -1&2&-1&0&\ldots&0\\ 0&-1&2&-1&0\ldots&0\\ \vdots&&\ddots&\ddots&\ddots&\\ 0&\ldots&0&-1&2&-1\\ 0&\ldots&&0&-1&1\end{array}\right].$ (40)
As an illustration, consider a 4-node directed line graph ${\mathcal{G}}^{d}$ with edge weights $1$; see Fig. 2(a) for an illustration. We see that the symmetrized directed graph Laplacian ${\mathcal{L}}^{d}_{r}$ defaults to the graph Laplacian ${\mathbf{L}}^{u}$ for a 4-node undirected line graph ${\mathcal{G}}^{u}$ (Fig. 2(b)):
${\mathbf{W}}^{d}_{r}=\left[\begin{array}[]{cccc}1&0&0&0\\ 1&0&0&0\\ 0&1&0&0\\ 0&0&1&0\end{array}\right],~{}{\mathbf{L}}^{d}_{r}=\left[\begin{array}[]{cccc}0&0&0&0\\ -1&1&0&0\\ 0&-1&1&0\\ 0&0&-1&1\end{array}\right],{\mathcal{L}}^{d}_{r}=\left[\begin{array}[]{cccc}1&-1&0&0\\ -1&2&-1&0\\ 0&-1&2&-1\\ 0&0&-1&1\end{array}\right].$ (53)
Appendix B Interpreting Linear System (10)
For notation simplicity, we first define
${\mathbf{r}}^{\tau}\triangleq\small({\mathbf{L}}^{d}_{r})^{\top}(\frac{{\boldsymbol{\gamma}}^{\tau}}{2}+\frac{\rho}{2}{\boldsymbol{\phi}}^{\tau})-\frac{{\boldsymbol{\gamma}}_{u}^{\tau}}{2}+\frac{\rho_{u}}{2}{\mathbf{z}}_{u}^{\tau}-\frac{{\boldsymbol{\gamma}}_{d}^{\tau}}{2}+\frac{\rho_{d}}{2}{\mathbf{z}}_{d}^{\tau}.$ (54)
Solution ${\mathbf{x}}^{\tau+1}$ to linear system (10) can now be written as
${\mathbf{x}}^{\tau+1}$ $=\left({\mathbf{H}}^{\top}{\mathbf{H}}+\frac{\rho}{2}{\mathcal{L}}^{d}_{r}+\frac{\rho_{u}+\rho_{d}}{2}{\mathbf{I}}\right)^{-1}({\mathbf{r}}^{\tau}+{\mathbf{H}}^{\top}{\mathbf{y}}).$ (55)
Define ~$\tilde{{\mathcal{L}}}_{r}^{d}\triangleq{\mathbf{H}}^{\top}{\mathbf{H}}+\frac{\rho}{2}{\mathcal{L}}^{d}_{r}$. Note that ${\mathbf{H}}^{\top}{\mathbf{H}}$ is a diagonal matrix with zeros and ones (corresponding to chosen sample indices) along its diagonal, and thus ~$\tilde{{\mathcal{L}}}^{d}_{r}$ is a scaled variant of ${\mathcal{L}}^{d}_{r}$ with the addition of selected self-loops of weight $1$. Given that ${\mathcal{L}}^{d}_{r}=({\mathbf{L}}^{d}_{r})^{\top}{\mathbf{L}}^{d}_{r}$ is real and symmetric, ~$\tilde{{\mathcal{L}}}^{d}_{r}$ is also real and symmetric. By eigen-decomposing ~$\tilde{{\mathcal{L}}}^{d}_{r}=\tilde{{\mathbf{U}}}\mathrm{diag}(\{\tilde{\xi}_{k}\})\tilde{{\mathbf{U}}}^{\top}$, we can rewrite ${\mathbf{x}}^{\tau+1}$ as
${\mathbf{x}}^{\tau+1}$ $=\tilde{{\mathbf{U}}}\mathrm{diag}\left(\left\{\frac{1}{\frac{\rho_{u}+\rho_{d}}{2}+\tilde{\xi}_{k}}\right\}\right)\tilde{{\mathbf{U}}}^{\top}\left({\mathbf{r}}^{\tau}+{\mathbf{H}}^{\top}{\mathbf{y}}\right)$ (56)
which demonstrates that ${\mathbf{x}}^{\tau+1}$ is a low-pass filter output of up-sampled ${\mathbf{H}}^{\top}{\mathbf{y}}$ (with bias ${\mathbf{r}}^{\tau}$).
Appendix C Deriving Solution to Optimization (15)
The optimization for ${\boldsymbol{\phi}}$ in objective (15) can be performed entry-by-entry. Specifically, for the $i$-th entry $\phi_{i}$:
$\min_{\phi_{i}}~{}~{}g(\phi_{i})=\mu_{d,1}|\phi_{i}|+\gamma_{i}^{\tau}\left(\phi_{i}-({\mathbf{L}}^{d}_{r})_{i}{\mathbf{x}}^{\tau+1}\right)+\frac{\rho}{2}\left(\phi_{i}-({\mathbf{L}}^{d}_{r})_{i}{\mathbf{x}}^{\tau+1}\right)^{2}$ (57)
where $({\mathbf{L}}^{d}_{r})_{i}$ denotes the $i$-row of ${\mathbf{L}}^{d}_{r}$. $g(\phi_{i})$ is convex and differentiable when $\phi_{i}\neq 0$. Specifically, when $\phi_{i}>0$, setting the derivative of $g(\phi_{i})$ w.r.t. $\phi_{i}$ to zero gives
$\mu_{d,1}+\gamma_{i}^{\tau}+\rho\left(\phi_{i}^{*}-({\mathbf{L}}^{d}_{r})_{i}{\mathbf{x}}^{\tau+1}\right)=0$
$\phi_{i}^{*}=({\mathbf{L}}^{d}_{r})_{i}{\mathbf{x}}^{\tau+1}-\rho^{-1}\gamma_{i}^{\tau}-\rho^{-1}\mu_{d,1}$ (58)
which holds only if $({\mathbf{L}}^{d}_{r})_{i}{\mathbf{x}}^{\tau+1}-\rho^{-1}\gamma_{i}^{\tau}>\rho^{-1}\mu_{d,1}$. On the other hand, when $\phi_{i}<0$,
$-\mu_{d,1}+\gamma_{i}^{\tau}+\rho\left(\phi_{i}^{*}-({\mathbf{L}}^{d}_{r})_{i}{\mathbf{x}}^{\tau+1}\right)=0$
$\phi_{i}^{*}=({\mathbf{L}}^{d}_{r})_{i}{\mathbf{x}}^{\tau+1}-\rho^{-1}\gamma_{i}^{\tau}+\rho^{-1}\mu_{d,1}$ (59)
which holds only if $({\mathbf{L}}^{d}_{r})_{i}{\mathbf{x}}^{\tau+1}-\rho^{-1}\gamma_{i}^{\tau}<-\rho^{-1}\mu_{d,1}$. Summarizing the two cases, we get
$\delta$ $=({\mathbf{L}}^{d}_{r})_{i}{\mathbf{x}}^{\tau+1}-\rho^{-1}\gamma_{i}^{\tau}$
$\phi_{i}^{*}$ $=\text{sign}(\delta)\cdot\max(|\delta|-\rho^{-1}\mu_{d,1},\,0)$ (60)
Alternative solution with subgradients: Denote $g_{1}({\boldsymbol{\phi}})=\mu_{d,1}\|{\boldsymbol{\phi}}\|_{1},g_{2}({\boldsymbol{\phi}})=({\boldsymbol{\gamma}}^{\tau})^{\top}({\boldsymbol{\phi}}-{\mathbf{L}}^{d}_{r}{\mathbf{x}}^{\tau+1})+\frac{\rho}{2}\|{\boldsymbol{\phi}}-{\mathbf{L}}^{d}_{r}{\mathbf{x}}^{\tau+1}\|_{2}^{2}$, the subgradient of $g({\boldsymbol{\phi}})$ is
$\partial g({\boldsymbol{\phi}})=\{{\mathbf{g}}+\nabla g_{2}({\boldsymbol{\phi}})|{\mathbf{g}}\in\partial g_{1}({\boldsymbol{\phi}})\}=\{\mu_{d,1}\sigma+{\boldsymbol{\gamma}}^{\tau}+\rho({\boldsymbol{\phi}}-{\mathbf{L}}^{d}_{r}{\mathbf{x}}^{\tau+1})|\sigma\in\partial(\|{\boldsymbol{\phi}}\|_{1})\},$ (61)
where
$\partial_{i}(\|{\boldsymbol{\phi}}\|_{1})=\left\{\begin{aligned} &{\text{sign}(\phi_{i})},&\phi_{i}\neq 0,\\ &[-1,1],&\phi_{i}=0.\end{aligned}\right.$ (62)
The minimal ${\boldsymbol{\phi}}^{*}$ satisfies ${\mathbf{0}}\in\partial g({\boldsymbol{\phi}}^{*})$, thus
$\mu_{d,1}+\gamma^{\tau}_{i}+\rho(\phi^{*}_{i}-({\mathbf{L}}^{d}_{r})_{i}{\mathbf{x}})=0,$ $\text{if}~{}~{}\phi^{*}_{i}>0,$
$-\mu_{d,1}+\gamma^{\tau}_{i}+\rho(\phi^{*}_{i}-({\mathbf{L}}^{d}_{r})_{i}{\mathbf{x}})=0,$ $\text{if}~{}~{}\phi^{*}_{i}<0,$ (63)
$\exists\sigma\in[-1,1],\text{s.t.}~{}~{}\mu_{d,1}\sigma+\gamma^{\tau}_{i}+\rho(\phi^{*}_{i}-({\mathbf{L}}^{d}_{r})_{i}{\mathbf{x}})=0,$ $\text{if}~{}~{}\phi^{*}_{i}=0,$
which gives
$\phi_{i}^{*}=\left\{\begin{aligned} &({\mathbf{L}}^{d}_{r})_{i}{\mathbf{x}}-\frac{\gamma_{i}^{\tau}}{\rho}-\frac{\mu_{d,1}}{\rho},&&({\mathbf{L}}^{d}_{r})_{i}{\mathbf{x}}-\frac{\gamma_{i}^{\tau}}{\rho}>\frac{\mu_{d,1}}{\rho}\\ &0,&&-\frac{\mu_{d,1}}{\rho}\leq({\mathbf{L}}^{d}_{r})_{i}{\mathbf{x}}-\frac{\gamma_{i}^{\tau}}{\rho}\leq\frac{\mu_{d,1}}{\rho}\\ &({\mathbf{L}}^{d}_{r})_{i}{\mathbf{x}}-\frac{\gamma_{i}^{\tau}}{\rho}+\frac{\mu_{d,1}}{\rho},&&({\mathbf{L}}^{d}_{r})_{i}{\mathbf{x}}-\frac{\gamma_{i}^{\tau}}{\rho}<-\frac{\mu_{d,1}}{\rho}\\ \end{aligned}\right..$ (64)
Thus the solution of (15) is
${\boldsymbol{\phi}}^{\tau+1}={\boldsymbol{\phi}}^{*}=\text{soft}_{\frac{\mu_{d,1}}{\rho}}\left({\mathbf{L}}^{d}_{r}{\mathbf{x}}-\frac{\gamma_{i}^{\tau}}{\rho}\right)=\text{sign}\left({\mathbf{L}}^{d}_{r}{\mathbf{x}}-\frac{\gamma_{i}^{\tau}}{\rho}\right)\max\left(\left|{\mathbf{L}}^{d}_{r}{\mathbf{x}}-\frac{\gamma_{i}^{\tau}}{\rho}\right|-\frac{\mu_{d,1}}{\rho},0\right).$ (65)
Appendix D Parameterizing Conjugated Gradient Method in Solving Linear Equations
A linear system ${\mathbf{A}}{\mathbf{x}}={\mathbf{b}}$ where ${\mathbf{A}}$ is a PD matrix can be solved iteratively by the conjugated gradient (CG) method, as is described in Algorithm 1. We parameterize the coefficients of the “gradients” $\alpha_{k}$ and “momentum” $\beta_{k}$ in each CG iteration to transform the iterative algorithm to stacked neural net layers. Note that $\alpha,\beta\geq 0$ in every iteration in the original algorithm, and should be kept non-negative in the unrolled version.
Algorithm 1 Conjugate Gradient (CG) Method to Solve ${\mathbf{A}}{\mathbf{x}}={\mathbf{b}}$
1:PSD matrix ${\mathbf{A}}\in\mathbb{R}^{n\times n}$, vector ${\mathbf{b}}\in\mathbb{R}^{n}$, initial guess ${\mathbf{x}}_{0}$
2:Approximate solution ${\mathbf{x}}$ such that $\|{\mathbf{A}}{\mathbf{x}}-{\mathbf{b}}\|\leq\epsilon$, where $\epsilon$ is the tolerance of convergence
3:${\mathbf{r}}_{0}\leftarrow{\mathbf{b}}-{\mathbf{A}}{\mathbf{x}}_{0}$triangleright Initial residual
4:${\mathbf{p}}_{0}\leftarrow{\mathbf{r}}_{0}$
5:$k\leftarrow 0$
6:while $\|{\mathbf{A}}{\mathbf{x}}_{k}-{\mathbf{b}}\|>\epsilon$ do
7:     $\boxed{\alpha_{k}\leftarrow\dfrac{{\mathbf{r}}_{k}^{\top}{\mathbf{r}}_{k}}{{\mathbf{p}}_{k}^{\top}{\mathbf{A}}{\mathbf{p}}_{k}}}$
8:     ${\mathbf{x}}_{k+1}\leftarrow{\mathbf{x}}_{k}+\alpha_{k}{\mathbf{p}}_{k}$
9:     ${\mathbf{r}}_{k+1}\leftarrow{\mathbf{r}}_{k}-\alpha_{k}{\mathbf{A}}{\mathbf{p}}_{k}$triangleright Dual residual
10:     $\boxed{\beta_{k}\leftarrow\dfrac{{\mathbf{r}}_{k+1}^{\top}{\mathbf{r}}_{k+1}}{{\mathbf{r}}_{k}^{\top}{\mathbf{r}}_{k}}}$
11:     ${\mathbf{p}}_{k+1}\leftarrow{\mathbf{r}}_{k+1}+\beta_{k}{\mathbf{p}}_{k}$
12:     $k\leftarrow k+1$
13:end while
14:return ${\mathbf{x}}_{k+1}$
Appendix E Model Setup
E.1 ADMM Blocks Setup
We initialize $\mu_{u},\mu_{d,1},\mu_{d,2}$ with 3, and the $\rho,\rho_{u},\rho_{d}$ with $\sqrt{N/(T+S+1)}$, where $N$ is the number of sensors and $T+S+1$ is the full length of the recovered signal (both the observed part and the part to be predicted). Those parameters are trained every ADMM layer. We initialize the CGD parameters (${\boldsymbol{\alpha}}$, ${\boldsymbol{\beta}}$)s as 0.08, and tune them for every CGD iteration. We clamp each element of the CGD step size ${\boldsymbol{\alpha}}$ to [0, 0.8] for stability. Each element of the momentum term coefficient ${\boldsymbol{\beta}}$ is clamped to be non-negative.
E.2 Graph Learning Modules Setup
We define multiple PD matrices ${\mathbf{P}}$s and ${\mathbf{M}}$s in our unrolling model due to the different impact strengths at different times. For the undirected graph, we define ${\mathbf{M}}^{t}$ as the Mahanabolis distance matrix for the undirected graph slice at time $t$ to vary the impacting strength between spatial neighbors in different time steps. For the directed graph, we assume that the impact strength varies with different intervals but remains uniform for different monitoring stations. Thus, we define ${\mathbf{P}}^{w}$ as the Mahanabolis distance matrix to represent the temporal influence of interval $w$ for each node. Therefore, we define $(T+S+1)$ learnable PSD matrices ${\mathbf{M}}^{0},\dots,{\mathbf{M}}^{T+S}$ for undirected graph learning, and $W$ PSD learnable matrices ${\mathbf{P}}^{1},\dots,{\mathbf{P}}^{W}$ for directed graph learning in total.
We learn each PSD matrix ${\mathbf{M}}$ in each UGL by learning a square matrix ${\mathbf{M}}_{0}\in\mathbb{R}^{K\times K}$ by ${\mathbf{M}}={\mathbf{M}}_{0}^{\top}{\mathbf{M}}_{0}$, and initialize ${\mathbf{M}}_{0}$ with a matrix whose diagonal is filled with 1.5. Similarly, we learn each PSD matrix ${\mathbf{P}}^{w}$ in DGL by learning ${\mathbf{P}}_{0}^{w}$ where ${\mathbf{P}}^{w}=({\mathbf{P}}_{0}^{w})^{\top}{\mathbf{P}}_{0}^{w}$. We initialize ${\mathbf{P}}_{0}^{w}$ with $(1+0.2w/W){\mathbf{I}}$ to set the edge weights representing influences of longer intervals slightly smaller.
We learn 4 mixed graphs in parallel with 4 sets of matrices ${\mathbf{M}}^{0},\dots,{\mathbf{M}}^{T+S}$ and ${\mathbf{P}}^{1},\dots,{\mathbf{P}}^{W}$ to implement the multi-head attention mechanism in conventional transformers. The output signals of all 4 channels are integrated to one by a linear layer at the end of each ADMM block.
E.3 Feature Extractors Setup
Input Laplacian embeddings Position embeddings Concatenate embedded input
(a) Input embedding layer.
Input Embedding $x$ Spatial GraphSAGE $\alpha(\cdot)$ Temporal History Aggregation $\alpha(\cdot)$features
(b) A layer of eature extractor.
Figure 4: Framework of our feature extractor. (a) We first propose a non-parametric input embedding layer to combine signal with spatial/temporal embeddings. (b) Then aggregate embedded inputs of spatial neighbors and history information in a time window to generate a layer of feature extractor.
We detailedly introduce our design of the feature extractor described in Section 4.3. Figure 4(b) shows the overall structure of our feature extractors.
E.3.1 Input Embeddings
We generate non-parametric spatial and temporal embeddings for each node in our mixed spatio-temporal graph, see Figure 4(a). For spatial embeddings, we first generate the adjacency matrix ${\mathbf{W}}^{s}$ for the real-world connection of monitor stations using the physical distances $\text{d}^{S}(i,j)$ with
$w_{i,j}^{S}=w_{j,i}^{S}=\exp\left[-\frac{(\text{d}^{S}(i,j))^{2}}{50\times\text{Var}(\text{d}^{S})}\right],$ (66)
where $\text{Var}(\text{d}^{S})$ denotes the variance of the physical distances. We use Laplacian embedding supposed by Dwivedi and Bresson [38], which uses the eigenvectors of the $k^{S}$-smallest non-zero eigenvalues of the Laplacian matrix ${\mathbf{L}}^{s}={\mathbf{D}}^{s}-{\mathbf{W}}^{s}$ to generate $k$-dimensional embeddings. Our experiments set $k^{S}=5$.
For temporal embeddings, we use the sine-cosine position embeddings[39]:
$e[t,2i]=$ $\sin(t/10000^{i})$
$e[t,2i+1]=$ $\cos(t/10000^{i}),i=0,1,\dots,\left\lfloor\frac{k^{T}}{2}\right\rfloor-1.$ (67)
We follow Feng and Tassiulas [9] to use real-time-stamp, time-in-day and day-in-week embeddings of dimensions 10, 6, 4.
We directly concatenate these spatial and temporal embedding to our signals ${\mathbf{x}}^{t}_{j}$ for node $j$ at time $t$ as ${\mathbf{e}}_{j}^{t}=[{\mathbf{x}}^{t}_{j};{\mathbf{e}}^{S}_{j};{\mathbf{e}}^{T}_{t}]$ for feature extractors, where ${\mathbf{e}}^{S}_{j}$ denote the spatial Laplacian embeddings of node $j$ of the physical graph, and ${\mathbf{e}}^{T}_{t}$ denotes the temporal position embeddings of time step $t$.
E.3.2 Feature Extraction from embeddings
We use a variant of GraphSAGE [40] to aggregate spatial features and introduce Temporal History Aggregation (THA) to aggregate spatial features. The GraphSAGE module aggregates the inputs of each node’s $k$-NNs with a simple and uniformed linear layer to generate $K$-dimension spatial features for each node of each time step. The THA module aggregates the embedded inputs of past $W$ time steps for each node with a uniformed linear layer, where $W$ is the size of the time window. For both GraphSAGE and THA modules, we pad zeros to inputs as placeholders, and use Swish [41] as acitvation functions with $\beta=0.8$.
E.4 Constructing Graphs ${\mathbf{W}}^{u,0},{\mathbf{W}}^{d,0}$ for the First ADMM Block
Our ADMM block recovers the full signal ${\mathbf{x}}$ from observed signal ${\mathbf{y}}$ with mixed graph ${\mathcal{G}}^{u},{\mathcal{G}}^{d}$ constructed from embedded full signal ${\mathbf{e}}=[{\mathbf{x}};{\mathbf{e}}^{S};{\mathbf{e}}^{T}]$. Thus, for the first block in ADMM, we need an initial extrapolation for the signals to predict.
We propose a simple module to make an initial guess of the signals: we modify the feature extractor described in Appendix E.3.2 to generate intermediate features from the embedded input of the observed part. Then we combine the feature and temporal dimensions, and use a simple linear layer with Swish activation to generate the initially extrapolated signal.
Note: Both the feature extractor and the initial prediction module are carefully designed to involve as few parameters as possible to make the entire model lightweight. Furthermore, low-parameter and simple feature engineering ensures that the effectiveness of our model is attributed to the unrolled blocks and graph learning modules, rather than to the trivial feature engineering parts.
E.5 Training Setup
We preprocess the dataset by standardizing the entire time series for each node in space, but minimize a loss between the real groundtruth signal and the rescaled whole output signal. We use an Adam optimizer with learning rate $5\times 10^{-4}$, together with a scheduler of reduction on validation loss with a reduction rate of 0.2 and patience of 5. We set different batch sizes for different datasets to leverage the most of the GPU memory (in this case, we use an entire Nvidia GeForce RTX 3090). Detailed selection of parameters such as number of nearest spatial neighbors $k$ and time window size $W$ is shown in Table 4.
Table 4: Training setup of graph construction and batch size for all traffic datasets.
Dataset PeMS03 PeMS04 METR-LA PEMS-BAY
$k$ 4 6 6 6
$W$ 6 6 6 4
Batch size 12 12 16 12
Appendix F Optimization Formulation and ADMM Algorithm in Ablation Studies
F.1 Optimization & Algorithm without DGTV Term
By removing the DGTV term $\|{\mathbf{L}}^{d}_{r}{\mathbf{x}}\|_{1}$ in (4), and introducing the auxiliary variables ${\mathbf{z}}_{u},{\mathbf{z}}_{d}$ as (9), we rewrite the unconstrained version of the optimization by augmented Lagrangian method as
$\min_{{\mathbf{x}},{\mathbf{z}}_{u},{\mathbf{z}}_{d}}$ $\|{\mathbf{y}}-{\mathbf{H}}{\mathbf{x}}\|_{2}^{2}+\mu_{u}{\mathbf{z}}_{u}^{\top}{\mathbf{L}}^{u}{\mathbf{z}}_{u}+\mu_{d,2}{\mathbf{z}}_{d}^{\top}{\mathcal{L}}^{d}_{r}{\mathbf{z}}_{d}+({\boldsymbol{\gamma}}_{u}^{\tau})^{\top}({\mathbf{x}}-{\mathbf{z}}_{u})+\frac{\rho_{u}}{2}\|{\mathbf{x}}-{\mathbf{z}}_{u}\|_{2}^{2}$
$+({\boldsymbol{\gamma}}_{d}^{\tau})^{\top}({\mathbf{x}}-{\mathbf{z}}_{d})+\frac{\rho_{d}}{2}\|{\mathbf{x}}-{\mathbf{z}}_{d}\|_{2}^{2}.$ (68)
We optimize ${\mathbf{x}},{\mathbf{z}}_{u}$ and ${\mathbf{z}}_{d}$ in turn at iteration $\tau$ as follows:
$\left({\mathbf{H}}^{\top}{\mathbf{H}}+\frac{\rho_{u}+\rho_{d}}{2}{\mathbf{I}}\right){\mathbf{x}}^{\tau+1}=$ ${\mathbf{H}}^{\top}{\mathbf{y}}-\frac{{\boldsymbol{\gamma}}_{u}^{\tau}+{\boldsymbol{\gamma}}_{d}^{\tau}}{2}+\frac{\rho_{u}}{2}{\mathbf{z}}_{u}^{\tau}+\frac{\rho_{d}}{2}{\mathbf{z}}_{d}^{\tau},$ (69)
$\left(\mu_{u}{\mathbf{L}}^{u}+\frac{\rho_{u}}{2}{\mathbf{I}}\right){\mathbf{z}}_{u}^{\tau+1}=$ $\frac{{\boldsymbol{\gamma}}_{u}^{\tau}}{2}+\frac{\rho_{u}}{2}{\mathbf{x}}^{\tau+1},$ (70)
$\left(\mu_{d,2}{\mathcal{L}}^{d}_{r}+\frac{\rho_{d}}{2}{\mathbf{I}}\right){\mathbf{z}}_{d}^{\tau+1}=$ $\frac{{\boldsymbol{\gamma}}_{d}^{\tau}}{2}+\frac{\rho_{d}}{2}{\mathbf{x}}^{\tau+1}.$ (71)
The updating of ${\boldsymbol{\gamma}}_{u}^{\tau},{\boldsymbol{\gamma}}_{d}^{\tau}$ follows (18).
F.2 Optimization & Algorithm without DGLR Term
By removing the DGLR term ${\mathbf{x}}{\mathcal{L}}^{d}_{r}{\mathbf{x}}$ in (4) and introduce only the auxiliary variables ${\boldsymbol{\phi}}$ in (5) and ${\mathbf{z}}_{u}$ in (9), we rewrite the unconstrained version of the optimization of ${\mathbf{x}}^{\tau+1}$ by augmented Lagrangian method as
$\min_{x,{\mathbf{z}}_{u}}\|{\mathbf{y}}-{\mathbf{H}}{\mathbf{x}}\|_{2}^{2}+\mu_{u}{\mathbf{z}}_{u}^{\top}{\mathbf{L}}^{u}{\mathbf{z}}_{u}+({\boldsymbol{\gamma}}^{\tau})^{\top}({\boldsymbol{\phi}}^{\tau}-{\mathbf{L}}^{d}_{r}{\mathbf{x}})+\frac{\rho}{2}\|{\boldsymbol{\phi}}^{\tau}-{\mathbf{L}}^{d}_{r}{\mathbf{x}}\|_{2}^{2}+({\boldsymbol{\gamma}}_{u}^{\tau})^{\top}({\mathbf{x}}-{\mathbf{z}}_{u})+\frac{\rho_{u}}{2}\|{\mathbf{x}}-{\mathbf{z}}_{u}\|_{2}^{2}$ (72)
We optimize ${\mathbf{x}}$ and ${\mathbf{z}}_{u}$ in turn at iteration $\tau$ as follows:
$\left({\mathbf{H}}^{\top}{\mathbf{H}}+\frac{\rho}{2}{\mathcal{L}}^{d}_{r}+\frac{\rho_{u}}{2}{\mathbf{I}}\right){\mathbf{x}}^{\tau+1}$ $={\mathbf{H}}^{\top}{\mathbf{y}}+({\mathbf{L}}^{d}_{r})^{\top}(\frac{\gamma^{\tau}}{2}+\frac{\rho}{2}{\boldsymbol{\phi}}^{\tau})-\frac{{\boldsymbol{\gamma}}_{u}^{\tau}}{2}+\frac{\rho_{u}}{2}{\mathbf{z}}_{u}^{\tau}$ (73)
$\left(\mu_{u}{\mathbf{L}}^{u}+\frac{\rho_{u}}{2}{\mathbf{I}}\right){\mathbf{z}}_{u}^{\tau+1}$ $=\frac{{\boldsymbol{\gamma}}_{u}^{\tau}}{2}+\frac{\rho_{u}}{2}{\mathbf{x}}^{\tau+1}$ (74)
The optimization of ${\boldsymbol{\phi}}^{\tau+1}$ is the same as (15), and the solution is the same as (16). The updating of ${\boldsymbol{\gamma}},{\boldsymbol{\gamma}}_{u}$ follows (17) and (18).
F.3 Algorithm Unrolling with an Undirected Temporal Graph
In this section, we are changing the directed temporal graph ${\mathcal{G}}^{d}$ into an undirected graph while maintaining the connections. Denote the undirected temporal graph as ${\mathcal{G}}^{n}$. We introduce the GLR operator ${\mathbf{x}}^{\top}{\mathbf{L}}^{n}{\mathbf{x}}$ for ${\mathcal{G}}^{n}$ on signal ${\mathbf{x}}$ where ${\mathbf{L}}^{n}$ is the normalized Laplacian matrix of ${\mathcal{G}}^{n}$. Thus the optimization problem is as follows:
$\min_{x}\|{\mathbf{y}}-{\mathbf{H}}{\mathbf{x}}\|_{2}^{2}+\mu_{u}{\mathbf{x}}^{\top}{\mathbf{L}}^{u}{\mathbf{x}}+\mu_{n}{\mathbf{x}}^{T}{\mathbf{L}}^{n}{\mathbf{x}}$ (75)
where $\mu_{u},\mu_{n}\in\mathbb{R}_{+}$ are weight parameters for the two GLR priors. The opitimization is similar to that in Appendix F.1. By a similar approach which introduces auxiliary variables ${\mathbf{z}}_{u}={\mathbf{x}},{\mathbf{z}}_{n}={\mathbf{x}}$, we optimize ${\mathbf{x}},{\mathbf{z}}_{u},{\mathbf{z}}_{n}$ as follows:
$\left({\mathbf{H}}^{\top}{\mathbf{H}}+\frac{\rho_{u}+\rho_{n}}{2}{\mathbf{I}}\right){\mathbf{x}}^{\tau+1}=$ ${\mathbf{H}}^{\top}{\mathbf{y}}-\frac{{\boldsymbol{\gamma}}_{u}^{\tau}+{\boldsymbol{\gamma}}_{n}^{\tau}}{2}+\frac{\rho_{u}}{2}{\mathbf{z}}_{u}^{\tau}+\frac{\rho_{n}}{2}{\mathbf{z}}_{n}^{\tau},$ (76)
$\left(\mu_{u}{\mathbf{L}}^{u}+\frac{\rho_{u}}{2}{\mathbf{I}}\right){\mathbf{z}}_{u}^{\tau+1}=$ $\frac{{\boldsymbol{\gamma}}_{u}^{\tau}}{2}+\frac{\rho_{u}}{2}{\mathbf{x}}^{\tau+1},$ (77)
$\left(\mu_{n}{\mathbf{L}}^{n}+\frac{\rho_{n}}{2}{\mathbf{I}}\right){\mathbf{z}}_{n}^{\tau+1}=$ $\frac{{\boldsymbol{\gamma}}_{n}^{\tau}}{2}+\frac{\rho_{n}}{2}{\mathbf{x}}^{\tau+1}.$ (78)
The updating of ${\boldsymbol{\gamma}}_{u}^{\tau},{\boldsymbol{\gamma}}_{n}^{\tau}$ is also similar to (18) as follows:
${\boldsymbol{\gamma}}_{u}^{\tau+1}={\boldsymbol{\gamma}}_{u}^{\tau}+\rho_{u}({\mathbf{x}}^{\tau+1}-{\mathbf{z}}_{u}^{\tau+1}),\quad{\boldsymbol{\gamma}}_{n}^{\tau+1}={\boldsymbol{\gamma}}_{n}^{\tau}+\rho_{n}({\mathbf{x}}^{\tau+1}-{\mathbf{z}}_{n}^{\tau+1}).$ (79)
The undirected temporal graph ${\mathcal{G}}^{n}$ is learned by an UGL described in Section 4.3, and is initialized similar to the spatial UGL described in Appendix E.2. The initialization of $\mu_{n},\rho_{n}$ and the CGD parameters in solving (78) follows the setup in Appendix E.1.
Appendix G Further Ablation Studies
G.1 Sensitivity of $k$, $W$ in Graph Construction
We change the number of spatial nearest neighbors $k=4,6,8$ and time window size $W=4,6,8$ and present the results for PEMS03 and METR-LA dataset in Table 5. The selected hyperparameters are marked with “*”. Decreasing or increasing $k,W$ leads to performance degradation. Our selection of $k,W$ are optimal among all test hyperparameters, considering the tradeoff with the cost of time and memory.
Table 5: Performance on 60-minute forecasting with PEMS03 (left) and METR-LA (right) under different $k,W$ settings.
PEMS03, 60-minute forecast METR-LA, 60-minute forecast
$k$ $W$ RMSE MAE MAPE(%) $k$ $W$ RMSE MAE MAPE(%)
 4*  6* 27.67 17.46 17.72  6*  6* 12.37 5.27 11.91
4 4 29.54 17.09 17.92 4 6 12.43 5.24 11.87
4 8 27.45 17.25 18.35 8 6 12.61 5.30 11.92
6 6 28.98 17.33 18.23 6 4 12.44 5.40 11.85
8 6 45.37 18.15 20.59 6 8 12.44 5.3525 11.96
G.2 Sensitivity of Number of Blocks and Layers in the Unrolling Model
We vary the number of blocks and ADMM layers for PEMS03 dataset and present the results in Table 6. The results show that reducing number of blocks or layers will increase the error in prediction. Our selected settings $n_{\text{blocks}}=5,n_{\text{layers}}=25$ is the optimal setting so far in consideration of the trade-off in computing memory and time cost.
Table 6: Performance on 60-minute forecast with PEMS03 dataset under different number of blocks and layers.
$n_{\text{blocks}}$ $n_{\text{layers}}$ PEMS03, 60-minute forecast METR-LA, 60-minute forecast
RMSE MAE MAPE(%) RMSE MAE MAPE(%)
 5*  25* 27.67 17.46 17.72 12.37 5.27 11.91
5 20 27.95 17.32 18.02 12.53 5.35 12.15
4 25 27.77 17.63 18.61 12.45 5.33 12.00
G.3 Ablation Study for the Splitting-$\ell_{2}$-term Strategy
We also ran ablations of the splitting-term strategy by optimizing ${\mathbf{x}}^{\tau+1}$ using only (7). CG in solving (7) converges much slower than in solving (10)-(12), and the unrolled model can become numerically unstable in training, indicating that splitting the $\ell_{2}$ terms can stabilize the training process.
Appendix H Error Bar for Our Model in Table 1
[…]
Table 7: $2\sigma$ error bar of the performance metrics of our model in 60-minute traffic forecast in all three datasets.
Dataset PeMS03 METR-LA PEMS-BAY
$2\sigma_{\text{RMSE}}$ 0.05 0.04 0.01
$2\sigma_{\text{MAE}}$ 0.02 0.02 0.01
$2\sigma_{\text{MAPE(\%)}}$ 0.03 0.04 0.10
Appendix I Reconstruction Error
We perform the spatial/temporal forecast by recovering the entire time-vertex signal, which is unusual compared to other deep models. Table 8 shows that the reconstruction error is much smaller than the predicted part, indicating that our model has extracted all spatial/temporal correlations from the observed part for prediction.
Table 8: Reconstruction RMSE for the observed signals in 60-minute forecast in all three datasets.
Dataset PEMS03 METR-LA PEMS-BAY
rec RMSE 1.44 0.49 0.39

### v2

@@
-Lightweight Transformer via Unrolling of Mixed Graph Algorithms for Traffic Forecast
+Lightweight and Interpretable Transformer via Mixed Graph Algorithm Unrolling for Traffic Forecast
 Abstract
-To forecast traffic with both spatial and temporal dimensions, we unroll a mixed-graph-based optimization algorithm into a lightweight and interpretable transformer-like neural net.
-Specifically, we construct two graphs: an undirected graph ${\mathcal{G}}^{u}$ capturing spatial correlations across geography, and a directed graph ${\mathcal{G}}^{d}$ capturing sequential relationships over time.
-We formulate a prediction problem for the future samples of signal ${\mathbf{x}}$, assuming it is “smooth” with respect to both ${\mathcal{G}}^{u}$ and ${\mathcal{G}}^{d}$, where we design new $\ell_{2}$ and $\ell_{1}$-norm variational terms to quantify and promote signal smoothness (low-frequency reconstruction) on a directed graph.
-We construct an iterative algorithm based on alternating direction method of multipliers (ADMM), and unroll it into a feed-forward network for data-driven parameter learning.
-We insert graph learning modules for ${\mathcal{G}}^{u}$ and ${\mathcal{G}}^{d}$, which are akin to the self-attention mechanism in classical transformers.
+Unlike conventional “black-box” transformers with classical self-attention mechanism, we build a lightweight and interpretable transformer-like neural net by unrolling a mixed-graph-based optimization algorithm to forecast traffic with spatial and temporal dimensions.
+We construct two graphs: an undirected graph ${\mathcal{G}}^{u}$ capturing spatial correlations across geography, and a directed graph ${\mathcal{G}}^{d}$ capturing sequential relationships over time.
+We predict future samples of signal $\mathbf{x}$}, assuming it is “smooth” with respect to both ${\mathcal{G}}^{u}$ and ${\mathcal{G}}^{d}$, where we design new $\ell_{2}$ and $\ell_{1}$-norm variational terms to quantify and promote signal smoothness (low-frequency reconstruction) on a directed graph.
+We design an iterative algorithm based on alternating direction method of multipliers (ADMM), and unroll it into a feed-forward network for data-driven parameter learning.
+We insert graph learning modules for ${\mathcal{G}}^{u}$ and ${\mathcal{G}}^{d}$ that play the role of self-attention.
 Experiments show that our unrolled networks achieve competitive traffic forecast performance as state-of-the-art prediction schemes, while reducing parameter counts drastically.
@@
 1 Introduction
-We study the problem of forecasting city traffic with spatial and temporal dimensions, e.g., predicting car density at a set of intersections 5 minutes into the future given measurements in the past.
-Optimizing prediction here means exploiting both spatial correlations across geography and sequential relationships in time of how the past measurements inform future observations.
-When modeling the pairwise relationships as a graph, the problem is also called time-vertex signal processing (TVSP) [1].
-Among model-based methods, one approach to TVSP is to filter the signal along the spatial and temporal dimensions independently [2]e.g., employ graph spectral filters in the graph signal processing (GSP) literature [3, 4] along the irregular spatial dimension, and traditional Fourier filters along the regular time dimension.
-Another approach is to model the spatial correlations across time as a sequence of slowly time-varying graphs, and then process each spatial signal at time $t$ separately, where the graph changes over time are constrained by the Frobenius norm [5], $\ell_{1}$-norm [6], or low-rankness [7] of consecutive adjacency matrices.
-However, neither approach exploits spatial and temporal relations simultaneously for optimal performance.
-Leveraging the prowess of modern deep learning (DL) models like transformers [8], recent data-driven efforts address TVSP by building elaborate transformer architectures to capture spatial and temporal relations for prediction [9, 10, 11].
-However, these schemes are inherently uninterpretable “black boxes” requiring large training data to tune millions of parameters.
-They are also fragile to covariate shift [12]: statistical mismatch between training and testing data.
-Graph attention networks (GAT) [13] and graph transformers [14] are adaptations of the basic self-attention mechanism in transformers for graph-structured data (e.g., computing output embeddings using only input embeddings in local graph neighborhoods).
-However, they are still intrinsically uninterpretable and maintain large parameter sizes.
-In contrast, algorithm unrolling [15] offers an alternative hybrid model-based / data-driven paradigm; by “unrolling” iterations of a model-based algorithm minimizing a well-defined optimization objective into neural layers stacked together to form a feed-forward network for data-driven parameter learning via back-propagation, the resulting network can be both mathematically interpretable and competitive in performance to state-of-the-art (SOTA) DL models.
-As examples, Yang et al.
-[16], Yu et al.
-[17] design optimization algorithms that unroll into transformer-like neural nets.
-Inspired by Yu et al.
-[17] and extending from Thuc et al.
-[18] [^1] While Thuc et al.
-[18] focuses on unrolling of undirected-graph-based algorithms into transformer-like neural nets for image interpolation, we extend the study to mixed graphs with both undirected and directed edges to capture spatial and temporal pairwise relations, respectively, for traffic forecast., we unroll a dual-graph-based optimization algorithm into a lightweight interpretable transformer-like neural net for the traffic forecast problem.
-Specifically, we propose a novel mixed-graph approach and learn undirected graph ${\mathcal{G}}^{u}$ and directed graph ${\mathcal{G}}^{d}$ from data, to capture spatial correlations across geography and sequential relationships over time, respectively.
-We show that our graph learning modules are akin to the basic self-attention mechanism [19], and thus our unrolled graph-based neural nets are transformers.
-Given the two learned graphs, we define a prediction objective for future samples in signal ${\mathbf{x}}$, assuming ${\mathbf{x}}$ is “smooth” with respect to (w.r.t.) both ${\mathcal{G}}^{u}$ and ${\mathcal{G}}^{d}$ in variational terms: graph Laplacian regularizer (GLR) [20] for undirected ${\mathcal{G}}^{u}$, and newly designed directed graph Laplacian regularizer (DGLR) and directed graph total variation (DGTV) for directed ${\mathcal{G}}^{d}$.
-We construct a corresponding iterative optimization algorithm based on alternating direction method of multipliers (ADMM) [21] that unrolls into neural layers for parameter learning.
-For the first time, our proposed DGLR and DGTV variational terms quantify and promote signal smoothness on a directed graph, with intuitive low-pass filter interpretations.
-Experiments show that our unrolled networks achieve competitive traffic forecast performance as SOTA prediction schemes, while employing drastically fewer parameters (6.4% of transformer-based PDFormer [22]).
+Transformer, based on the classical self-attention mechanism (Vaswani et al., 2017), is now a dominant deep learning (DL) architecture that has achieved state-of-the-art (SOTA) performance across multiple fields (Woźniak et al., 2020; Zamir et al., 2022).
+However, like other “off-the-shelf” DL models, transformer has a massive number of parameters, and its internal structure is not easily interpretable.
+Parameter reduction is a major industrial concern for practical computation and/or memory-constrained environments [^1] See a brief review of lightweight learning models for different applications in Appendix A.1..
+Algorithm unrolling (Monga et al., 2021) offers an alternative hybrid model-based / data-driven paradigm; by “unrolling” iterations of a model-based algorithm minimizing a well-defined optimization objective into neural layers stacked together to form a feed-forward network for data-driven parameter learning, the resulting network can be both mathematically interpretable [^2] Common in algorithm unrolling (Monga et al., 2021), “interpretability” here means that each neural layer corresponds to an iteration of an optimization algorithm minimizing a mathematically-defined objective.
+and competitive in performance.
+Notably, Yu et al.
+(2023) designs an algorithm minimizing a sparse rate reduction (SRR) objective that unrolls into a transformer-like neural net for image classification.
+Thuc et al.
+(2024) designs graph-based algorithms minimizing graph smoothness priors that unroll into transformer-like neural nets for image interpolation while reducing parameters.
+However, only a positive undirected graph model was used in Thuc et al.
+(2024) to capture simple pairwise correlations between neighboring pixels in a static image.
+Spatial-temporal data (e.g., traffic and weather) contains more complex node-to-node relationships: geographically near stations exhibit spatial correlations, while past observations influences future data, resulting in sequential relationships.
+In this paper, we study the design of transformers via graph algorithm unrolling for spatial-temporal data, using traffic forecast as a concrete example application.
+Among model-based methods, one approach to traffic forecast is to filter the signal along the spatial and temporal dimensions independently (Ramakrishna et al., 2020)e.g., employ graph spectral filters in the graph signal processing (GSP) literature (Ortega et al., 2018; Cheung et al., 2018) along the irregular spatial dimension, and traditional Fourier filters along the regular time dimension.
+Another approach is to model the spatial correlations across time as a sequence of slowly time-varying graphs, and then filter each spatial signal at time $t$ separately, where the changes in consecutive graph adjacency matrices in time are constrained by the Frobenius norm (Kalofolias et al., 2017), $\ell_{1}$-norm (Yamada et al., 2019), or low-rankness (Bagheri et al., 2024).
+However, neither approach exploits spatial and temporal relationships simultaneously for optimal performance.
+Recent DL-based efforts address traffic forecast by building elaborate transformer architectures to capture spatial and temporal relations (Feng & Tassiulas, 2022; Gao et al., 2022; Jin et al., 2024) [^3] A detailed review of DL models on traffic forecast is included in Appendix A.2..
+Like transformers used in other fields (Zamir et al., 2022), these are also black boxes with large parameter counts.
+Graph attention networks (GAT) (Velickovic et al., 2018) and graph transformers (Dwivedi & Bresson, 2021) are adaptations of the self-attention mechanism in transformers for graph-structured data (e.g., computing output embeddings using only input embeddings in local graph neighborhoods).
+However, they are still uninterpretable and maintain large parameter sizes.
+In contrast, we extend Thuc et al.
+(2024) to build lightweight “white-box” transformers that are fully interpretable via mixed-graph algorithm unrolling for traffic forecast.
+Specifically, we learn two graphs from data: i) undirected graph ${\mathcal{G}}^{u}$ to capture spatial correlations across geography, and ii) a directed graph ${\mathcal{G}}^{d}$ to capture sequential relationships over time.
+We show that our graph learning modules play the role of self-attention (Bahdanau et al., 2014), and thus our unrolled graph-based neural nets are transformers.
+Given the two learned graphs, we define a prediction objective for future samples in signal $\mathbf{x}$}, assuming $\mathbf{x}$} is “smooth” with respect to (w.r.t.) both ${\mathcal{G}}^{u}$ and ${\mathcal{G}}^{d}$ in variational terms: graph Laplacian regularizer (GLR) (Pang & Cheung, 2017) for undirected ${\mathcal{G}}^{u}$, and newly designed directed graph Laplacian regularizer (DGLR) and directed graph total variation (DGTV) for directed ${\mathcal{G}}^{d}$.
+We design a corresponding linear-time optimization algorithm based on alternating direction method of multipliers (ADMM) (Boyd et al., 2011) that unrolls into neural layers for parameter learning.
+For the first time, our DGLR and DGTV variational terms quantify and promote signal smoothness on a directed graph, with intuitive low-pass filter interpretations.
+Experiments show that our networks achieve competitive traffic forecast performance as SOTA prediction schemes, while employing drastically fewer parameters (6.4% of transformer-based PDFormer (Jiang et al., 2023)).
 2 Preliminaries
@@
 Denote by ${\mathbf{D}}^{u}\in\mathbb{R}^{N\times N}$ a diagonal degree matrix, where $D_{i,i}^{u}=\sum_{j}W_{j,i}^{u}$.
-The symmetric graph Laplacian matrix is ${\mathbf{L}}^{u}\triangleq{\mathbf{D}}^{u}-{\mathbf{W}}^{u}$ [3].
-${\mathbf{L}}^{u}$ is positive semi-definite (PSD) if all edge weights are non-negative $w^{u}_{i,j}\geq 0,\forall i,j$ [4].
+The symmetric graph Laplacian matrix is ${\mathbf{L}}^{u}\triangleq{\mathbf{D}}^{u}-{\mathbf{W}}^{u}$ (Ortega et al., 2018).
+${\mathbf{L}}^{u}$ is positive semi-definite (PSD) if all edge weights are non-negative $w^{u}_{i,j}\geq 0,\forall i,j$ (Cheung et al., 2018).
 The symmetric normalized graph Laplacian matrix is ${\mathbf{L}}^{u}_{n}\triangleq{\mathbf{I}}-({\mathbf{D}}^{u})^{-1/2}{\mathbf{W}}^{u}({\mathbf{D}}^{u})^{-1/2}$, while the asymmetric random-walk graph Laplacian matrix is ${\mathbf{L}}^{u}_{r}\triangleq{\mathbf{I}}-({\mathbf{D}}^{u})^{-1}{\mathbf{W}}^{u}$.
-One can define a graph spectrum by eigen-decomposing the real and symmetric graph Laplacian ${\mathbf{L}}^{u}$ (or normalized graph Laplacian ${\mathbf{L}}_{n}^{u}$), where the $k$-th eigen-pair $(\lambda_{k},{\mathbf{v}}_{k})$ is interpreted as the $k$-th graph frequency and Fourier mode, respectively [3].
-Given $\{{\mathbf{v}}_{k}\}$ are orthonormal vectors, one can decompose a signal ${\mathbf{x}}$ into its graph frequency components as ${\boldsymbol{\alpha}}={\mathbf{V}}^{\top}{\mathbf{x}}$, where ${\mathbf{L}}^{u}={\mathbf{V}}\mathrm{diag}(\{\lambda_{k}\}){\mathbf{V}}^{\top}$, and ${\mathbf{V}}^{\top}$ is the graph Fourier transform (GFT).
-The $\ell_{2}$-norm graph Laplacian regularizer (GLR) [20], ${\mathbf{x}}^{\top}{\mathbf{L}}^{u}{\mathbf{x}}=\sum_{(i,j)\in{\mathcal{E}}^{u}}w^{u}_{i,j}(x_{i}-x_{j})^{2}$, is used to regularize an ill-posed signal restoration problem, like denoising, to bias low-frequency signal reconstruction (i.e., signals consistent with similarity graph ${\mathcal{G}}^{u}$) given observation ${\mathbf{y}}$:
+One can define a graph spectrum by eigen-decomposing the real and symmetric graph Laplacian ${\mathbf{L}}^{u}$ (or normalized graph Laplacian ${\mathbf{L}}_{n}^{u}$), where the $k$-th eigen-pair $(\lambda_{k},{\mathbf{v}}_{k})$ is interpreted as the $k$-th graph frequency and Fourier mode, respectively (Ortega et al., 2018).
+Given $\{{\mathbf{v}}_{k}\}$ are orthonormal vectors, one can decompose a signal $\mathbf{x}$} into its graph frequency components as ${\bm{\alpha}}={\mathbf{V}}^{\top}{\mathbf{x}}$, where ${\mathbf{L}}^{u}={\mathbf{V}}\mathrm{diag}(\{\lambda_{k}\}){\mathbf{V}}^{\top}$, and $\mathbf{V}$$^{\top}$ is the graph Fourier transform (GFT).
+The $\ell_{2}$-norm graph Laplacian regularizer (GLR) (Pang & Cheung, 2017), ${\mathbf{x}}^{\top}{\mathbf{L}}^{u}{\mathbf{x}}=\sum_{(i,j)\in{\mathcal{E}}^{u}}w^{u}_{i,j}(x_{i}-x_{j})^{2}$, is used to regularize an ill-posed signal restoration problem, like denoising, to bias low-frequency signal reconstruction (i.e., signals consistent with similarity graph ${\mathcal{G}}^{u}$) given observation $\mathbf{y}$}:
 ${\mathbf{x}}^{*}$ $=\arg\min_{\mathbf{x}}\|{\mathbf{y}}-{\mathbf{x}}\|^{2}_{2}+\mu\,{\mathbf{x}}^{\top}{\mathbf{L}}^{u}{\mathbf{x}}$
 $=({\mathbf{I}}+\mu{\mathbf{L}}^{u})^{-1}{\mathbf{y}}={\mathbf{V}}\mathrm{diag}\left(\frac{1}{1+\mu\lambda_{1}},\ldots,\frac{1}{1+\mu\lambda_{N}}\right){\mathbf{V}}^{\top}{\mathbf{y}}.$ (1)
-$\mu$ is a bias-variance tradeoff parameter [23], and the low-pass filter response is $f(\lambda)=(1+\mu\lambda)^{-1}$.
+$\mu$ is a bias-variance tradeoff parameter, and the low-pass filter response is $f(\lambda)=(1+\mu\lambda)^{-1}$.
 Directed Graph Definitions:
 We define analogous notations for a directed graph, denoted by ${\mathcal{G}}^{d}({\mathcal{V}},{\mathcal{E}}^{d},{\mathbf{W}}^{d})$.
-$[i,j]\in{\mathcal{E}}^{d}$ implies a directed edge exists from node $i$ to $j$ with weight $w_{j,i}^{d}=W_{j,i}^{d}$.
+$i,j]\in{\mathcal{E}}^{d}$ implies a directed edge exists from node $i$ to $j$ with weight $w_{j,i}^{d}=W_{j,i}^{d}$.
 ${\mathbf{W}}^{d}\in\mathbb{R}^{N\times N}$ is the asymmetric adjacency matrix.
@@
 The directed random-walk graph Laplacian matrix is ${\mathbf{L}}^{d}_{r}\triangleq{\mathbf{I}}-({\mathbf{D}}^{d})^{-1}{\mathbf{W}}^{d}$.
-Similar frequency notion cannot be simply defined for directed graphs via eigen-decomposition of directed graph Laplacian ${\mathbf{L}}^{d}$, because ${\mathbf{L}}^{d}$ is asymmetric, and thus the Spectral Theorem [24] does not apply.
+Similar frequency notion cannot be simply defined for directed graphs via eigen-decomposition of directed graph Laplacian ${\mathbf{L}}^{d}$, because ${\mathbf{L}}^{d}$ is asymmetric, and thus the Spectral Theorem (Hawkins, 1975) does not apply.
 We circumvent this problem by designing new variational terms in Section 3.2.
@@
 3.2.1 $\ell_{2}$-norm Variational Term
-Interpreting ${\mathbf{W}}_{r}^{d}$ as a graph shift operator (GSO) [25] on ${\mathcal{G}}^{d}$, we first define an $\ell_{2}$-norm variational term [^2] A similar directed graph variational term was defined in Li et al.
-[26] towards an objective for directed graph sampling.
+Interpreting ${\mathbf{W}}_{r}^{d}$ as a graph shift operator (GSO) (Chen et al., 2015) on ${\mathcal{G}}^{d}$, we first define an $\ell_{2}$-norm variational term [^4] A similar directed graph variational term was defined in Li et al.
+(2023) towards an objective for directed graph sampling.
 We focus instead on directed graph signal restoration.
@@
 where symmetrized directed graph Laplacian matrix ${\mathcal{L}}_{r}^{d}\triangleq({\mathbf{L}}^{d}_{r})^{\top}{\mathbf{L}}^{d}_{r}$ is symmetric and PSD.
-We call ${\mathbf{x}}^{\top}{\mathcal{L}}^{d}_{r}{\mathbf{x}}$ the DGLR, which computes the sum of squared differences between each child $j\in\bar{{\mathcal{S}}}$ and its parents $i$ for $[i,j]\in{\mathcal{E}}^{d}$.
+We call ${\mathbf{x}}^{\top}{\mathcal{L}}^{d}_{r}{\mathbf{x}}$ the DGLR, which computes the sum of squared differences between each child $j\in\bar{{\mathcal{S}}}$ and its parents $i$ for $i,j]\in{\mathcal{E}}^{d}$.
 Note that DGLR computes to zero for the constant vector $\mathbf{1}$}:
@@
 Importantly, we prove that our frequency definition using DGLR defaults to pure sinusoids as frequencies in the unweighted directed line graph case.
-Thus, our algorithm minimizing DGLR includes TVSP schemes like Ramakrishna et al.
-[2] that analyze signals along the time dimension using classical Fourier filters as special cases.
+Thus, our algorithm minimizing DGLR includes GSP schemes like Ramakrishna et al.
+(2020) that analyze signals along the time dimension using classical Fourier filters as special cases.
 Theorem 3.1.
@@
 The symmetrized directed graph Laplacian ${\mathcal{L}}^{d}_{r}=({\mathbf{L}}^{d}_{r})^{\top}{\mathbf{L}}^{d}_{r}$, where ${\mathbf{L}}^{d}_{r}$ is the random-walk graph Laplacian for ${\mathcal{G}}^{d}$, is the same as graph Laplacian ${\mathbf{L}}^{u}$ for an undirected line graph ${\mathcal{G}}^{u}$ of $N$ nodes with edge weights equal to $1$.
-See Appendix A for a formal proof.
+See Appendix B for a formal proof.
 Figure 2: Example of a 4-node DAG ${\mathcal{G}}^{d}$, with an added self-loop at node $1$, specified by ${\mathbf{L}}^{d}_{r}$ (a), and corresponding undirected graph ${\mathcal{G}}^{u}$, specified by ${\mathcal{L}}^{d}_{r}=({\mathbf{L}}^{d}_{r})^{\top}{\mathbf{L}}^{d}_{r}={\mathbf{L}}^{u}$ (b).
@@
 For the 3-node DAG in Fig.
-2(c), the variational term is $\|{\mathbf{L}}^{d}_{r}{\mathbf{x}}\|^{2}_{2}=(x_{3}-\frac{1}{2}(x_{1}+x_{2}))^{2}$ and computes to 00 for ${\mathbf{x}}=[2~{}~{}0~{}~{}1]^{\top}$.
+2(c), the variational term is $\|{\mathbf{L}}^{d}_{r}{\mathbf{x}}\|^{2}_{2}=(x_{3}-\frac{1}{2}(x_{1}+x_{2}))^{2}$ and computes to 0 for ${\mathbf{x}}=[2~~0~~1]^{\top}$.
 On the other hand, for the DAG in Fig.
-2(d) with opposite directions, the variational term is $\|{\mathbf{L}}^{d}_{r}{\mathbf{x}}\|^{2}_{2}=(x_{1}-x_{3})^{2}+(x_{2}-x_{3})^{2}$ and computes to 00 only if ${\mathbf{x}}$ is a constant vector.
+2(d) with opposite directions, the variational term is $\|{\mathbf{L}}^{d}_{r}{\mathbf{x}}\|^{2}_{2}=(x_{1}-x_{3})^{2}+(x_{2}-x_{3})^{2}$ and computes to 0 only if $\mathbf{x}$} is a constant vector.
 3.2.2 $\ell_{1}$-Norm Variational Term
@@
 $\|{\mathbf{x}}-{\mathbf{W}}^{d}_{r}{\mathbf{x}}\|_{1}=\|{\mathbf{L}}^{d}_{r}{\mathbf{x}}\|_{1}=\sum_{j\in\bar{{\mathcal{S}}}}|x_{j}-\sum_{i}w_{j,i}x_{i}|.$ (3)
-Unlike DGLR in (2), DGTV is not symmetric and has no obvious frequency interpretation. Nonetheless, we provide a two-channel filterbank interpretation in the sequel.
+Unlike DGLR in eq. 2, DGTV is not symmetric and has no obvious frequency interpretation. Nonetheless, we provide a two-channel filterbank interpretation in the sequel.
 3.3 Optimization Formulation
@@
 For simplicity, we focus on designing new regularization terms for directed graphs in this paper.
-—is called elastic net regularization in statistics [27] with demonstrable improved robustness.
+—is called elastic net regularization in statistics (Zou & Hastie, 2005) with demonstrable improved robustness.
 Further, employing both regularization terms means that, after algorithm unrolling, we can adapt an appropriate mixture by learning weights $\mu_{d,2}$ and $\mu_{d,1}$ per neural layer.
-(4) is a convex minimization problem composed of convex smooth $\ell_{2}$-norm terms and one convex non-smooth $\ell_{1}$-norm term.
-We pursue a divide-and-conquer approach and solve (4) via an ADMM framework [21] by minimizing $\ell_{2}$- and $\ell_{1}$-norm terms in (4) alternately.
+Convex minimization in eq.
+4 is composed of smooth $\ell_{2}$-norm terms and one non-smooth $\ell_{1}$-norm term.
+We pursue a divide-and-conquer approach and solve eq.
+4 via an ADMM framework (Boyd et al., 2011) by minimizing $\ell_{2}$- and $\ell_{1}$-norm terms in eq.
+4 alternately.
 3.4 ADMM Optimization Algorithm
-We first introduce auxiliary variable ${\boldsymbol{\phi}}$ and rewrite the optimization in (4) as
+We first introduce auxiliary variable $\boldsymbol{\phi}$} and rewrite the optimization in eq. 4 as
 $\min_{{\mathbf{x}},{\boldsymbol{\phi}}}$ $\|{\mathbf{y}}-{\mathbf{H}}{\mathbf{x}}\|^{2}_{2}+\mu_{u}{\mathbf{x}}^{\top}{\mathbf{L}}^{u}{\mathbf{x}}+\mu_{d,2}{\mathbf{x}}^{\top}{\mathcal{L}}_{r}^{d}{\mathbf{x}}+\mu_{d,1}\|{\boldsymbol{\phi}}\|_{1},\qquad\mbox{s.t.}~~~{\boldsymbol{\phi}}={\mathbf{L}}^{d}_{r}{\mathbf{x}}.$ (5)
-Using the augmented Lagrangian method [28], we rewrite (5) in an unconstrained form:
-$\min_{{\mathbf{x}},{\boldsymbol{\phi}}}$ $\|{\mathbf{y}}-{\mathbf{H}}{\mathbf{x}}\|^{2}_{2}+\mu_{u}{\mathbf{x}}^{\top}{\mathbf{L}}^{u}{\mathbf{x}}+\mu_{d,2}{\mathbf{x}}^{\top}{\mathcal{L}}_{r}^{d}{\mathbf{x}}+\mu_{d,1}\|{\boldsymbol{\phi}}\|_{1}+{\boldsymbol{\gamma}}^{\top}({\boldsymbol{\phi}}-{\mathbf{L}}^{d}_{r}{\mathbf{x}})+\frac{\rho}{2}\|{\boldsymbol{\phi}}-{\mathbf{L}}^{d}_{r}{\mathbf{x}}\|^{2}_{2}$ (6)
-where ${\boldsymbol{\gamma}}\in\mathbb{R}^{N(T+S+1)}$ is a Lagrange multiplier vector, and $\rho\in\mathbb{R}_{+}$ is a non-negative ADMM parameter.
-We solve (6) iteratively by minimizing variables ${\mathbf{x}}$ and ${\boldsymbol{\phi}}$ in alternating steps till convergence.
+Using the augmented Lagrangian method (Boyd & Vandenberghe, 2004), we rewrite eq. 5 in an unconstrained form:
+$\min_{{\mathbf{x}},{\boldsymbol{\phi}}}$ $\|{\mathbf{y}}-{\mathbf{H}}{\mathbf{x}}\|^{2}_{2}+\mu_{u}{\mathbf{x}}^{\top}{\mathbf{L}}^{u}{\mathbf{x}}+\mu_{d,2}{\mathbf{x}}^{\top}{\mathcal{L}}_{r}^{d}{\mathbf{x}}+\mu_{d,1}\|{\boldsymbol{\phi}}\|_{1}+{\bm{\gamma}}^{\top}({\boldsymbol{\phi}}-{\mathbf{L}}^{d}_{r}{\mathbf{x}})+\frac{\rho}{2}\|{\boldsymbol{\phi}}-{\mathbf{L}}^{d}_{r}{\mathbf{x}}\|^{2}_{2}$ (6)
+where ${\bm{\gamma}}\in\mathbb{R}^{N(T+S+1)}$ is a Lagrange multiplier vector, and $\rho\in\mathbb{R}_{+}$ is a non-negative ADMM parameter.
+We solve eq.
+6 iteratively by minimizing $\mathbf{x}$} and $\boldsymbol{\phi}$} in alternating steps till convergence.
 3.4.1 Minimizing ${\mathbf{x}}^{\tau+1}$
-Fixing ${\boldsymbol{\phi}}^{\tau}$ and ${\boldsymbol{\gamma}}^{\tau}$ at iteration $\tau$, optimization in (6) for ${\mathbf{x}}^{\tau+1}$ becomes an unconstrained convex quadratic objective.
+Fixing ${\boldsymbol{\phi}}^{\tau}$ and ${\bm{\gamma}}^{\tau}$ at iteration $\tau$, optimization in eq.
+6 for ${\mathbf{x}}^{\tau+1}$ becomes an unconstrained convex quadratic objective.
 The solution is a system of linear equations:
-$\left({\mathbf{H}}^{\top}{\mathbf{H}}+\mu_{u}{\mathbf{L}}^{u}+(\mu_{d,2}+\frac{\rho}{2}){\mathcal{L}}^{d}_{r}\right){\mathbf{x}}^{\tau+1}$ $=({\mathbf{L}}_{r}^{d})^{\top}(\frac{\rho}{2}\,{\boldsymbol{\phi}}^{\tau}+\frac{{\boldsymbol{\gamma}}^{\tau}}{2})+{\mathbf{H}}^{\top}{\mathbf{y}}.$ (7)
-Because coefficient matrix ${\mathbf{H}}^{\top}{\mathbf{H}}+\mu_{u}{\mathbf{L}}^{u}+(\mu_{d,2}+\frac{\rho}{2}){\mathcal{L}}^{d}_{r}$ is sparse, symmetric and PD, (7) can be solved in linear time via conjugate gradient (CG) [29] without matrix inversion.
-Splitting $\ell_{2}$-norm Terms: Instead of solving (7) directly for ${\mathbf{x}}^{\tau+1}$ at iteration $\tau$, we introduce again auxiliary variables ${\mathbf{z}}_{u}$ and ${\mathbf{z}}_{d}$ in (5) for the $\ell_{2}$-norm GLR and DGLR terms respectively, and rewrite the optimization as
-$\min_{{\mathbf{x}},{\mathbf{z}}_{u},{\mathbf{z}}_{d}}{}{}$ $\|{\mathbf{y}}-{\mathbf{H}}{\mathbf{x}}\|^{2}_{2}+\mu_{u}{\mathbf{z}}_{u}^{\top}{\mathbf{L}}^{u}{\mathbf{z}}_{u}+\mu_{d,2}{\mathbf{z}}_{d}^{\top}{\mathcal{L}}_{r}^{d}{\mathbf{z}}_{d}+({\boldsymbol{\gamma}}^{\tau})^{\top}({\boldsymbol{\phi}}^{\tau}-{\mathbf{L}}^{d}_{r}{\mathbf{x}})+\frac{\rho}{2}\|{\boldsymbol{\phi}}^{\tau}-{\mathbf{L}}^{d}_{r}{\mathbf{x}}\|^{2}_{2}$ (8)
+$\hskip-10.00002pt\left({\mathbf{H}}^{\top}{\mathbf{H}}+\mu_{u}{\mathbf{L}}^{u}+(\mu_{d,2}+\frac{\rho}{2}){\mathcal{L}}^{d}_{r}\right){\mathbf{x}}^{\tau+1}$ $=({\mathbf{L}}_{r}^{d})^{\top}(\frac{\rho}{2}\,{\boldsymbol{\phi}}^{\tau}+\frac{{\bm{\gamma}}^{\tau}}{2})+{\mathbf{H}}^{\top}{\mathbf{y}}.$ (7)
+Because coefficient matrix ${\mathbf{H}}^{\top}{\mathbf{H}}+\mu_{u}{\mathbf{L}}^{u}+(\mu_{d,2}+\frac{\rho}{2}){\mathcal{L}}^{d}_{r}$ is sparse, symmetric and PD, eq.
+7 can be solved in linear time via conjugate gradient (CG) (Shewchuk, 1994) without matrix inversion.
+Splitting $\ell_{2}$-norm Terms: Instead of solving eq.
+7 directly for ${\mathbf{x}}^{\tau+1}$ at iteration $\tau$, we introduce again auxiliary variables ${\mathbf{z}}_{u}$ and ${\mathbf{z}}_{d}$ in eq.
+5 for the $\ell_{2}$-norm GLR and DGLR terms respectively, and rewrite the optimization as
+$\min_{{\mathbf{x}},{\mathbf{z}}_{u},{\mathbf{z}}_{d}}$ $\|{\mathbf{y}}-{\mathbf{H}}{\mathbf{x}}\|^{2}_{2}+\mu_{u}{\mathbf{z}}_{u}^{\top}{\mathbf{L}}^{u}{\mathbf{z}}_{u}+\mu_{d,2}{\mathbf{z}}_{d}^{\top}{\mathcal{L}}_{r}^{d}{\mathbf{z}}_{d}+({\bm{\gamma}}^{\tau})^{\top}({\boldsymbol{\phi}}^{\tau}-{\mathbf{L}}^{d}_{r}{\mathbf{x}})+\frac{\rho}{2}\|{\boldsymbol{\phi}}^{\tau}-{\mathbf{L}}^{d}_{r}{\mathbf{x}}\|^{2}_{2}$ (8)
 s.t. $~{\mathbf{x}}={\mathbf{z}}_{u}={\mathbf{z}}_{d}.$
 Using again the augmented Lagrangian method, we rewrite the unconstrained version as
-$\min_{{\mathbf{x}},{\mathbf{z}}_{u},{\mathbf{z}}_{d}}$ $\|{\mathbf{y}}-{\mathbf{H}}{\mathbf{x}}\|^{2}_{2}+\mu_{u}{\mathbf{z}}_{u}^{\top}{\mathbf{L}}^{u}{\mathbf{z}}_{u}+\mu_{d,2}{\mathbf{z}}_{d}^{\top}{\mathcal{L}}_{r}^{d}{\mathbf{z}}_{d}+({\boldsymbol{\gamma}}^{\tau})^{\top}({\boldsymbol{\phi}}^{\tau}-{\mathbf{L}}^{d}_{r}{\mathbf{x}})+\frac{\rho}{2}\|{\boldsymbol{\phi}}^{\tau}-{\mathbf{L}}^{d}_{r}{\mathbf{x}}\|^{2}_{2}$ (9)
-$+({\boldsymbol{\gamma}}_{u}^{\tau})^{\top}({\mathbf{x}}-{\mathbf{z}}_{u})+\frac{\rho_{u}}{2}\|{\mathbf{x}}-{\mathbf{z}}_{u}\|^{2}_{2}+({\boldsymbol{\gamma}}_{d}^{\tau})^{\top}({\mathbf{x}}-{\mathbf{z}}_{d})+\frac{\rho_{d}}{2}\|{\mathbf{x}}-{\mathbf{z}}_{d}\|^{2}_{2}$
-where ${\boldsymbol{\gamma}}_{u},{\boldsymbol{\gamma}}_{d}\in\mathbb{R}^{N(T+S+1)}$ are multipliers, and $\rho_{u},\rho_{d}\in\mathbb{R}_{+}$ are ADMM parameters.
-Optimizing ${\mathbf{x}}$, ${\mathbf{z}}_{u}$ and ${\mathbf{z}}_{d}$ in (9) in turn at iteration $\tau$ lead to three linear systems:
-${\footnotesize\left({\mathbf{H}}^{\top}{\mathbf{H}}+\frac{\rho}{2}{\mathcal{L}}^{d}_{r}+\frac{\rho_{u}+\rho_{d}}{2}{\mathbf{I}}\right)}{\mathbf{x}}^{\tau+1}$ $={\small({\mathbf{L}}^{d}_{r})^{\top}(\frac{{\boldsymbol{\gamma}}^{\tau}}{2}+\frac{\rho}{2}{\boldsymbol{\phi}}^{\tau})-\frac{{\boldsymbol{\gamma}}_{u}^{\tau}}{2}+\frac{\rho_{u}}{2}{\mathbf{z}}_{u}^{\tau}-\frac{{\boldsymbol{\gamma}}_{d}^{\tau}}{2}+\frac{\rho_{d}}{2}{\mathbf{z}}_{d}^{\tau}+{\mathbf{H}}^{\top}{\mathbf{y}}}$ (10)
-${\small\left(\mu_{u}{\mathbf{L}}^{u}+\frac{\rho_{u}}{2}{\mathbf{I}}\right)}{\mathbf{z}}_{u}^{\tau+1}$ $={\small\frac{{\boldsymbol{\gamma}}_{u}^{\tau}}{2}+\frac{\rho_{u}}{2}{\mathbf{x}}^{\tau+1}}$ (11)
-${\small\left(\mu_{d,2}{\mathcal{L}}^{d}_{r}+\frac{\rho_{d}}{2}{\mathbf{I}}\right)}{\mathbf{z}}_{d}^{\tau+1}$ $={\small\frac{{\boldsymbol{\gamma}}_{d}^{\tau}}{2}+\frac{\rho_{d}}{2}{\mathbf{x}}^{\tau+1}}.$ (12)
-Thus, instead of solving (7) for ${\mathbf{x}}^{\tau+1}$ directly at iteration $\tau$, ${\mathbf{x}}^{\tau+1}$, ${\mathbf{z}}_{u}^{\tau+1}$ and ${\mathbf{z}}_{d}^{\tau+1}$ are optimized in turn via (10) through (12) using CG.
+$\min_{{\mathbf{x}},{\mathbf{z}}_{u},{\mathbf{z}}_{d}}$ $\|{\mathbf{y}}-{\mathbf{H}}{\mathbf{x}}\|^{2}_{2}+\mu_{u}{\mathbf{z}}_{u}^{\top}{\mathbf{L}}^{u}{\mathbf{z}}_{u}+\mu_{d,2}{\mathbf{z}}_{d}^{\top}{\mathcal{L}}_{r}^{d}{\mathbf{z}}_{d}+({\bm{\gamma}}^{\tau})^{\top}({\boldsymbol{\phi}}^{\tau}-{\mathbf{L}}^{d}_{r}{\mathbf{x}})+\frac{\rho}{2}\|{\boldsymbol{\phi}}^{\tau}-{\mathbf{L}}^{d}_{r}{\mathbf{x}}\|^{2}_{2}$ (9)
+$+({\bm{\gamma}}_{u}^{\tau})^{\top}({\mathbf{x}}-{\mathbf{z}}_{u})+\frac{\rho_{u}}{2}\|{\mathbf{x}}-{\mathbf{z}}_{u}\|^{2}_{2}+({\bm{\gamma}}_{d}^{\tau})^{\top}({\mathbf{x}}-{\mathbf{z}}_{d})+\frac{\rho_{d}}{2}\|{\mathbf{x}}-{\mathbf{z}}_{d}\|^{2}_{2}$
+where ${\bm{\gamma}}_{u},{\bm{\gamma}}_{d}\in\mathbb{R}^{N(T+S+1)}$ are multipliers, and $\rho_{u},\rho_{d}\in\mathbb{R}_{+}$ are ADMM parameters.
+Optimizing $\mathbf{x}$}, ${\mathbf{z}}_{u}$ and ${\mathbf{z}}_{d}$ in eq.
+9 in turn at iteration $\tau$ lead to three linear systems:
+$\hskip-4.2679pt{\left({\mathbf{H}}^{\top}{\mathbf{H}}+\frac{\rho}{2}{\mathcal{L}}^{d}_{r}+\frac{\rho_{u}+\rho_{d}}{2}{\mathbf{I}}\right)}{\mathbf{x}}^{\tau+1}$ $={({\mathbf{L}}^{d}_{r})^{\top}\!\left(\frac{{\bm{\gamma}}^{\tau}}{2}+\frac{\rho}{2}{\boldsymbol{\phi}}^{\tau}\right)-\frac{{\bm{\gamma}}_{u}^{\tau}}{2}+\frac{\rho_{u}}{2}{\mathbf{z}}_{u}^{\tau}-\frac{{\bm{\gamma}}_{d}^{\tau}}{2}+\frac{\rho_{d}}{2}{\mathbf{z}}_{d}^{\tau}+{\mathbf{H}}^{\top}{\mathbf{y}}}$ (10)
+${\left(\mu_{u}{\mathbf{L}}^{u}+\frac{\rho_{u}}{2}{\mathbf{I}}\right)}{\mathbf{z}}_{u}^{\tau+1}$ $={\frac{{\bm{\gamma}}_{u}^{\tau}}{2}+\frac{\rho_{u}}{2}{\mathbf{x}}^{\tau+1}}$ (11)
+${\left(\mu_{d,2}{\mathcal{L}}^{d}_{r}+\frac{\rho_{d}}{2}{\mathbf{I}}\right)}{\mathbf{z}}_{d}^{\tau+1}$ $={\frac{{\bm{\gamma}}_{d}^{\tau}}{2}+\frac{\rho_{d}}{2}{\mathbf{x}}^{\tau+1}}.$ (12)
+Thus, instead of solving eq.
+7 for ${\mathbf{x}}^{\tau+1}$ directly at iteration $\tau$, ${\mathbf{x}}^{\tau+1}$, ${\mathbf{z}}_{u}^{\tau+1}$ and ${\mathbf{z}}_{d}^{\tau+1}$ are optimized in turn via eq.
+10 through eq.
+12 using CG.
 This splitting induces more network parameters for data-driven learning after algorithm unrolling (see Section 4.1) towards better performance, and affords us spectral filter interpretations of the derived linear systems.
-Interpretation: To interpret solution ${\mathbf{z}}_{u}^{\tau+1}$ spectrally, we rewrite (11) as
-${\mathbf{z}}_{u}^{\tau+1}=\left(\frac{2\mu_{u}}{\rho_{u}}{\mathbf{L}}^{u}+{\mathbf{I}}\right)^{-1}\left(\frac{{\boldsymbol{\gamma}}_{u}}{\rho_{u}}+{\mathbf{x}}^{\tau+1}\right)={\mathbf{V}}\mathrm{diag}\left(\left\{\frac{1}{1+\frac{2\mu_{u}}{\rho_{u}}\lambda_{k}}\right\}\right){\mathbf{V}}^{\top}\left(\frac{{\boldsymbol{\gamma}}_{u}}{\rho_{u}}+{\mathbf{x}}^{\tau+1}\right)$ (13)
+Interpretation: To interpret solution ${\mathbf{z}}_{u}^{\tau+1}$ spectrally, we rewrite eq. 11 as
+$\hskip-2.84526pt{\mathbf{z}}_{u}^{\tau+1}=\left(\frac{2\mu_{u}}{\rho_{u}}{\mathbf{L}}^{u}+{\mathbf{I}}\right)^{-1}\left(\frac{{\bm{\gamma}}_{u}}{\rho_{u}}+{\mathbf{x}}^{\tau+1}\right)={\mathbf{V}}\mathrm{diag}\left(\left\{\frac{1}{1+\frac{2\mu_{u}}{\rho_{u}}\lambda_{k}}\right\}\right){\mathbf{V}}^{\top}\left(\frac{{\bm{\gamma}}_{u}}{\rho_{u}}+{\mathbf{x}}^{\tau+1}\right)$ (13)
 where ${\mathbf{L}}^{u}={\mathbf{V}}\mathrm{diag}(\{\lambda_{k}\}){\mathbf{V}}^{\top}$ is the eigen-decomposition.
-Thus, using frequencies defined by eigen-pairs $\{(\lambda_{k},{\mathbf{v}}_{k})\}$ of ${\mathbf{L}}^{u}$, ${\mathbf{z}}_{u}^{\tau+1}$ is the low-pass filter output of ${\mathbf{x}}^{\tau+1}$ offset by $\frac{{\boldsymbol{\gamma}}_{u}}{\rho_{u}}$, with frequency response $f(\lambda)=(1+\frac{2\mu_{u}}{\rho_{u}}\lambda)^{-1}$.
-As done for ${\mathbf{z}}_{u}^{\tau+1}$, we can rewrite ${\mathbf{z}}_{d}^{\tau+1}$ in (12) as
-${\mathbf{z}}_{d}^{\tau+1}$ $={\mathbf{U}}\mathrm{diag}\left(\left\{\frac{1}{1+\frac{2\mu_{d,2}}{\rho_{d}}\xi_{k}}\right\}\right){\mathbf{U}}^{\top}\left(\frac{{\boldsymbol{\gamma}}_{d}}{\rho_{d}}+{\mathbf{x}}^{\tau+1}\right)$ (14)
+Thus, using frequencies defined by eigen-pairs $\{(\lambda_{k},{\mathbf{v}}_{k})\}$ of ${\mathbf{L}}^{u}$, ${\mathbf{z}}_{u}^{\tau+1}$ is the low-pass filter output of ${\mathbf{x}}^{\tau+1}$ offset by $\frac{{\bm{\gamma}}_{u}}{\rho_{u}}$, with frequency response $f(\lambda)=(1+\frac{2\mu_{u}}{\rho_{u}}\lambda)^{-1}$.
+As done for ${\mathbf{z}}_{u}^{\tau+1}$, we can rewrite ${\mathbf{z}}_{d}^{\tau+1}$ in eq. 12 as
+${\mathbf{z}}_{d}^{\tau+1}$ $={\mathbf{U}}\mathrm{diag}\left(\left\{\frac{1}{1+\frac{2\mu_{d,2}}{\rho_{d}}\xi_{k}}\right\}\right){\mathbf{U}}^{\top}\left(\frac{{\bm{\gamma}}_{d}}{\rho_{d}}+{\mathbf{x}}^{\tau+1}\right)$ (14)
 where ${\mathcal{L}}^{d}_{r}={\mathbf{U}}\mathrm{diag}(\{\xi_{k}\}){\mathbf{U}}^{\top}$ is the eigen-decomposition.
-Analogously, ${\mathbf{z}}_{d}^{\tau+1}$ is the low-pass filter output of ${\mathbf{x}}^{\tau+1}$ offset by $\frac{{\boldsymbol{\gamma}}_{d}}{\rho_{d}}$, with frequency response $f(\xi)=(1+\frac{2\mu_{d,2}}{\rho_{d}}\xi)^{-1}$.
-To interpret solution ${\mathbf{x}}^{\tau+1}$ in (10) as a low-pass filter output, see Appendix B.
+Analogously, ${\mathbf{z}}_{d}^{\tau+1}$ is the low-pass filter output of ${\mathbf{x}}^{\tau+1}$ offset by $\frac{{\bm{\gamma}}_{d}}{\rho_{d}}$, with frequency response $f(\xi)=(1+\frac{2\mu_{d,2}}{\rho_{d}}\xi)^{-1}$.
+To interpret solution ${\mathbf{x}}^{\tau+1}$ in eq. 10 as a low-pass filter output, see Appendix C.
 3.4.2 Minimizing ${\boldsymbol{\phi}}^{\tau+1}$
-Fixing ${\mathbf{x}}^{\tau+1}$ at iteration $\tau$, optimizing ${\boldsymbol{\phi}}^{\tau+1}$ in (6) means
-${\boldsymbol{\phi}}^{\tau+1}=$ $\arg\min_{{\boldsymbol{\phi}}}g({\boldsymbol{\phi}})=\mu_{d,1}\|{\boldsymbol{\phi}}\|_{1}+({\boldsymbol{\gamma}}^{\tau})^{\top}({\boldsymbol{\phi}}-{\mathbf{L}}^{d}_{r}{\mathbf{x}}^{\tau+1})+\frac{\rho}{2}\|{\boldsymbol{\phi}}-{\mathbf{L}}^{d}_{r}{\mathbf{x}}^{\tau+1}\|^{2}_{2}.$ (15)
-(15) can be computed element-wise as follows (see Appendix C for a derivation):
+Fixing ${\mathbf{x}}^{\tau+1}$ at iteration $\tau$, optimizing ${\boldsymbol{\phi}}^{\tau+1}$ in eq. 6 means
+${\boldsymbol{\phi}}^{\tau+1}=$ $\arg\min_{{\boldsymbol{\phi}}}g({\boldsymbol{\phi}})=\mu_{d,1}\|{\boldsymbol{\phi}}\|_{1}+({\bm{\gamma}}^{\tau})^{\top}({\boldsymbol{\phi}}-{\mathbf{L}}^{d}_{r}{\mathbf{x}}^{\tau+1})+\frac{\rho}{2}\|{\boldsymbol{\phi}}-{\mathbf{L}}^{d}_{r}{\mathbf{x}}^{\tau+1}\|^{2}_{2}.$ (15)
+We can compute eq. 15 element-wise as follows (see Appendix D for a derivation):
 $\delta$ $=({\mathbf{L}}^{d}_{r})_{i}{\mathbf{x}}^{\tau+1}-\rho^{-1}\gamma_{i}^{\tau}$ (16)
 $\phi_{i}^{\tau+1}$ $=\text{sign}(\delta)\cdot\max(|\delta|-\rho^{-1}\mu_{d,1},0).$
-Interpretation: Soft-thresholding in (16) to compute $\phi_{i}^{\tau+1}$ attenuates the $i$-th entry of ${\mathbf{L}}^{d}_{r}{\mathbf{x}}^{\tau+1}$.
-Given that graph Laplacians—second difference matrices on graphs—are high-pass filters [3], we interpret ${\mathbf{L}}^{d}_{r}$ as the high-pass channel of a two-channel filterbank [30], where the corresponding low-pass channel is ${\mathbf{I}}-{\mathbf{L}}^{d}_{r}={\mathbf{W}}^{d}_{r}$.
-Using (16) for processing means that the low-pass channel ${\mathbf{W}}^{d}_{r}$ is unused and thus preserved.
+Interpretation: Soft-thresholding in eq.
+16 to compute $\phi_{i}^{\tau+1}$ attenuates the $i$-th entry of ${\mathbf{L}}^{d}_{r}{\mathbf{x}}^{\tau+1}$.
+Given that graph Laplacians—second difference matrices on graphs—are high-pass filters (Ortega et al., 2018), we interpret ${\mathbf{L}}^{d}_{r}$ as the high-pass channel of a two-channel filterbank (Vetterli & Kovacevic, 2013), where the corresponding low-pass channel is ${\mathbf{I}}-{\mathbf{L}}^{d}_{r}={\mathbf{W}}^{d}_{r}$.
+Using eq.
+16 for processing means that the low-pass channel ${\mathbf{W}}^{d}_{r}$ is unused and thus preserved.
 For example, the constant (low-frequency) signal $\mathbf{1}$} passes through the low-pass channel undisturbed, i.e., ${\mathbf{1}}={\mathbf{W}}^{d}_{r}{\mathbf{1}}$, while it is filtered out completely by the high-pass channel, i.e., ${\mathbf{0}}={\mathbf{L}}^{d}_{r}{\mathbf{1}}$.
-Thus, attenuation of the high-pass channel means (16) is a low-pass filter.
+Thus, attenuation of the high-pass channel means eq.
+16 is a low-pass filter.
 3.4.3 Updating Lagrange Multipliers
-For objective (6), we follow standard ADMM procedure to update Lagrange multiplier ${\boldsymbol{\gamma}}^{\tau+1}$ after each computation of ${\mathbf{x}}^{\tau+1}$ and ${\boldsymbol{\phi}}^{\tau+1}$:
-${\boldsymbol{\gamma}}^{\tau+1}$ $={\boldsymbol{\gamma}}^{\tau}+\rho({\boldsymbol{\phi}}^{\tau+1}-{\mathbf{L}}^{d}_{r}{\mathbf{x}}^{\tau+1}).$ (17)
-Splitting the $\ell_{2}$-norm terms means that additional multipliers ${\boldsymbol{\gamma}}_{u}^{\tau+1}$ and ${\boldsymbol{\gamma}}_{d}^{\tau+1}$ in objective (9) must also be updated after each computation of ${\mathbf{x}}^{\tau+1}$, ${\mathbf{z}}_{u}^{\tau+1}$, ${\mathbf{z}}_{d}^{\tau+1}$, and ${\boldsymbol{\phi}}^{\tau+1}$:
-${\boldsymbol{\gamma}}_{u}^{\tau+1}={\boldsymbol{\gamma}}_{u}^{\tau}+\rho_{u}({\mathbf{x}}^{\tau+1}-{\mathbf{z}}_{u}^{\tau+1}),~{}~{}~{}~{}~{}~{}~{}~{}~{}{\boldsymbol{\gamma}}_{d}^{\tau+1}={\boldsymbol{\gamma}}_{d}^{\tau}+\rho_{d}({\mathbf{x}}^{\tau+1}-{\mathbf{z}}_{d}^{\tau+1}).$ (18)
+For objective eq. 6, we follow standard ADMM procedure to update Lagrange multiplier ${\bm{\gamma}}^{\tau+1}$ after each computation of ${\mathbf{x}}^{\tau+1}$ and ${\boldsymbol{\phi}}^{\tau+1}$:
+${\bm{\gamma}}^{\tau+1}$ $={\bm{\gamma}}^{\tau}+\rho({\boldsymbol{\phi}}^{\tau+1}-{\mathbf{L}}^{d}_{r}{\mathbf{x}}^{\tau+1}).$ (17)
+Splitting the $\ell_{2}$-norm terms means that additional multipliers ${\bm{\gamma}}_{u}^{\tau+1}$ and ${\bm{\gamma}}_{d}^{\tau+1}$ in objective eq.
+9 must also be updated after each computation of ${\mathbf{x}}^{\tau+1}$, ${\mathbf{z}}_{u}^{\tau+1}$, ${\mathbf{z}}_{d}^{\tau+1}$, and ${\boldsymbol{\phi}}^{\tau+1}$:
+${\bm{\gamma}}_{u}^{\tau+1}={\bm{\gamma}}_{u}^{\tau}+\rho_{u}({\mathbf{x}}^{\tau+1}-{\mathbf{z}}_{u}^{\tau+1}),~~~~~~~~~{\bm{\gamma}}_{d}^{\tau+1}={\bm{\gamma}}_{d}^{\tau}+\rho_{d}({\mathbf{x}}^{\tau+1}-{\mathbf{z}}_{d}^{\tau+1}).$ (18)
 The ADMM loop is repeated till ${\mathbf{x}}^{\tau+1}$ and ${\boldsymbol{\phi}}^{\tau+1}$ converge.
@@
 We unroll our iterative ADMM algorithm to an interpretable neural net.
-The crux to our unrolled neural net is two periodically inserted graph learning modules for undirected and directed edge weight computation that are akin to the self-attention mechanism in classical transformers [19].
+The crux to our network is two periodically inserted graph learning modules for undirected and directed edge weight computation that are akin to the classical self-attention mechanism in transformers (Bahdanau et al., 2014).
 4.1 ADMM Algorithm Unrolling
 We implement each iteration of our algorithm in Section 3.4 as a neural layer.
-Specifically, each ADMM iteration is implemented as four sub-layers in sequence: solving linear systems (10) for ${\mathbf{x}}^{\tau+1}$, (11) for ${\mathbf{z}}_{u}^{\tau+1}$, and (12) for ${\mathbf{z}}_{d}^{\tau+1}$ via CG, and computing ${\boldsymbol{\phi}}^{\tau+1}$ via (16).
-CG is itself an iterative descent algorithm, where parameters $\alpha$ and $\beta$ corresponding to step size and momentum can be tuned end-to-end (see Appendix D for detailed implementation).
+Specifically, each ADMM iteration is implemented as four sub-layers in sequence: solving linear systems eq.
+10 for ${\mathbf{x}}^{\tau+1}$, eq.
+11 for ${\mathbf{z}}_{u}^{\tau+1}$, and eq.
+12 for ${\mathbf{z}}_{d}^{\tau+1}$ via CG, and computing ${\boldsymbol{\phi}}^{\tau+1}$ via eq.
+16.
+CG is itself an iterative descent algorithm, where parameters $\alpha$ and $\beta$ corresponding to step size and momentum can be tuned end-to-end (see Appendix E for detailed implementation).
 Weight parameters $\mu_{u},\mu_{d,2},\mu_{d,1}$ for prior terms and ADMM parameters $\rho,\rho_{u},\rho_{d}$ are also tuned per layer.
-Multipliers ${\boldsymbol{\gamma}}^{\tau+1},{\boldsymbol{\gamma}}_{u}^{\tau+1},{\boldsymbol{\gamma}}_{d}^{\tau+1}$ are then updated via (17) and (18) to complete one layer.
-Parameters learned during back-propagation for each ADMM block $b$ are denoted by ${\boldsymbol{\Theta}}_{b}$.
+Multipliers ${\bm{\gamma}}^{\tau+1},{\bm{\gamma}}_{u}^{\tau+1},{\bm{\gamma}}_{d}^{\tau+1}$ are then updated via eq.
+17 and eq.
+18 to complete one layer.
+Parameters learned during back-propagation for each ADMM block $b$ are denoted by ${\bm{\Theta}}_{b}$.
 See Fig.
@@
 4.2 Self-Attention Operator in Transformer
-We first review the basic self-attention mechanism: a scaled dot product following by a softmax operation [19].
+We first review the basic self-attention mechanism: a scaled dot product following by a softmax operation (Bahdanau et al., 2014).
 Denote by ${\mathbf{x}}_{i}\in\mathbb{R}^{E}$ an embedding for token $i$.
@@
 Undirected Graph Learning: We learn an undirected graph ${\mathcal{G}}^{u}$ in a module $\texttt{UGL}_{b}$ at block $b$.
-For each node $i$, we compute a low-dimensional feature vector ${\mathbf{f}}_{i}^{u}=F^{u}({\mathbf{e}}_{i})\in\mathbb{R}^{K}$, where ${\mathbf{e}}_{i}\in\mathbb{R}^{E}$ is a vector of signal values, position embeddings in time and Laplacian embeddings of the physical graph as done in Feng and Tassiulas [9], and $F^{u}(\cdot)$ is a chosen learned nonlinear feature function $F^{u}:\mathbb{R}^{E}\mapsto\mathbb{R}^{K}$, e.g., a shallow GNN (see Appendix E.3 for details).
+For each node $i$, we compute a low-dimensional feature vector ${\mathbf{f}}_{i}^{u}=F^{u}({\mathbf{e}}_{i})\in\mathbb{R}^{K}$, where ${\mathbf{e}}_{i}\in\mathbb{R}^{E}$ is a vector of signal values, position embeddings in time and Laplacian embeddings of the physical graph as done in Feng & Tassiulas (2022), and $F^{u}(\cdot)$ is a chosen learned nonlinear feature function $F^{u}:\mathbb{R}^{E}\mapsto\mathbb{R}^{K}$, e.g., a shallow GNN (see Appendix F.3 for details).
 As done in Thuc et al.
-[18], the feature distance between nodes $i$ and $j$ is computed as the Mahalanobis distance $\text{d}^{u}(i,j)$:
+(2024), the feature distance between nodes $i$ and $j$ is computed as the Mahalanobis distance $\text{d}^{u}(i,j)$:
 $\text{d}^{u}(i,j)=({\mathbf{f}}_{i}^{u}-{\mathbf{f}}_{j}^{u})^{\top}{\mathbf{M}}({\mathbf{f}}_{i}^{u}-{\mathbf{f}}_{j}^{u})$ (21)
 where ${\mathbf{M}}\succeq 0$ is a symmetric PSD metric matrix of dimension $\mathbb{R}^{K\times K}$.
-Distance in (21) is non-negative and symmetric, i.e., $\text{d}^{u}(i,j)=\text{d}^{u}(j,i)$.
+Distance in eq.
+21 is non-negative and symmetric, i.e., $\text{d}^{u}(i,j)=\text{d}^{u}(j,i)$.
 Weight $w^{u}_{i,j}$ of undirected edge $(i,j)\in{\mathcal{E}}^{u}$ is computed as
@@
 where the denominator summing over 1-hop neighborhood ${\mathcal{N}}_{i}$ is inserted for normalization.
-Remark: Interpreting distance $\text{d}^{u}(i,j)$ as $-e(i,j)$, edge weights $w^{u}_{i,j}$’s in (22) are essentially attention weights $a_{i,j}$’s in (19), and thus the undirected graph learning module constitutes a self-attention mechanism, albeit requiring fewer parameters, since parameters for function $F^{u}(\cdot)$ and metric matrix ${\mathbf{M}}$ can be much smaller than large and dense query and key matrices, ${\mathbf{Q}}$ and ${\mathbf{K}}$.
-Further, there is no value matrix ${\mathbf{V}}$ required, since the output signal ${\mathbf{x}}^{\tau+1}$ is computed through a set of low-pass filters—(10), (11), (12) and (16)—derived from our optimization of objective (4).
+Remark: Interpreting distance $\text{d}^{u}(i,j)$ as e(i,j)-e(i,j), edge weights $w^{u}_{i,j}$’s in eq.
+22 are essentially attention weights $a_{i,j}$’s in eq.
+19, and thus the undirected graph learning module constitutes a self-attention mechanism, albeit requiring fewer parameters, since parameters for function $F^{u}(\cdot)$ and metric matrix $\mathbf{M}$} can be much smaller than large and dense query and key matrices, $\mathbf{Q}$} and $\mathbf{K}$}.
+Further, no value matrix $\mathbf{V}$} is required, since output signal ${\mathbf{x}}^{\tau+1}$ is computed through a set of low-pass filters—eq.
+10, eq.
+11, eq.
+12 and eq.
+16—derived from our optimization of objective eq.
+4.
 Directed Graph Learning: Similarly, we learn a directed graph ${\mathcal{G}}^{d}$ in a module $\texttt{DGL}_{b}$ at block $b$.
 For each node $i$, we compute a feature vector ${\mathbf{f}}^{d}_{i}=F^{d}({\mathbf{e}}_{i})\in\mathbb{R}^{K}$ using feature function $F^{d}(\cdot)$.
-Like (21), we compute Mahalanobis distance $\text{d}^{d}(i,j)$ between nodes $i$ and $j$, given feature vectors ${\mathbf{f}}_{i}^{d}$ and ${\mathbf{f}}_{j}^{d}$, using a learned PSD metric matrix ${\mathbf{P}}$.
-Weight $w_{j,i}^{d}$ of directed edge $[i,j]\in{\mathcal{E}}^{d}$ is then computed using exponential functions and normalization, as done in (22) for an undirected edge.
-Parameters learned via back-propagation by the two graph learning modules at block $b$—parameters of $F^{u}(\cdot)$, $F^{d}(\cdot)$, ${\mathbf{M}}$ and ${\mathbf{P}}$—are denoted by ${\boldsymbol{\Phi}}_{b}$.
+Like eq.
+21, the Mahalanobis distance $\text{d}^{d}(i,j)$ between nodes $i$ and $j$ is computed from ${\mathbf{f}}_{i}^{d}$ and ${\mathbf{f}}_{j}^{d}$ with a learned PSD metric $\mathbf{P}$}.
+Weight $w_{j,i}^{d}$ of directed edge $i,j]\in{\mathcal{E}}^{d}$ is then computed using exponentials and normalization, as done in eq.
+22 for an undirected edge.
+Parameters learned via back-propagation by the two graph learning modules at block bb—parameters of $F^{u}(\cdot)$, $F^{d}(\cdot)$, $\mathbf{M}$} and $\mathbf{P}$}—are denoted by ${\bm{\Phi}}_{b}$.
 5 Experiments
 5.1 Experimental Setup
-We evaluate our model’s performance on traffic speed datasets METR-LA, PEMS-BAY [31], and traffic flow dataset PEMS03 [32], each of which contains traffic data with a sampling interval of 5 minutes in a local region, together with the real connections and distances (or travel cost) between sensors.
+We evaluate our model’s performance on commonly used traffic speed dataset METR-LA (Li et al., 2017) and traffic flow dataset PEMS03 (Guo et al., 2022), each of which contains city-size traffic data with a sampling interval of 5 minutes in a local region, together with the real connections and distances (or travel cost) between sensors.
 We reduce the size of the original dataset by sub-sampling 1/3 of all data uniformly, and split it into training, validation, and test sets by 6:2:2.
-We predict the traffic speed/flow in the following 30 / 60 minutes (6 / 12 steps) from the previously observed 60 minutes (12 steps) for all datasets.
+We predict the traffic speed/flow in the following 30 / 60 /120 minutes (6 / 12 /24 steps) from the previously observed 60 minutes (12 steps) for all datasets.
 We restrict our model to learn a sparse graph based on the real road connections and a limited time window to model only the local spatial/temporal influence.
-For undirected graphs ${\mathcal{G}}^{u}$, we connect each node with its $k$ nearest neighbors (NNs).
-We construct 5 ADMM blocks, each contains 25 ADMM layers, and insert a graph learning module before each ADMM block.
+For undirected graphs ${\mathcal{G}}^{u}$, we connect each node to its $k$ nearest neighbors (NNs).
+We construct 5 ADMM blocks, each contains 25 ADMM layers, and insert a pair of graph learning modules before each ADMM block.
 We set feature dimensions $K=6$, and stack vertically 4 graph learning modules to learn in parallel as an implementation of the multi-head attention mechanism.
 We minimize the Huber loss with $\delta=1$ for the entire reconstructed sequence and the groundtruths.
-See Appendix E for more details.
+See Appendix F for more details.
 5.2 Experimental Results
-Table 1: Comparison of RMSE / MAE / MAPE(%) metrics of our lightweight transformer to baseline models on 30 / 60-minute forecasting in PeMS03, METR-LA and PEMS-BAY datasets.
-We use boldface for the smallest error and underline the next 3 smallest errors.
-Dataset & PEMS03 (358 nodes, 547 edges) METR-LA (207 nodes, 1315 edges) PEMS-BAY (325 nodes, 2090 edges)
-Horizon 30 minutes 60 minutes 30 minutes 60 minutes 30 minutes 60 minutes
-VAR [33] 28.07 / 16.53 / 17.49 30.54 / 18.31 / 19.61 10.72 / 5.55 / 11.29 12.59 / 6.99 / 13.46 3.81 / 2.08 / 4.46 3.98 / 2.56 / 5.52
-STGCN [34] 28.06 / 18.24 / 17.76 37.31 / 24.29 / 23.00 11.97 / 5.43 / 13.45 14.77 / 6.86 / 17.51 4.32 / 2.65 / 5.21 7.39 / 4.26 / 8.25
-STSGCN [35] 26.41 / 16.75 / 16.86 30.62 / 19.35 / 19.15 10.25 / 4.05 / 9.18 12.65 / 5.18 / 11.48 3.43 / 1.54 / 3.28 4.50 / 1.93 / 4.42
-ST-Wave [36] 25.57 / 15.11 / 15.04 28.65 / 16.81 / 19.24 10.81 / 4.11 / 9.16 13.24 / 5.33 / 11.32 3.53 / 1.49 / 3.18 4.30 / 1.82 / 4.33
-ASTTN [9] 26.84 / 17.04 / 21.46 27.24 / 16.43 / 15.63 11.49 / 4.80 / 10.97 11.10 / 5.08 / 12.51 5.07 / 2.23 / 4.65 4.82 / 2.13 / 4.40
-PDFormer [22] 23.71 / 15.05 / 18.16 27.16 / 17.26 / 21.21 10.21 / 3.89 / 8.50 12.17 / 4.81 / 10.92 3.18 / 1.41 / 2.99 4.55 / 2.00 / 4.88
-STAEformer [37] 30.22 / 18.85 / 26.62 38.36 / 23.68 / 29.21 10.16 / 3.73 / 8.25 12.58 / 4.79 / 10.08 4.07 / 1.75 / 3.95 5.28 / 2.28 / 5.59
-Ours 26.10 / 17.03 / 18.85 27.67 / 17.46 / 17.72 10.35 / 4.22 / 9.70 12.37 / 5.27 / 11.91 3.78 / 1.70 / 4.08 4.60 / 2.03 / 4.79
-Table 2: Comparison of the number of parameters between our model and baseline deep learning models for 60-minute forecasting in PEMS03 dataset.
-Model Ours STGCN STSGCN ST-Wave ASTTN PDFormer STAEformer
-Params # 34,096 321,941 3,496,212 882,558 161,635 531,165 1,403,940
-We train each run of our model on NVIDIA GeForce RTX 3090.
-We evaluate our unrolling model against model-based method VAR [33], GNN-based models STGCN [^4] STGCN predicts only one time step at a time.
-We train the model to predict only the next step, and predict the entire sequence recursively.
-[34] and STSGCN [35], GAT-based models ST-Wave [36] and ASTTN [9], and transformer-based models PDFormer [22] and STAEformer [37] with 70 epochs of training on the reduced dataset with the Root Mean Squared Error (RMSE), Mean Absolute Error (MAE), and Mean Absolute Percentage Error (MAPE) in the predicted 6 / 12 steps.
-Table 1 shows the performance of our models and baselines in traffic forecasting, and Table 2 compares the number of parameters.
+Table 1: Comparison of categorized baselines and model parameters for 60-minute forecasting on the PEMS03 dataset against ours.
+Category Selected Models Params #
+Model-based VAR (Reinsel, 2003) -
+[…]
+GATs GMAN (Zheng et al., 2020), ST-Wave (Fang et al., 2023) 210K, 883K
+Transformers PDFormer (Jiang et al., 2023), STAEformer (Liu et al., 2023) 531K, 1404K
+Adaptive-graph models Graph WaveNet (GWN) (Wu et al., 2019), AGCRN (Bai et al., 2020) 277K, 749K
+Simple linear models STID (Shao et al., 2022), SimpleTM (Chen et al., 2025) 123K, 540K
+Mixed-graph unrolling Ours 34K
+  • $\dagger$
+    STGCN predicts only one time step at a time. We train the model to predict only the next step, and predict the entire sequence recursively.
+Table 2: Comparison of RMSE / MAE / MAPE(%) metrics of our lightweight transformer to baseline models on 30 / 60-minute forecasting in PeMS03 and METR-LA datasets.
+We use boldface for the smallest error, color the 2nd and 3nd small error in blue, and underline the next 2 smallest errors.
+Dataset & PEMS03 (358 nodes, 547 edges) METR-LA (207 nodes, 1315 edges)
+Horizon 30 minutes 60 minutes 120 minutes 30 minutes 60 minutes 120 minutes
+VAR 28.07 / 16.53 / 17.49 30.54 / 18.31 / 19.61 36.65 / 22.40 / 24.64 10.72 / 5.55 / 11.29 12.59 / 6.99 / 13.46 14.83 / 8.90 / 16.54
+[…]
+STSGCN 26.41 / 16.75 / 16.86 30.62 / 19.35 / 19.15 38.04 / 23.86 / 23.62 10.25 / 4.05 / 9.18 12.65 / 5.18 / 11.48 15.74 / 6.84 / 15.33
+[…]
+ST-Wave 25.57 / 15.11 / 15.04 28.65 / 16.81 / 19.24 29.88 / 17.11 / 17.71 10.81 / 4.11 / 9.16 13.24 / 5.33 / 11.32 23.18 / 11.22 / 12.71
+[…]
+STAEformer 30.22 / 18.85 / 26.62 38.36 / 23.68 / 29.21 48.72 / 31.96 / 44.71 10.16 / 3.73 / 8.25 12.58 / 4.79 / 10.08 14.63 / 5.82 / 11.87
+[…]
+AGCRN 27.40 / 15.19 / 14.35 29.90 / 16.76 / 15.32 32.79 / 18.72 / 17.03 10.11 / 3.82 / 8.61 12.56 / 5.00 / 11.25 14.77 / 6.33 / 14.05
+[…]
+SimpleTM 23.75 / 15.13 / 15.59 25.56 / 15.97 / 15.49 30.58 / 18.73 / 18.18 8.76 / 4.12 / 7.63 11.96 / 5.49 / 9.65 15.44 / 7.64 / 12.27
+Ours 26.10 / 17.03 / 18.85 27.67 / 17.46 / 17.72 34.40 / 21.16 / 23.47 10.35 / 4.22 / 9.70 12.34 / 5.18 / 11.80 15.19 / 7.56 / 16.12
+We train each run of our model on an entire NVIDIA GeForce RTX 3090.
+We evaluate our unrolling model against the baselines shown in Table 1.
+Each model is trained for 70 epochs on the reduced dataset, and the metrics are the Root Mean Squared Error (RMSE), Mean Absolute Error (MAE), and Mean Absolute Percentage Error (MAPE) in the predicted 6 / 12 / 24 steps.
+Table 2 shows the performance of our models and baselines in traffic forecasting, and the last column of Table 1 compares the number of parameters.
 We observe that our models reach comparable performance to most baseline models, while employ drastically fewer parameters (6.4% of transformer-based PDFormer).
-Among all the models, ST-Wave generally achieves the best performance, which is attributed to an application-specific usage of wavelet transforms, but requires even more parameters than transformers.
-The largest model STSGCN fails to achieve satisfying performance for PEMS03, where the real connections are highly sparse.
-STAEformer, the largest transformer, also overfits heavily in the reduced PEMS03 and PEMS-BAY datasets.
-We also observe significant oscillations in validation loss during training in ASTTN due to the adaptive attention maps.
-In summary, none of the DL-based baseline schemes unambiguously achieves the best performance across the three datasets, while all require significantly more parameters than our models.
-5.3 Ablation Study
-Next, we discuss the effectiveness of the new directed graph priors, DGLR and DGTV, as well as the benefit of directed-graph modeling of temporal relations.
-First, we separately test the effects of DGLR and DGTV by removing the prior and unrolling the ADMM algorithm.
-We denote these models as “w/o DGTV” and “w/o DGLR”.
-To test the benefit of the directed-graph modeling, we change the temporal graph ${\mathcal{G}}^{d}$ into a fully undirected graph ${\mathcal{G}}^{n}$ while maintaining the connections, and replace the directed graph priors with a GLR prior ${\mathbf{x}}^{\top}{\mathbf{L}}^{n}{\mathbf{x}}$ with the same coefficient $\mu_{n}=\mu_{d,2}$.
-We denote the undirected-time unrolling model as “UT”.
-Detailed derivations of the ADMM algorithm for all the above ablation studies are in Appendix F.
-Table 3: Performance comparison in RMSE / MAE / MAPE(%) of the ablation studies of directed-graph modeling and directed graph priors in 60-minute forecasting with PEMS03 and METR-LA.
-Dataset Mixed-Graph UT w/o DGTV w/o DGLR
-PEMS03 27.67 / 17.46 / 17.72 27.52 / 17.87 / 18.82 27.78 / 17.85 / 17.90 30.89 / 20.02 / 21.10
-METR-LA 12.37 / 5.27 / 11.91 12.51 / 5.42 /12.11 12.36 / 5.40 / 12.31 12.41 / 5.35 / 12.20
-As is summarized in Table 3, removing the DGLR or DGTV term leads to noticeable drops in performance, which underscores their importance as directed graph priors.
-Moreover, our model outperforms the fully-undirected-graph-based model, demonstrating that sequential relationships—such as time—are better modeled as directed graphs.
-For further ablation studies in the number of nearest neighbors $k$, size of time window $W$, number of ADMM blocks and layers, and the splitting-term strategy in (8), see Appendix G.
+Among all the models, ST-Wave and SimpleTM generally achieve the best performance.
+However, ST-Wave relies on an application-specific usage of wavelet transforms, and SimpleTM employs linear models with no inductive bias, which is parameter-heavy (Battaglia et al., 2018); both models require even more parameters than PDFormer.
+Our unrolled model also generally outperforms the multi-attention model GMAN.
+Though both our model and GMAN leverage local graph neighborhoods when computing attention, we dramatically reduce parameters by replacing classical self-attention mechanisms in transformers with unrolled ADMM iterations with no observable degradation in performance.
+The largest GCN model, STSGCN, fails to achieve a satisfying performance for PEMS03, where the real connections are highly sparse.
+STAEformer, the largest transformer, also overfits heavily in the reduced PEMS03 datasets.
+In summary, none of the DL-based baseline schemes unambiguously achieves the best performance across the two datasets, while all require significantly more parameters than our models.
+In contrast, our model captures the complex and evolving spatial-temporal relationship accurately using a mixed graph, demonstrating consistent performance for all predicting horizons; notably, the performance gap to the best model decreases as the horizon grows.
+Further experiments on data efficiency are presented in Appendix G, and a comprehensive ablation study is provided in Appendix H.
 6 Conclusion
-To forecast city traffic, we unroll a mixed-graph optimization algorithm into a lightweight transformer-like neural net, where an undirected graph models spatial correlations and a directed graph models temporal relationships.
-We show that the two graph learning modules are akin to the classical self-attention mechanism.
+To capture complex node-to-node relations in spatial-temporal data, we unroll a mixed-graph optimization algorithm into a lightweight transformer-like neural net, where an undirected graph models spatial correlations and a directed graph models temporal relationships.
+We show that the two graph learning modules play the role of self-attention, meaning that our unrolled network is a transformer.
 We design new $\ell_{2}$ and $\ell_{1}$-norm regularizers to quantify and promote signal smoothness on directed graphs.
-We interpret the derived processing operations from a constructed ADMM algorithm as low-pass filters.
+We interpret the derived processing operations from a designed ADMM algorithm as low-pass filters.
 Experiments demonstrate competitive prediction performance at drastically reduced parameter counts.
 One limitation of our approach is that our computed Mahalanobis distances $\text{d}^{u}(i,j)$ and $\text{d}^{d}(i,j)$ for respective undirected and directed graphs are non-negative (due to learned PSD metric matrices $\mathbf{M}$} and $\mathbf{P}$}), whereas affinity $e(i,j)$ in traditional self-attention can be negative.
-We will study extensions to signed distances as future work.
-Appendix A Proof of Theorem 3.1
+We will study extensions to signed distances and more complex graph modeling as future work.
+Appendix A Related Works
+A.1 Lightweight Learning Models
+Given the importance of parameter reduction for increasingly large deep learning models, there exist various studies towards this common goal across application domains.
+For large language models (LLMs), smaller transformer models have been achieved via low-rank assumption (Hu et al., 2022) or parameter quantization (Frantar et al., 2023).
+For speech enhancement, Oostermeijer et al.
+(2021) proposes local causal self-attention, while Zhao & Madhu (2025) proposes parameter sharing across spectral and time dimensions.
+For low-light image enhancement, Brateanu et al.
+(2025) processes the Y-channel via convolution, pooling, and multi-head self-attention, separately from the U- and V-channels.
+For hyperspectral image classification, Sun et al.
+(2024) proposes a novel multiscale 3D-2D mixed CNN design.
+For cervical cancer detection, Pacal (2024) substitutes larger MBConv blocks in the MaxViT architecture with smaller ConvNeXtv2 blocks, and MLP blocks with GRN-based MLPs.
+Similar architecture miniaturization techniques have been proposed for edge devices (Mishra & Gupta, 2024) and resource-constrained environments (Liu et al., 2024).
+For traffic prediction, there are practical use cases where parameter reduction is important.
+For example, drivers in a city predict local traffic conditions (in an area-specific neighborhood graph) on their memory-constrained mobile devices every five minutes to compute optimal routes to their destinations, based on updated congestion information shared in their local neighborhoods, instead of sending individual requests to the cloud and overwhelming the central server.
+Thuc et al.
+(2024) achieves a lightweight model for image interpolation in a more principled manner: first design a linear-time optimization algorithm minimizing a chosen undirected graph smoothness prior (such as graph Laplacian regularizer (GLR) (Pang & Cheung, 2017) or graph total variation (GTV) (Bai et al., 2019)), then unroll its iterations into neural layers to compose a “white-box” transformer-like neural net for data-driven parameter learning.
+This algorithm unrolling approach results in a lightweight and mathematically interpretable network.
+Our current work differs from Thuc et al.
+(2024) in that we model the more complex node-to-node relationships in spatial-temporal data using a mixed graph: i) an undirected graph ${\mathcal{G}}^{u}$ to capture spatial correlations among geographically near stations, and ii) a directed graph ${\mathcal{G}}^{d}$ to capture sequential relationships along the temporal dimension.
+In particular, the notion of “frequencies” for directed graphs is not well understood in the graph signal processing (GSP) community.
+We are the first to define this notion and promote low-pass signal reconstruction via our designed $\ell_{2}$-norm directed graph Laplacian regularizer (DGLR) variational term in eq.
+2 and $\ell_{1}$-norm directed graph total variation (DGTV) variational term in eq.
+3, both derived from the graph shift operator (GSO) (Chen et al., 2015).
+To the best of our knowledge, we are the first define and promote smooth signals on a directed graph in a data-driven transformer setting.
+A.2 Deep Models for Traffic Forecasting
+Traffic forecasting problems contain spatial and temporal correlations between stations and time steps; thus, Graph Neural Network (GNN) is naturally employed in early research.
+Early research, such as STGCN (Yu et al., 2018), STSGCN (Song et al., 2020) and GGRNN (Ruiz et al., 2020), employs Graph Convolutional Networks (GCNs) to capture spatial relations and uses 1D convolutions and RNNs to process temporal relationships.
+ASTGNN (Ta et al., 2022), Graph WaveNet (Wu et al., 2019) and AGCRN (Bai et al., 2019) explored adaptive graph settings to capture potential non-local signal relations over the graph, trading in interpretability for improved performance.
+Graph Attention Networks (GATs) learn graph weights from embedded signals to enable flexible and data-targeted graph learning, typical examples include GMAN (Zheng et al., 2020) and ST-Wave (Fang et al., 2023).
+All GNNs above focus only on the spatial correlations, while the product graph (PG) is an intuitive model that describes local spatial-temporal relations at the same time.
+Einizade et al.
+(2024) presents a continuous product graph neural network for spatial-temporal forecasting, which achieves superior performance to the previous most-popular graph methods.
+Transformers extend the powerful self-attention mechanism to all spatial-temporal data points, which enhances model performance with an extra-large number of parameters.
+Popular transformer-based models for traffic forecasting employ two transformers to process spatial and temporal dimensions separately (Xu et al., 2020; Chen et al., 2022; Huo et al., 2023), while state-of-the-art transformers introduce adaptive local attentions (Liu et al., 2023), time-delay (Jiang et al., 2023), or pre-training (Gao et al., 2024).
+In recent years, simple linear models have been proposed to challenge the effectiveness and efficiency of heavy transformers.
+DLinear (Zeng et al., 2023) simply separates the input signals as trends and residuals, and employs 2 simple linear layers shared across nodes for prediction, but beats some of the transformer baselines in long-term forecasting.
+Other representative simple linear models include STID (Shao et al., 2022) and SimpleTM Chen et al.
+(2025).
+As shown in Table 1, we select two most representative and well-performing models from each category as baselines for comparison.
+Appendix B Proof of Theorem 3.1
 We prove Theorem 3.1 by construction as follows.
@@
 ${\mathcal{L}}^{d}_{r}={\mathbf{L}}^{u}$ $=\left[\begin{array}[]{cccccc}1&-1&0&\ldots&&0\\ -1&2&-1&0&\ldots&0\\ 0&-1&2&-1&0\ldots&0\\ \vdots&&\ddots&\ddots&\ddots&\\ 0&\ldots&0&-1&2&-1\\ 0&\ldots&&0&-1&1\end{array}\right].$ (40)
-As an illustration, consider a 4-node directed line graph ${\mathcal{G}}^{d}$ with edge weights $1$; see Fig.
+As an illustration, consider a 4-node directed line graph ${\mathcal{G}}^{d}$ with edge weights 11; see Fig.
 2(a) for an illustration.
@@
 ${\mathbf{W}}^{d}_{r}=\left[\begin{array}[]{cccc}1&0&0&0\\ 1&0&0&0\\ 0&1&0&0\\ 0&0&1&0\end{array}\right],~{\mathbf{L}}^{d}_{r}=\left[\begin{array}[]{cccc}0&0&0&0\\ -1&1&0&0\\ 0&-1&1&0\\ 0&0&-1&1\end{array}\right],{\mathcal{L}}^{d}_{r}=\left[\begin{array}[]{cccc}1&-1&0&0\\ -1&2&-1&0\\ 0&-1&2&-1\\ 0&0&-1&1\end{array}\right].$ (53)
-Appendix B Interpreting Linear System (10)
+Appendix C Interpreting Linear System eq. 10
 For notation simplicity, we first define
-${\mathbf{r}}^{\tau}\triangleq\small({\mathbf{L}}^{d}_{r})^{\top}(\frac{{\boldsymbol{\gamma}}^{\tau}}{2}+\frac{\rho}{2}{\boldsymbol{\phi}}^{\tau})-\frac{{\boldsymbol{\gamma}}_{u}^{\tau}}{2}+\frac{\rho_{u}}{2}{\mathbf{z}}_{u}^{\tau}-\frac{{\boldsymbol{\gamma}}_{d}^{\tau}}{2}+\frac{\rho_{d}}{2}{\mathbf{z}}_{d}^{\tau}.$ (54)
-Solution ${\mathbf{x}}^{\tau+1}$ to linear system (10) can now be written as
+${\mathbf{r}}^{\tau}\triangleq({\mathbf{L}}^{d}_{r})^{\top}(\frac{{\bm{\gamma}}^{\tau}}{2}+\frac{\rho}{2}{\boldsymbol{\phi}}^{\tau})-\frac{{\bm{\gamma}}_{u}^{\tau}}{2}+\frac{\rho_{u}}{2}{\mathbf{z}}_{u}^{\tau}-\frac{{\bm{\gamma}}_{d}^{\tau}}{2}+\frac{\rho_{d}}{2}{\mathbf{z}}_{d}^{\tau}.$ (54)
+Solution ${\mathbf{x}}^{\tau+1}$ to linear system eq. 10 can now be written as
 ${\mathbf{x}}^{\tau+1}$ $=\left({\mathbf{H}}^{\top}{\mathbf{H}}+\frac{\rho}{2}{\mathcal{L}}^{d}_{r}+\frac{\rho_{u}+\rho_{d}}{2}{\mathbf{I}}\right)^{-1}({\mathbf{r}}^{\tau}+{\mathbf{H}}^{\top}{\mathbf{y}}).$ (55)
-Define ~$\tilde{{\mathcal{L}}}_{r}^{d}\triangleq{\mathbf{H}}^{\top}{\mathbf{H}}+\frac{\rho}{2}{\mathcal{L}}^{d}_{r}$.
-Note that ${\mathbf{H}}^{\top}{\mathbf{H}}$ is a diagonal matrix with zeros and ones (corresponding to chosen sample indices) along its diagonal, and thus ~$\tilde{{\mathcal{L}}}^{d}_{r}$ is a scaled variant of ${\mathcal{L}}^{d}_{r}$ with the addition of selected self-loops of weight $1$.
-Given that ${\mathcal{L}}^{d}_{r}=({\mathbf{L}}^{d}_{r})^{\top}{\mathbf{L}}^{d}_{r}$ is real and symmetric, ~$\tilde{{\mathcal{L}}}^{d}_{r}$ is also real and symmetric.
-By eigen-decomposing ~$\tilde{{\mathcal{L}}}^{d}_{r}=\tilde{{\mathbf{U}}}\mathrm{diag}(\{\tilde{\xi}_{k}\})\tilde{{\mathbf{U}}}^{\top}$, we can rewrite ${\mathbf{x}}^{\tau+1}$ as
+Define $\tilde{{\mathcal{L}}}_{r}^{d}\triangleq{\mathbf{H}}^{\top}{\mathbf{H}}+\frac{\rho}{2}{\mathcal{L}}^{d}_{r}$.
+Note that ${\mathbf{H}}^{\top}{\mathbf{H}}$ is a diagonal matrix with zeros and ones (corresponding to chosen sample indices) along its diagonal, and thus $\tilde{{\mathcal{L}}}^{d}_{r}$ is a scaled variant of ${\mathcal{L}}^{d}_{r}$ with the addition of selected self-loops of weight $1$.
+Given that ${\mathcal{L}}^{d}_{r}=({\mathbf{L}}^{d}_{r})^{\top}{\mathbf{L}}^{d}_{r}$ is real and symmetric, $\tilde{{\mathcal{L}}}^{d}_{r}$ is also real and symmetric.
+By eigen-decomposing $\tilde{{\mathcal{L}}}^{d}_{r}=\tilde{{\mathbf{U}}}\mathrm{diag}(\{\tilde{\xi}_{k}\})\tilde{{\mathbf{U}}}^{\top}$, we can rewrite ${\mathbf{x}}^{\tau+1}$ as
 ${\mathbf{x}}^{\tau+1}$ $=\tilde{{\mathbf{U}}}\mathrm{diag}\left(\left\{\frac{1}{\frac{\rho_{u}+\rho_{d}}{2}+\tilde{\xi}_{k}}\right\}\right)\tilde{{\mathbf{U}}}^{\top}\left({\mathbf{r}}^{\tau}+{\mathbf{H}}^{\top}{\mathbf{y}}\right)$ (56)
 which demonstrates that ${\mathbf{x}}^{\tau+1}$ is a low-pass filter output of up-sampled ${\mathbf{H}}^{\top}{\mathbf{y}}$ (with bias ${\mathbf{r}}^{\tau}$).
-Appendix C Deriving Solution to Optimization (15)
-The optimization for ${\boldsymbol{\phi}}$ in objective (15) can be performed entry-by-entry. Specifically, for the $i$-th entry $\phi_{i}$:
+Appendix D Deriving Solution to Optimization eq. 15
+The optimization for $\boldsymbol{\phi}$} in objective eq. 15 can be performed entry-by-entry. Specifically, for the $i$-th entry $\phi_{i}$:
 $\min_{\phi_{i}}~~g(\phi_{i})=\mu_{d,1}|\phi_{i}|+\gamma_{i}^{\tau}\left(\phi_{i}-({\mathbf{L}}^{d}_{r})_{i}{\mathbf{x}}^{\tau+1}\right)+\frac{\rho}{2}\left(\phi_{i}-({\mathbf{L}}^{d}_{r})_{i}{\mathbf{x}}^{\tau+1}\right)^{2}$ (57)
@@
 $\phi_{i}^{*}$ $=\text{sign}(\delta)\cdot\max(|\delta|-\rho^{-1}\mu_{d,1},\,0)$ (60)
-Alternative solution with subgradients: Denote $g_{1}({\boldsymbol{\phi}})=\mu_{d,1}\|{\boldsymbol{\phi}}\|_{1},g_{2}({\boldsymbol{\phi}})=({\boldsymbol{\gamma}}^{\tau})^{\top}({\boldsymbol{\phi}}-{\mathbf{L}}^{d}_{r}{\mathbf{x}}^{\tau+1})+\frac{\rho}{2}\|{\boldsymbol{\phi}}-{\mathbf{L}}^{d}_{r}{\mathbf{x}}^{\tau+1}\|_{2}^{2}$, the subgradient of $g({\boldsymbol{\phi}})$ is
-$\partial g({\boldsymbol{\phi}})=\{{\mathbf{g}}+\nabla g_{2}({\boldsymbol{\phi}})|{\mathbf{g}}\in\partial g_{1}({\boldsymbol{\phi}})\}=\{\mu_{d,1}\sigma+{\boldsymbol{\gamma}}^{\tau}+\rho({\boldsymbol{\phi}}-{\mathbf{L}}^{d}_{r}{\mathbf{x}}^{\tau+1})|\sigma\in\partial(\|{\boldsymbol{\phi}}\|_{1})\},$ (61)
+Alternative solution with subgradients: Denote $g_{1}({\boldsymbol{\phi}})=\mu_{d,1}\|{\boldsymbol{\phi}}\|_{1},g_{2}({\boldsymbol{\phi}})=({\bm{\gamma}}^{\tau})^{\top}({\boldsymbol{\phi}}-{\mathbf{L}}^{d}_{r}{\mathbf{x}}^{\tau+1})+\frac{\rho}{2}\|{\boldsymbol{\phi}}-{\mathbf{L}}^{d}_{r}{\mathbf{x}}^{\tau+1}\|_{2}^{2}$, the subgradient of $g({\boldsymbol{\phi}})$ is
+$\partial g({\boldsymbol{\phi}})=\{{\mathbf{g}}+\nabla g_{2}({\boldsymbol{\phi}})|{\mathbf{g}}\in\partial g_{1}({\boldsymbol{\phi}})\}=\{\mu_{d,1}\sigma+{\bm{\gamma}}^{\tau}+\rho({\boldsymbol{\phi}}-{\mathbf{L}}^{d}_{r}{\mathbf{x}}^{\tau+1})|\sigma\in\partial(\|{\boldsymbol{\phi}}\|_{1})\},$ (61)
 where
@@
 which gives
-$\phi_{i}^{*}=\left\{\begin{aligned} &({\mathbf{L}}^{d}_{r})_{i}{\mathbf{x}}-\frac{\gamma_{i}^{\tau}}{\rho}-\frac{\mu_{d,1}}{\rho},&&({\mathbf{L}}^{d}_{r})_{i}{\mathbf{x}}-\frac{\gamma_{i}^{\tau}}{\rho}>\frac{\mu_{d,1}}{\rho}\\ &0,&&-\frac{\mu_{d,1}}{\rho}\leq({\mathbf{L}}^{d}_{r})_{i}{\mathbf{x}}-\frac{\gamma_{i}^{\tau}}{\rho}\leq\frac{\mu_{d,1}}{\rho}\\ &({\mathbf{L}}^{d}_{r})_{i}{\mathbf{x}}-\frac{\gamma_{i}^{\tau}}{\rho}+\frac{\mu_{d,1}}{\rho},&&({\mathbf{L}}^{d}_{r})_{i}{\mathbf{x}}-\frac{\gamma_{i}^{\tau}}{\rho}<-\frac{\mu_{d,1}}{\rho}\\ \end{aligned}\right..$ (64)
-Thus the solution of (15) is
+$\phi_{i}^{*}=\left\{\begin{aligned} &({\mathbf{L}}^{d}_{r})_{i}{\mathbf{x}}-\frac{\gamma_{i}^{\tau}}{\rho}-\frac{\mu_{d,1}}{\rho},&&({\mathbf{L}}^{d}_{r})_{i}{\mathbf{x}}-\frac{\gamma_{i}^{\tau}}{\rho}>\frac{\mu_{d,1}}{\rho}\\ &0,&&-\frac{\mu_{d,1}}{\rho}\leq({\mathbf{L}}^{d}_{r})_{i}{\mathbf{x}}-\frac{\gamma_{i}^{\tau}}{\rho}\leq\frac{\mu_{d,1}}{\rho}\\ &({\mathbf{L}}^{d}_{r})_{i}{\mathbf{x}}-\frac{\gamma_{i}^{\tau}}{\rho}+\frac{\mu_{d,1}}{\rho},&&({\mathbf{L}}^{d}_{r})_{i}{\mathbf{x}}-\frac{\gamma_{i}^{\tau}}{\rho}<-\frac{\mu_{d,1}}{\rho}\\ \end{aligned}\right.$.
+(64)
+Thus the solution of eq. 15 is
 ${\boldsymbol{\phi}}^{\tau+1}={\boldsymbol{\phi}}^{*}=\text{soft}_{\frac{\mu_{d,1}}{\rho}}\left({\mathbf{L}}^{d}_{r}{\mathbf{x}}-\frac{\gamma_{i}^{\tau}}{\rho}\right)=\text{sign}\left({\mathbf{L}}^{d}_{r}{\mathbf{x}}-\frac{\gamma_{i}^{\tau}}{\rho}\right)\max\left(\left|{\mathbf{L}}^{d}_{r}{\mathbf{x}}-\frac{\gamma_{i}^{\tau}}{\rho}\right|-\frac{\mu_{d,1}}{\rho},0\right).$ (65)
-Appendix D Parameterizing Conjugated Gradient Method in Solving Linear Equations
+Appendix E Parameterizing Conjugated Gradient Method in Solving Linear Equations
 A linear system ${\mathbf{A}}{\mathbf{x}}={\mathbf{b}}$ where $\mathbf{A}$} is a PD matrix can be solved iteratively by the conjugated gradient (CG) method, as is described in Algorithm 1.
@@
 2:Approximate solution $\mathbf{x}$} such that $\|{\mathbf{A}}{\mathbf{x}}-{\mathbf{b}}\|\leq\epsilon$, where $\epsilon$ is the tolerance of convergence
-3:${\mathbf{r}}_{0}\leftarrow{\mathbf{b}}-{\mathbf{A}}{\mathbf{x}}_{0}$triangleright Initial residual
-4:${\mathbf{p}}_{0}\leftarrow{\mathbf{r}}_{0}$
-5:$k\leftarrow 0$
+[…]
+[…]
+[…]
 6:while $\|{\mathbf{A}}{\mathbf{x}}_{k}-{\mathbf{b}}\|>\epsilon$ do
@@
 8:  ${\mathbf{x}}_{k+1}\leftarrow{\mathbf{x}}_{k}+\alpha_{k}{\mathbf{p}}_{k}$
-9:     ${\mathbf{r}}_{k+1}\leftarrow{\mathbf{r}}_{k}-\alpha_{k}{\mathbf{A}}{\mathbf{p}}_{k}$triangleright Dual residual
+9:  ${\mathbf{r}}_{k+1}\leftarrow{\mathbf{r}}_{k}-\alpha_{k}{\mathbf{A}}{\mathbf{p}}_{k}$ $\triangleright$ Dual residual
 10:  $\boxed{\beta_{k}\leftarrow\dfrac{{\mathbf{r}}_{k+1}^{\top}{\mathbf{r}}_{k+1}}{{\mathbf{r}}_{k}^{\top}{\mathbf{r}}_{k}}}$
@@
 14:return ${\mathbf{x}}_{k+1}$
-Appendix E Model Setup
-E.1 ADMM Blocks Setup
+Appendix F Model Setup
+F.1 ADMM Blocks Setup
 We initialize $\mu_{u},\mu_{d,1},\mu_{d,2}$ with 3, and the $\rho,\rho_{u},\rho_{d}$ with $\sqrt{N/(T+S+1)}$, where $N$ is the number of sensors and $T+S+1$ is the full length of the recovered signal (both the observed part and the part to be predicted).
 Those parameters are trained every ADMM layer.
-We initialize the CGD parameters (${\boldsymbol{\alpha}}$, ${\boldsymbol{\beta}}$)s as 0.08, and tune them for every CGD iteration.
-We clamp each element of the CGD step size ${\boldsymbol{\alpha}}$ to [0, 0.8] for stability.
-Each element of the momentum term coefficient ${\boldsymbol{\beta}}$ is clamped to be non-negative.
-E.2 Graph Learning Modules Setup
+We initialize the CGD parameters ($\bm{\alpha}$}, $\bm{\beta}$})s as 0.08, and tune them for every CGD iteration.
+We clamp each element of the CGD step size $\bm{\alpha}$} to [0, 0.8] for stability.
+Each element of the momentum term coefficient $\bm{\beta}$} is clamped to be non-negative.
+F.2 Graph Learning Modules Setup
 We define multiple PD matrices $\mathbf{P}$}s and $\mathbf{M}$}s in our unrolling model due to the different impact strengths at different times.
@@
 The output signals of all 4 channels are integrated to one by a linear layer at the end of each ADMM block.
-E.3 Feature Extractors Setup
+F.3 Feature Extractors Setup
 Input Laplacian embeddings Position embeddings Concatenate embedded input
@@
 Input Embedding $x$ Spatial GraphSAGE $\alpha(\cdot)$ Temporal History Aggregation $\alpha(\cdot)$features
-(b) A layer of eature extractor.
+(b) A layer of feature extractor.
 Figure 4: Framework of our feature extractor.
@@
 We detailedly introduce our design of the feature extractor described in Section 4.3. Figure 4(b) shows the overall structure of our feature extractors.
-E.3.1 Input Embeddings
+F.3.1 Input Embeddings
 We generate non-parametric spatial and temporal embeddings for each node in our mixed spatio-temporal graph, see Figure 4(a).
@@
 where $\text{Var}(\text{d}^{S})$ denotes the variance of the physical distances.
-We use Laplacian embedding supposed by Dwivedi and Bresson [38], which uses the eigenvectors of the $k^{S}$-smallest non-zero eigenvalues of the Laplacian matrix ${\mathbf{L}}^{s}={\mathbf{D}}^{s}-{\mathbf{W}}^{s}$ to generate $k$-dimensional embeddings.
+We use Laplacian embedding supposed by Dwivedi & Bresson (2020), which uses the eigenvectors of the $k^{S}$-smallest non-zero eigenvalues of the Laplacian matrix ${\mathbf{L}}^{s}={\mathbf{D}}^{s}-{\mathbf{W}}^{s}$ to generate $k$-dimensional embeddings.
 Our experiments set $k^{S}=5$.
-For temporal embeddings, we use the sine-cosine position embeddings[39]:
+For temporal embeddings, we use the sine-cosine position embeddings (Vaswani et al., 2018):
 $e[t,2i]=$ $\sin(t/10000^{i})$
 $e[t,2i+1]=$ $\cos(t/10000^{i}),i=0,1,\dots,\left\lfloor\frac{k^{T}}{2}\right\rfloor-1.$ (67)
-We follow Feng and Tassiulas [9] to use real-time-stamp, time-in-day and day-in-week embeddings of dimensions 10, 6, 4.
+We follow Feng & Tassiulas (2022) to use real-time-stamp, time-in-day and day-in-week embeddings of dimensions 10, 6, 4.
 We directly concatenate these spatial and temporal embedding to our signals ${\mathbf{x}}^{t}_{j}$ for node $j$ at time $t$ as ${\mathbf{e}}_{j}^{t}=[{\mathbf{x}}^{t}_{j};{\mathbf{e}}^{S}_{j};{\mathbf{e}}^{T}_{t}]$ for feature extractors, where ${\mathbf{e}}^{S}_{j}$ denote the spatial Laplacian embeddings of node $j$ of the physical graph, and ${\mathbf{e}}^{T}_{t}$ denotes the temporal position embeddings of time step $t$.
-E.3.2 Feature Extraction from embeddings
-We use a variant of GraphSAGE [40] to aggregate spatial features and introduce Temporal History Aggregation (THA) to aggregate spatial features.
+F.3.2 Feature Extraction from embeddings
+We use a variant of GraphSAGE (Hamilton et al., 2017) to aggregate spatial features and introduce Temporal History Aggregation (THA) to aggregate spatial features.
 The GraphSAGE module aggregates the inputs of each node’s $k$-NNs with a simple and uniformed linear layer to generate $K$-dimension spatial features for each node of each time step.
 The THA module aggregates the embedded inputs of past $W$ time steps for each node with a uniformed linear layer, where $W$ is the size of the time window.
-For both GraphSAGE and THA modules, we pad zeros to inputs as placeholders, and use Swish [41] as acitvation functions with $\beta=0.8$.
-E.4 Constructing Graphs ${\mathbf{W}}^{u,0},{\mathbf{W}}^{d,0}$ for the First ADMM Block
+For both GraphSAGE and THA modules, we pad zeros to inputs as placeholders, and use Swish (Ramachandran et al., 2017) as acitvation functions with $\beta=0.8$.
+F.4 Constructing Graphs ${\mathbf{W}}^{u,0},{\mathbf{W}}^{d,0}$ for the First ADMM Block
 Our ADMM block recovers the full signal $\mathbf{x}$} from observed signal $\mathbf{y}$} with mixed graph ${\mathcal{G}}^{u},{\mathcal{G}}^{d}$ constructed from embedded full signal ${\mathbf{e}}=[{\mathbf{x}};{\mathbf{e}}^{S};{\mathbf{e}}^{T}]$.
 Thus, for the first block in ADMM, we need an initial extrapolation for the signals to predict.
-We propose a simple module to make an initial guess of the signals: we modify the feature extractor described in Appendix E.3.2 to generate intermediate features from the embedded input of the observed part.
+We propose a simple module to make an initial guess of the signals: we modify the feature extractor described in Appendix F.3.2 to generate intermediate features from the embedded input of the observed part.
 Then we combine the feature and temporal dimensions, and use a simple linear layer with Swish activation to generate the initially extrapolated signal.
@@
 Furthermore, low-parameter and simple feature engineering ensures that the effectiveness of our model is attributed to the unrolled blocks and graph learning modules, rather than to the trivial feature engineering parts.
-E.5 Training Setup
+F.5 Training Setup
 We preprocess the dataset by standardizing the entire time series for each node in space, but minimize a loss between the real groundtruth signal and the rescaled whole output signal.
@@
 Table 3: Training setup of graph construction and batch size for all traffic datasets.
-Dataset PeMS03 PeMS04 METR-LA PEMS-BAY
-$k$ 4 6 6 6
-$W$ 6 6 6 4
-Batch size 12 12 16 12
-Appendix F Optimization Formulation and ADMM Algorithm in Ablation Studies
-F.1 Optimization & Algorithm without DGTV Term
-By removing the DGTV term $\|{\mathbf{L}}^{d}_{r}{\mathbf{x}}\|_{1}$ in (4), and introducing the auxiliary variables ${\mathbf{z}}_{u},{\mathbf{z}}_{d}$ as (9), we rewrite the unconstrained version of the optimization by augmented Lagrangian method as
-$\min_{{\mathbf{x}},{\mathbf{z}}_{u},{\mathbf{z}}_{d}}$ $\|{\mathbf{y}}-{\mathbf{H}}{\mathbf{x}}\|_{2}^{2}+\mu_{u}{\mathbf{z}}_{u}^{\top}{\mathbf{L}}^{u}{\mathbf{z}}_{u}+\mu_{d,2}{\mathbf{z}}_{d}^{\top}{\mathcal{L}}^{d}_{r}{\mathbf{z}}_{d}+({\boldsymbol{\gamma}}_{u}^{\tau})^{\top}({\mathbf{x}}-{\mathbf{z}}_{u})+\frac{\rho_{u}}{2}\|{\mathbf{x}}-{\mathbf{z}}_{u}\|_{2}^{2}$
-$+({\boldsymbol{\gamma}}_{d}^{\tau})^{\top}({\mathbf{x}}-{\mathbf{z}}_{d})+\frac{\rho_{d}}{2}\|{\mathbf{x}}-{\mathbf{z}}_{d}\|_{2}^{2}.$ (68)
-We optimize ${\mathbf{x}},{\mathbf{z}}_{u}$ and ${\mathbf{z}}_{d}$ in turn at iteration $\tau$ as follows:
-$\left({\mathbf{H}}^{\top}{\mathbf{H}}+\frac{\rho_{u}+\rho_{d}}{2}{\mathbf{I}}\right){\mathbf{x}}^{\tau+1}=$ ${\mathbf{H}}^{\top}{\mathbf{y}}-\frac{{\boldsymbol{\gamma}}_{u}^{\tau}+{\boldsymbol{\gamma}}_{d}^{\tau}}{2}+\frac{\rho_{u}}{2}{\mathbf{z}}_{u}^{\tau}+\frac{\rho_{d}}{2}{\mathbf{z}}_{d}^{\tau},$ (69)
-$\left(\mu_{u}{\mathbf{L}}^{u}+\frac{\rho_{u}}{2}{\mathbf{I}}\right){\mathbf{z}}_{u}^{\tau+1}=$ $\frac{{\boldsymbol{\gamma}}_{u}^{\tau}}{2}+\frac{\rho_{u}}{2}{\mathbf{x}}^{\tau+1},$ (70)
-$\left(\mu_{d,2}{\mathcal{L}}^{d}_{r}+\frac{\rho_{d}}{2}{\mathbf{I}}\right){\mathbf{z}}_{d}^{\tau+1}=$ $\frac{{\boldsymbol{\gamma}}_{d}^{\tau}}{2}+\frac{\rho_{d}}{2}{\mathbf{x}}^{\tau+1}.$ (71)
-The updating of ${\boldsymbol{\gamma}}_{u}^{\tau},{\boldsymbol{\gamma}}_{d}^{\tau}$ follows (18).
-F.2 Optimization & Algorithm without DGLR Term
-By removing the DGLR term ${\mathbf{x}}{\mathcal{L}}^{d}_{r}{\mathbf{x}}$ in (4) and introduce only the auxiliary variables ${\boldsymbol{\phi}}$ in (5) and ${\mathbf{z}}_{u}$ in (9), we rewrite the unconstrained version of the optimization of ${\mathbf{x}}^{\tau+1}$ by augmented Lagrangian method as
-$\min_{x,{\mathbf{z}}_{u}}\|{\mathbf{y}}-{\mathbf{H}}{\mathbf{x}}\|_{2}^{2}+\mu_{u}{\mathbf{z}}_{u}^{\top}{\mathbf{L}}^{u}{\mathbf{z}}_{u}+({\boldsymbol{\gamma}}^{\tau})^{\top}({\boldsymbol{\phi}}^{\tau}-{\mathbf{L}}^{d}_{r}{\mathbf{x}})+\frac{\rho}{2}\|{\boldsymbol{\phi}}^{\tau}-{\mathbf{L}}^{d}_{r}{\mathbf{x}}\|_{2}^{2}+({\boldsymbol{\gamma}}_{u}^{\tau})^{\top}({\mathbf{x}}-{\mathbf{z}}_{u})+\frac{\rho_{u}}{2}\|{\mathbf{x}}-{\mathbf{z}}_{u}\|_{2}^{2}$ (72)
-We optimize ${\mathbf{x}}$ and ${\mathbf{z}}_{u}$ in turn at iteration $\tau$ as follows:
-$\left({\mathbf{H}}^{\top}{\mathbf{H}}+\frac{\rho}{2}{\mathcal{L}}^{d}_{r}+\frac{\rho_{u}}{2}{\mathbf{I}}\right){\mathbf{x}}^{\tau+1}$ $={\mathbf{H}}^{\top}{\mathbf{y}}+({\mathbf{L}}^{d}_{r})^{\top}(\frac{\gamma^{\tau}}{2}+\frac{\rho}{2}{\boldsymbol{\phi}}^{\tau})-\frac{{\boldsymbol{\gamma}}_{u}^{\tau}}{2}+\frac{\rho_{u}}{2}{\mathbf{z}}_{u}^{\tau}$ (73)
-$\left(\mu_{u}{\mathbf{L}}^{u}+\frac{\rho_{u}}{2}{\mathbf{I}}\right){\mathbf{z}}_{u}^{\tau+1}$ $=\frac{{\boldsymbol{\gamma}}_{u}^{\tau}}{2}+\frac{\rho_{u}}{2}{\mathbf{x}}^{\tau+1}$ (74)
-The optimization of ${\boldsymbol{\phi}}^{\tau+1}$ is the same as (15), and the solution is the same as (16). The updating of ${\boldsymbol{\gamma}},{\boldsymbol{\gamma}}_{u}$ follows (17) and (18).
-F.3 Algorithm Unrolling with an Undirected Temporal Graph
-In this section, we are changing the directed temporal graph ${\mathcal{G}}^{d}$ into an undirected graph while maintaining the connections.
-Denote the undirected temporal graph as ${\mathcal{G}}^{n}$.
-We introduce the GLR operator ${\mathbf{x}}^{\top}{\mathbf{L}}^{n}{\mathbf{x}}$ for ${\mathcal{G}}^{n}$ on signal ${\mathbf{x}}$ where ${\mathbf{L}}^{n}$ is the normalized Laplacian matrix of ${\mathcal{G}}^{n}$.
-Thus the optimization problem is as follows:
-$\min_{x}\|{\mathbf{y}}-{\mathbf{H}}{\mathbf{x}}\|_{2}^{2}+\mu_{u}{\mathbf{x}}^{\top}{\mathbf{L}}^{u}{\mathbf{x}}+\mu_{n}{\mathbf{x}}^{T}{\mathbf{L}}^{n}{\mathbf{x}}$ (75)
-where $\mu_{u},\mu_{n}\in\mathbb{R}_{+}$ are weight parameters for the two GLR priors.
-The opitimization is similar to that in Appendix F.1.
-By a similar approach which introduces auxiliary variables ${\mathbf{z}}_{u}={\mathbf{x}},{\mathbf{z}}_{n}={\mathbf{x}}$, we optimize ${\mathbf{x}},{\mathbf{z}}_{u},{\mathbf{z}}_{n}$ as follows:
-$\left({\mathbf{H}}^{\top}{\mathbf{H}}+\frac{\rho_{u}+\rho_{n}}{2}{\mathbf{I}}\right){\mathbf{x}}^{\tau+1}=$ ${\mathbf{H}}^{\top}{\mathbf{y}}-\frac{{\boldsymbol{\gamma}}_{u}^{\tau}+{\boldsymbol{\gamma}}_{n}^{\tau}}{2}+\frac{\rho_{u}}{2}{\mathbf{z}}_{u}^{\tau}+\frac{\rho_{n}}{2}{\mathbf{z}}_{n}^{\tau},$ (76)
-$\left(\mu_{u}{\mathbf{L}}^{u}+\frac{\rho_{u}}{2}{\mathbf{I}}\right){\mathbf{z}}_{u}^{\tau+1}=$ $\frac{{\boldsymbol{\gamma}}_{u}^{\tau}}{2}+\frac{\rho_{u}}{2}{\mathbf{x}}^{\tau+1},$ (77)
-$\left(\mu_{n}{\mathbf{L}}^{n}+\frac{\rho_{n}}{2}{\mathbf{I}}\right){\mathbf{z}}_{n}^{\tau+1}=$ $\frac{{\boldsymbol{\gamma}}_{n}^{\tau}}{2}+\frac{\rho_{n}}{2}{\mathbf{x}}^{\tau+1}.$ (78)
-The updating of ${\boldsymbol{\gamma}}_{u}^{\tau},{\boldsymbol{\gamma}}_{n}^{\tau}$ is also similar to (18) as follows:
-${\boldsymbol{\gamma}}_{u}^{\tau+1}={\boldsymbol{\gamma}}_{u}^{\tau}+\rho_{u}({\mathbf{x}}^{\tau+1}-{\mathbf{z}}_{u}^{\tau+1}),\quad{\boldsymbol{\gamma}}_{n}^{\tau+1}={\boldsymbol{\gamma}}_{n}^{\tau}+\rho_{n}({\mathbf{x}}^{\tau+1}-{\mathbf{z}}_{n}^{\tau+1}).$ (79)
-The undirected temporal graph ${\mathcal{G}}^{n}$ is learned by an UGL described in Section 4.3, and is initialized similar to the spatial UGL described in Appendix E.2.
-The initialization of $\mu_{n},\rho_{n}$ and the CGD parameters in solving (78) follows the setup in Appendix E.1.
-Appendix G Further Ablation Studies
-G.1 Sensitivity of $k$, $W$ in Graph Construction
+Dataset PeMS03 PeMS04 METR-LA
+$k$ 4 6 6
+$W$ 6 6 6
+Batch size 12 12 16
+Appendix G Training Performance w.r.t. Training Data Cost
+Due to the parameter efficiency of our model, our unrolling model is also data-efficient under the same observation length.
+We trained our models and baselines with reduced training datasets by selecting sample windows from the sequence with larger strides of 5 and 10 in the PEMS03 dataset for 60-minute forecasting.
+The results below show that baseline models with large parameter counts suffer from steep performance degradation when the training data becomes scarce.
+In contrast, our model maintains stable performance under limited data conditions, demonstrating both lower data requirements and better generalizability.
+This makes it particularly suitable for deployment in data-scarce or resource-constrained environments.
+Table 4: Performance under different sampling strides (metrics: MAE / RMSE / MAPE(%)) on 60-minute forecasting for PEMS03 dataset.
+Model Stride = 3 (original) Stride = 5 Stride = 10
+Ours 27.67 / 17.46 / 17.72 30.42 / 18.60 / 20.05 30.65 / 19.02 / 19.67
+VAR 30.54 / 18.31 / 19.61 30.68 / 18.37 / 19.74 31.25 / 18.71 / 21.22
+STSGCN 30.62 / 19.35 / 19.15 32.37 / 20.67 / 20.66 33.22 / 21.30 / 21.58
+PDFormer 27.16 / 17.26 / 21.21 25.70 / 15.37 / 15.99 77.29 / 44.36 / 126.63
+AGCRN 29.90 / 16.76 / 15.32 39.25 / 20.60 / 16.95 62.58 / 109.27 / 30.72
+Appendix H Ablation Studies
+H.1 Ablation Studies for $\ell_{2}$- and $\ell_{1}$- norms in Direct Graph Modeling
+First, we discuss the effectiveness of the new directed graph priors, DGLR and DGTV, as well as the benefit of directed-graph modeling of temporal relations.
+We separately test the effects of DGLR and DGTV by removing the prior and unrolling the ADMM algorithm.
+We denote these models as “w/o DGTV” and “w/o DGLR”.
+To test the benefit of the directed-graph modeling, we change the temporal graph ${\mathcal{G}}^{d}$ into a fully undirected graph ${\mathcal{G}}^{n}$ while maintaining the connections, and replace the directed graph priors with a GLR prior ${\mathbf{x}}^{\top}{\mathbf{L}}^{n}{\mathbf{x}}$ with the same coefficient $\mu_{n}=\mu_{d,2}$.
+We denote the undirected-time unrolling model as “UT”.
+Detailed derivations of the ADMM algorithm for all the above ablation studies are in Appendix I.
+Table 5: Performance comparison in RMSE / MAE / MAPE(%) of the ablation studies of directed-graph modeling and directed graph priors in 60-minute forecasting with PEMS03 and METR-LA.
+Dataset Mixed-Graph UT w/o DGTV w/o DGLR
+PEMS03 27.67 / 17.46 / 17.72 27.52 / 17.87 / 18.82 27.78 / 17.85 / 17.90 30.89 / 20.02 / 21.10
+METR-LA 12.34 / 5.18 / 11.80 12.51 / 5.42 / 12.11 12.36 / 5.40 / 12.31 12.41 / 5.35 / 12.20
+As is summarized in Table 5, removing the DGLR or DGTV term leads to noticeable drops in performance, which underscores their importance as directed graph priors.
+Moreover, our model outperforms the fully-undirected-graph-based model, demonstrating that sequential relationships—such as time—are better modeled as directed graphs.
+H.2 Sensitivity of $k$, $W$ in Graph Construction
 We change the number of spatial nearest neighbors $k=4,6,8$ and time window size $W=4,6,8$ and present the results for PEMS03 and METR-LA dataset in Table 6.
@@
 6 6 28.98 17.33 18.23 6 4 12.44 5.40 11.85
-8 6 45.37 18.15 20.59 6 8 12.44 5.3525 11.96
-G.2 Sensitivity of Number of Blocks and Layers in the Unrolling Model
+8 6 45.37 18.15 20.59 6 8 12.44 5.35 11.96
+H.3 Sensitivity of Number of Blocks and Layers in the Unrolling Model
 We vary the number of blocks and ADMM layers for PEMS03 dataset and present the results in Table 7.
@@
 RMSE MAE MAPE(%) RMSE MAE MAPE(%)
- 5*  25* 27.67 17.46 17.72 12.37 5.27 11.91
+ 5*  25* 27.67 17.46 17.72 12.34 5.18 11.80
 5 20 27.95 17.32 18.02 12.53 5.35 12.15
+5 30 27.86 17.73 18.80 12.54 5.31 11.94
 4 25 27.77 17.63 18.61 12.45 5.33 12.00
-G.3 Ablation Study for the Splitting-$\ell_{2}$-term Strategy
-We also ran ablations of the splitting-term strategy by optimizing ${\mathbf{x}}^{\tau+1}$ using only (7).
-CG in solving (7) converges much slower than in solving (10)-(12), and the unrolled model can become numerically unstable in training, indicating that splitting the $\ell_{2}$ terms can stabilize the training process.
-Appendix H Error Bar for Our Model in Table 1
+6 25 27.84 17.54 18.61 12.37 5.27 11.90
+H.4 Ablation Study for the Splitting-$\ell_{2}$-term Strategy
+We also ran ablations of the splitting-term strategy by optimizing ${\mathbf{x}}^{\tau+1}$ using only eq.
+7.
+CG in solving eq.
+7 converges much slower than in solving eq.
+10-eq.
+12, and the unrolled model can become numerically unstable in training, indicating that splitting the $\ell_{2}$ terms can stabilize the training process.
+Appendix I Optimization Formulation and ADMM Algorithm in Ablation Studies
+I.1 Optimization & Algorithm without DGTV Term
+By removing the DGTV term $\|{\mathbf{L}}^{d}_{r}{\mathbf{x}}\|_{1}$ in eq.
+4, and introducing the auxiliary variables ${\mathbf{z}}_{u},{\mathbf{z}}_{d}$ as eq.
+9, we rewrite the unconstrained version of the optimization by augmented Lagrangian method as
+$\min_{{\mathbf{x}},{\mathbf{z}}_{u},{\mathbf{z}}_{d}}$ $\|{\mathbf{y}}-{\mathbf{H}}{\mathbf{x}}\|_{2}^{2}+\mu_{u}{\mathbf{z}}_{u}^{\top}{\mathbf{L}}^{u}{\mathbf{z}}_{u}+\mu_{d,2}{\mathbf{z}}_{d}^{\top}{\mathcal{L}}^{d}_{r}{\mathbf{z}}_{d}+({\bm{\gamma}}_{u}^{\tau})^{\top}({\mathbf{x}}-{\mathbf{z}}_{u})+\frac{\rho_{u}}{2}\|{\mathbf{x}}-{\mathbf{z}}_{u}\|_{2}^{2}$
+$+({\bm{\gamma}}_{d}^{\tau})^{\top}({\mathbf{x}}-{\mathbf{z}}_{d})+\frac{\rho_{d}}{2}\|{\mathbf{x}}-{\mathbf{z}}_{d}\|_{2}^{2}.$ (68)
+We optimize ${\mathbf{x}},{\mathbf{z}}_{u}$ and ${\mathbf{z}}_{d}$ in turn at iteration $\tau$ as follows:
+$\left({\mathbf{H}}^{\top}{\mathbf{H}}+\frac{\rho_{u}+\rho_{d}}{2}{\mathbf{I}}\right){\mathbf{x}}^{\tau+1}=$ ${\mathbf{H}}^{\top}{\mathbf{y}}-\frac{{\bm{\gamma}}_{u}^{\tau}+{\bm{\gamma}}_{d}^{\tau}}{2}+\frac{\rho_{u}}{2}{\mathbf{z}}_{u}^{\tau}+\frac{\rho_{d}}{2}{\mathbf{z}}_{d}^{\tau},$ (69)
+$\left(\mu_{u}{\mathbf{L}}^{u}+\frac{\rho_{u}}{2}{\mathbf{I}}\right){\mathbf{z}}_{u}^{\tau+1}=$ $\frac{{\bm{\gamma}}_{u}^{\tau}}{2}+\frac{\rho_{u}}{2}{\mathbf{x}}^{\tau+1},$ (70)
+$\left(\mu_{d,2}{\mathcal{L}}^{d}_{r}+\frac{\rho_{d}}{2}{\mathbf{I}}\right){\mathbf{z}}_{d}^{\tau+1}=$ $\frac{{\bm{\gamma}}_{d}^{\tau}}{2}+\frac{\rho_{d}}{2}{\mathbf{x}}^{\tau+1}.$ (71)
+The updating of ${\bm{\gamma}}_{u}^{\tau},{\bm{\gamma}}_{d}^{\tau}$ follows eq. 18.
+I.2 Optimization & Algorithm without DGLR Term
+By removing the DGLR term ${\mathbf{x}}{\mathcal{L}}^{d}_{r}{\mathbf{x}}$ in eq.
+4 and introduce only the auxiliary variables $\boldsymbol{\phi}$} in eq.
+5 and ${\mathbf{z}}_{u}$ in eq.
+9, we rewrite the unconstrained version of the optimization of ${\mathbf{x}}^{\tau+1}$ by augmented Lagrangian method as
+$\min_{x,{\mathbf{z}}_{u}}\|{\mathbf{y}}-{\mathbf{H}}{\mathbf{x}}\|_{2}^{2}+\mu_{u}{\mathbf{z}}_{u}^{\top}{\mathbf{L}}^{u}{\mathbf{z}}_{u}+({\bm{\gamma}}^{\tau})^{\top}({\boldsymbol{\phi}}^{\tau}-{\mathbf{L}}^{d}_{r}{\mathbf{x}})+\frac{\rho}{2}\|{\boldsymbol{\phi}}^{\tau}-{\mathbf{L}}^{d}_{r}{\mathbf{x}}\|_{2}^{2}+({\bm{\gamma}}_{u}^{\tau})^{\top}({\mathbf{x}}-{\mathbf{z}}_{u})+\frac{\rho_{u}}{2}\|{\mathbf{x}}-{\mathbf{z}}_{u}\|_{2}^{2}$ (72)
+We optimize $\mathbf{x}$} and ${\mathbf{z}}_{u}$ in turn at iteration $\tau$ as follows:
+$\left({\mathbf{H}}^{\top}{\mathbf{H}}+\frac{\rho}{2}{\mathcal{L}}^{d}_{r}+\frac{\rho_{u}}{2}{\mathbf{I}}\right){\mathbf{x}}^{\tau+1}$ $={\mathbf{H}}^{\top}{\mathbf{y}}+({\mathbf{L}}^{d}_{r})^{\top}(\frac{\gamma^{\tau}}{2}+\frac{\rho}{2}{\boldsymbol{\phi}}^{\tau})-\frac{{\bm{\gamma}}_{u}^{\tau}}{2}+\frac{\rho_{u}}{2}{\mathbf{z}}_{u}^{\tau}$ (73)
+$\left(\mu_{u}{\mathbf{L}}^{u}+\frac{\rho_{u}}{2}{\mathbf{I}}\right){\mathbf{z}}_{u}^{\tau+1}$ $=\frac{{\bm{\gamma}}_{u}^{\tau}}{2}+\frac{\rho_{u}}{2}{\mathbf{x}}^{\tau+1}$ (74)
+The optimization of ${\boldsymbol{\phi}}^{\tau+1}$ is the same as eq. 15, and the solution is the same as eq. 16. The updating of ${\bm{\gamma}},{\bm{\gamma}}_{u}$ follows eq. 17 and eq. 18.
+I.3 Algorithm Unrolling with an Undirected Temporal Graph
+In this section, we are changing the directed temporal graph ${\mathcal{G}}^{d}$ into an undirected graph while maintaining the connections.
+Denote the undirected temporal graph as ${\mathcal{G}}^{n}$.
+We introduce the GLR operator ${\mathbf{x}}^{\top}{\mathbf{L}}^{n}{\mathbf{x}}$ for ${\mathcal{G}}^{n}$ on signal $\mathbf{x}$} where ${\mathbf{L}}^{n}$ is the normalized Laplacian matrix of ${\mathcal{G}}^{n}$.
+Thus the optimization problem is as follows:
+$\min_{x}\|{\mathbf{y}}-{\mathbf{H}}{\mathbf{x}}\|_{2}^{2}+\mu_{u}{\mathbf{x}}^{\top}{\mathbf{L}}^{u}{\mathbf{x}}+\mu_{n}{\mathbf{x}}^{T}{\mathbf{L}}^{n}{\mathbf{x}}$ (75)
+where $\mu_{u},\mu_{n}\in\mathbb{R}_{+}$ are weight parameters for the two GLR priors.
+The opitimization is similar to that in Appendix I.1.
+By a similar approach which introduces auxiliary variables ${\mathbf{z}}_{u}={\mathbf{x}},{\mathbf{z}}_{n}={\mathbf{x}}$, we optimize ${\mathbf{x}},{\mathbf{z}}_{u},{\mathbf{z}}_{n}$ as follows:
+$\left({\mathbf{H}}^{\top}{\mathbf{H}}+\frac{\rho_{u}+\rho_{n}}{2}{\mathbf{I}}\right){\mathbf{x}}^{\tau+1}=$ ${\mathbf{H}}^{\top}{\mathbf{y}}-\frac{{\bm{\gamma}}_{u}^{\tau}+{\bm{\gamma}}_{n}^{\tau}}{2}+\frac{\rho_{u}}{2}{\mathbf{z}}_{u}^{\tau}+\frac{\rho_{n}}{2}{\mathbf{z}}_{n}^{\tau},$ (76)
+$\left(\mu_{u}{\mathbf{L}}^{u}+\frac{\rho_{u}}{2}{\mathbf{I}}\right){\mathbf{z}}_{u}^{\tau+1}=$ $\frac{{\bm{\gamma}}_{u}^{\tau}}{2}+\frac{\rho_{u}}{2}{\mathbf{x}}^{\tau+1},$ (77)
+$\left(\mu_{n}{\mathbf{L}}^{n}+\frac{\rho_{n}}{2}{\mathbf{I}}\right){\mathbf{z}}_{n}^{\tau+1}=$ $\frac{{\bm{\gamma}}_{n}^{\tau}}{2}+\frac{\rho_{n}}{2}{\mathbf{x}}^{\tau+1}.$ (78)
+The updating of ${\bm{\gamma}}_{u}^{\tau},{\bm{\gamma}}_{n}^{\tau}$ is also similar to eq. 18 as follows:
+${\bm{\gamma}}_{u}^{\tau+1}={\bm{\gamma}}_{u}^{\tau}+\rho_{u}({\mathbf{x}}^{\tau+1}-{\mathbf{z}}_{u}^{\tau+1}),\quad{\bm{\gamma}}_{n}^{\tau+1}={\bm{\gamma}}_{n}^{\tau}+\rho_{n}({\mathbf{x}}^{\tau+1}-{\mathbf{z}}_{n}^{\tau+1}).$ (79)
+The undirected temporal graph ${\mathcal{G}}^{n}$ is learned by an UGL described in Section 4.3, and is initialized similar to the spatial UGL described in Appendix F.2.
+The initialization of $\mu_{n},\rho_{n}$ and the CGD parameters in solving eq.
+78 follows the setup in Appendix F.1.
+Appendix J Error Bar for Our Model in Table 2
 […]
 Table 8: $2\sigma$ error bar of the performance metrics of our model in 60-minute traffic forecast in all three datasets.
-Dataset PeMS03 METR-LA PEMS-BAY
-$2\sigma_{\text{RMSE}}$ 0.05 0.04 0.01
-$2\sigma_{\text{MAE}}$ 0.02 0.02 0.01
-$2\sigma_{\text{MAPE(\%)}}$ 0.03 0.04 0.10
-Appendix I Reconstruction Error
-We perform the spatial/temporal forecast by recovering the entire time-vertex signal, which is unusual compared to other deep models.
-Table 8 shows that the reconstruction error is much smaller than the predicted part, indicating that our model has extracted all spatial/temporal correlations from the observed part for prediction.
-Table 8: Reconstruction RMSE for the observed signals in 60-minute forecast in all three datasets.
-Dataset PEMS03 METR-LA PEMS-BAY
-rec RMSE 1.44 0.49 0.39
+Dataset PEMS03 METR-LA
+$2\sigma_{\text{RMSE}}$ 0.05 0.04
+$2\sigma_{\text{MAE}}$ 0.02 0.02
+$2\sigma_{\text{MAPE(\%)}}$ 0.03 0.04

### v3

@@
 We design an iterative algorithm based on alternating direction method of multipliers (ADMM), and unroll it into a feed-forward network for data-driven parameter learning.
-We insert graph learning modules for ${\mathcal{G}}^{u}$ and ${\mathcal{G}}^{d}$ that play the role of self-attention.
-Experiments show that our unrolled networks achieve competitive traffic forecast performance as state-of-the-art prediction schemes, while reducing parameter counts drastically.
-Our code is available in https://github.com/SingularityUndefined/Unrolling-GSP-STForecast.
+We periodically insert graph learning modules for ${\mathcal{G}}^{u}$ and ${\mathcal{G}}^{d}$ that play the role of self-attention.
+Experiments show that our unrolled networks achieve competitive traffic forecast performance as state-of-the-art prediction schemes, while reducing parameter counts drastically [^1] https://github.com/SingularityUndefined/Unrolling-GSP-STForecast..
+algorithm unrolling, mixed graph, graph signal processing, lightweight model, traffic forecasting
 1 Introduction
-Transformer, based on the classical self-attention mechanism (Vaswani et al., 2017), is now a dominant deep learning (DL) architecture that has achieved state-of-the-art (SOTA) performance across multiple fields (Woźniak et al., 2020; Zamir et al., 2022).
+Transformer, based on the classical self-attention mechanism (vaswani17attention), is now a dominant deep learning (DL) architecture that has achieved state-of-the-art (SOTA) performance across multiple fields (transformer_overview_2020; zamir2022restormer).
 However, like other “off-the-shelf” DL models, transformer has a massive number of parameters, and its internal structure is not easily interpretable.
-Parameter reduction is a major industrial concern for practical computation and/or memory-constrained environments [^1] See a brief review of lightweight learning models for different applications in Appendix A.1..
-Algorithm unrolling (Monga et al., 2021) offers an alternative hybrid model-based / data-driven paradigm; by “unrolling” iterations of a model-based algorithm minimizing a well-defined optimization objective into neural layers stacked together to form a feed-forward network for data-driven parameter learning, the resulting network can be both mathematically interpretable [^2] Common in algorithm unrolling (Monga et al., 2021), “interpretability” here means that each neural layer corresponds to an iteration of an optimization algorithm minimizing a mathematically-defined objective.
+Model size reduction is a major industrial concern for practical computation and/or memory-constrained environments [^2] See a brief review of lightweight learning models for different applications in Appendix A.1..
+Algorithm unrolling (monga21) offers an alternative hybrid model-based / data-driven paradigm; by “unrolling” iterations of a model-based algorithm minimizing an optimization objective into neural layers stacked together to form a feed-forward network for data-driven parameter learning, the resulting network can be both mathematically interpretable [^3] Common in algorithm unrolling (monga21), “interpretability” here means that each neural layer corresponds to an iteration of an optimization algorithm minimizing a mathematically-defined objective.
 and competitive in performance.
-Notably, Yu et al.
-(2023) designs an algorithm minimizing a sparse rate reduction (SRR) objective that unrolls into a transformer-like neural net for image classification.
-Thuc et al.
-(2024) designs graph-based algorithms minimizing graph smoothness priors that unroll into transformer-like neural nets for image interpolation while reducing parameters.
-However, only a positive undirected graph model was used in Thuc et al.
-(2024) to capture simple pairwise correlations between neighboring pixels in a static image.
+Notably, yu23nips designs an algorithm minimizing a sparse rate reduction (SRR) objective that unrolls into a transformer-like neural net for image classification.
+Do2024 designs graph-based algorithms minimizing graph smoothness priors that unroll into transformer-like neural nets for image interpolation while reducing parameters.
+However, only a positive undirected graph model was used in Do2024 to capture simple pairwise correlations between neighboring pixels in a static image.
 Spatial-temporal data (e.g., traffic and weather) contains more complex node-to-node relationships: geographically near stations exhibit spatial correlations, while past observations influences future data, resulting in sequential relationships.
 In this paper, we study the design of transformers via graph algorithm unrolling for spatial-temporal data, using traffic forecast as a concrete example application.
-Among model-based methods, one approach to traffic forecast is to filter the signal along the spatial and temporal dimensions independently (Ramakrishna et al., 2020)e.g., employ graph spectral filters in the graph signal processing (GSP) literature (Ortega et al., 2018; Cheung et al., 2018) along the irregular spatial dimension, and traditional Fourier filters along the regular time dimension.
-Another approach is to model the spatial correlations across time as a sequence of slowly time-varying graphs, and then filter each spatial signal at time $t$ separately, where the changes in consecutive graph adjacency matrices in time are constrained by the Frobenius norm (Kalofolias et al., 2017), $\ell_{1}$-norm (Yamada et al., 2019), or low-rankness (Bagheri et al., 2024).
+We roughly categorize traffic forecast methods into model-based and DL-based methods.
+Among model-based methods, one approach to traffic forecast is to filter the signal along the spatial and temporal dimensions independently (ramakrishna2020)e.g., employ graph spectral filters in the graph signal processing (GSP) literature (ortega18ieee; cheung18) along the irregular spatial dimension, and traditional Fourier filters along the regular time dimension.
+Another approach is to model the spatial correlations across time as a sequence of slowly time-varying graphs, and then filter each spatial signal at time $t$ separately, where the changes in consecutive graph adjacency matrices in time are constrained by the Frobenius norm (kalofolias2017), $\ell_{1}$-norm (yamada2019), or low-rankness (bagheri2024).
 However, neither approach exploits spatial and temporal relationships simultaneously for optimal performance.
-Recent DL-based efforts address traffic forecast by building elaborate transformer architectures to capture spatial and temporal relations (Feng & Tassiulas, 2022; Gao et al., 2022; Jin et al., 2024) [^3] A detailed review of DL models on traffic forecast is included in Appendix A.2..
-Like transformers used in other fields (Zamir et al., 2022), these are also black boxes with large parameter counts.
-Graph attention networks (GAT) (Velickovic et al., 2018) and graph transformers (Dwivedi & Bresson, 2021) are adaptations of the self-attention mechanism in transformers for graph-structured data (e.g., computing output embeddings using only input embeddings in local graph neighborhoods).
+Recent DL-based efforts address traffic forecast by building elaborate transformer architectures to capture spatial and temporal relations (feng2022; gao2022; jin2024) [^4] A detailed review of DL models on traffic forecast is included in Appendix A.2..
+Like transformers used in other fields (zamir2022restormer), these are also black boxes with large parameter counts.
+Graph attention networks (GAT) (velickovic2018) and graph transformers (dwivedi2021) are adaptations of the self-attention mechanism in transformers for graph-structured data (e.g., computing output embeddings using only input embeddings in local graph neighborhoods).
 However, they are still uninterpretable and maintain large parameter sizes.
-In contrast, we extend Thuc et al.
-(2024) to build lightweight “white-box” transformers that are fully interpretable via mixed-graph algorithm unrolling for traffic forecast.
+In contrast, we extend Do2024 to build lightweight “white-box” transformers that are fully interpretable via mixed graph algorithm unrolling for traffic forecast.
 Specifically, we learn two graphs from data: i) undirected graph ${\mathcal{G}}^{u}$ to capture spatial correlations across geography, and ii) a directed graph ${\mathcal{G}}^{d}$ to capture sequential relationships over time.
-We show that our graph learning modules play the role of self-attention (Bahdanau et al., 2014), and thus our unrolled graph-based neural nets are transformers.
-Given the two learned graphs, we define a prediction objective for future samples in signal $\mathbf{x}$}, assuming $\mathbf{x}$} is “smooth” with respect to (w.r.t.) both ${\mathcal{G}}^{u}$ and ${\mathcal{G}}^{d}$ in variational terms: graph Laplacian regularizer (GLR) (Pang & Cheung, 2017) for undirected ${\mathcal{G}}^{u}$, and newly designed directed graph Laplacian regularizer (DGLR) and directed graph total variation (DGTV) for directed ${\mathcal{G}}^{d}$.
-We design a corresponding linear-time optimization algorithm based on alternating direction method of multipliers (ADMM) (Boyd et al., 2011) that unrolls into neural layers for parameter learning.
-For the first time, our DGLR and DGTV variational terms quantify and promote signal smoothness on a directed graph, with intuitive low-pass filter interpretations.
-Experiments show that our networks achieve competitive traffic forecast performance as SOTA prediction schemes, while employing drastically fewer parameters (6.4% of transformer-based PDFormer (Jiang et al., 2023)).
+We show that our graph learning modules play the role of self-attention (bahdanau14), and thus our unrolled graph-based neural nets are transformers.
+Given the two learned graphs, we define a prediction objective for future samples in signal $\mathbf{x}$}, assuming $\mathbf{x}$} is “smooth” with respect to (w.r.t.) both ${\mathcal{G}}^{u}$ and ${\mathcal{G}}^{d}$ in variational terms: graph Laplacian regularizer (GLR) (pang17) for undirected ${\mathcal{G}}^{u}$, and newly designed directed graph Laplacian regularizer (DGLR) and directed graph total variation (DGTV) for directed ${\mathcal{G}}^{d}$.
+We devise a corresponding linear-time optimization algorithm based on alternating direction method of multipliers (ADMM) (boyd11) that unrolls into neural layers of a feed-forward network for parameter learning.
+For the first time, our proposed DGLR and DGTV variational terms quantify and promote signal smoothness on a directed graph, with intuitive low-pass filter interpretations [^5] See a detailed review of related works on directed graph Laplacians and discussion of our novelty in directed graph frequency analysis in Appendix A.3..
+Extensive experiments show that our networks achieve competitive traffic forecast performance as SOTA prediction schemes, while employing drastically fewer parameters (7.2% of transformer-based PDFormer (jiang2023pdformer)).
 2 Preliminaries
@@
 Denote by ${\mathbf{D}}^{u}\in\mathbb{R}^{N\times N}$ a diagonal degree matrix, where $D_{i,i}^{u}=\sum_{j}W_{j,i}^{u}$.
-The symmetric graph Laplacian matrix is ${\mathbf{L}}^{u}\triangleq{\mathbf{D}}^{u}-{\mathbf{W}}^{u}$ (Ortega et al., 2018).
-${\mathbf{L}}^{u}$ is positive semi-definite (PSD) if all edge weights are non-negative $w^{u}_{i,j}\geq 0,\forall i,j$ (Cheung et al., 2018).
+The symmetric graph Laplacian matrix is ${\mathbf{L}}^{u}\triangleq{\mathbf{D}}^{u}-{\mathbf{W}}^{u}$ (ortega18ieee).
+${\mathbf{L}}^{u}$ is positive semi-definite (PSD) if all edge weights are non-negative $w^{u}_{i,j}\geq 0,\forall i,j$ (cheung18).
 The symmetric normalized graph Laplacian matrix is ${\mathbf{L}}^{u}_{n}\triangleq{\mathbf{I}}-({\mathbf{D}}^{u})^{-1/2}{\mathbf{W}}^{u}({\mathbf{D}}^{u})^{-1/2}$, while the asymmetric random-walk graph Laplacian matrix is ${\mathbf{L}}^{u}_{r}\triangleq{\mathbf{I}}-({\mathbf{D}}^{u})^{-1}{\mathbf{W}}^{u}$.
-One can define a graph spectrum by eigen-decomposing the real and symmetric graph Laplacian ${\mathbf{L}}^{u}$ (or normalized graph Laplacian ${\mathbf{L}}_{n}^{u}$), where the $k$-th eigen-pair $(\lambda_{k},{\mathbf{v}}_{k})$ is interpreted as the $k$-th graph frequency and Fourier mode, respectively (Ortega et al., 2018).
-Given $\{{\mathbf{v}}_{k}\}$ are orthonormal vectors, one can decompose a signal $\mathbf{x}$} into its graph frequency components as ${\bm{\alpha}}={\mathbf{V}}^{\top}{\mathbf{x}}$, where ${\mathbf{L}}^{u}={\mathbf{V}}\mathrm{diag}(\{\lambda_{k}\}){\mathbf{V}}^{\top}$, and $\mathbf{V}$$^{\top}$ is the graph Fourier transform (GFT).
-The $\ell_{2}$-norm graph Laplacian regularizer (GLR) (Pang & Cheung, 2017), ${\mathbf{x}}^{\top}{\mathbf{L}}^{u}{\mathbf{x}}=\sum_{(i,j)\in{\mathcal{E}}^{u}}w^{u}_{i,j}(x_{i}-x_{j})^{2}$, is used to regularize an ill-posed signal restoration problem, like denoising, to bias low-frequency signal reconstruction (i.e., signals consistent with similarity graph ${\mathcal{G}}^{u}$) given observation $\mathbf{y}$}:
-${\mathbf{x}}^{*}$ $=\arg\min_{\mathbf{x}}\|{\mathbf{y}}-{\mathbf{x}}\|^{2}_{2}+\mu\,{\mathbf{x}}^{\top}{\mathbf{L}}^{u}{\mathbf{x}}$
-$=({\mathbf{I}}+\mu{\mathbf{L}}^{u})^{-1}{\mathbf{y}}={\mathbf{V}}\mathrm{diag}\left(\frac{1}{1+\mu\lambda_{1}},\ldots,\frac{1}{1+\mu\lambda_{N}}\right){\mathbf{V}}^{\top}{\mathbf{y}}.$ (1)
+One can define a graph spectrum by eigen-decomposing the real and symmetric graph Laplacian ${\mathbf{L}}^{u}$ (or normalized graph Laplacian ${\mathbf{L}}_{n}^{u}$), where the $k$-th eigen-pair $(\lambda_{k},{\mathbf{v}}_{k})$ is interpreted as the $k$-th graph frequency and Fourier mode, respectively (ortega18ieee).
+Given $\{{\mathbf{v}}_{k}\}$ are orthonormal vectors, one can decompose a signal $\mathbf{x}$} into its graph frequency components as ${\boldsymbol{\alpha}}={\mathbf{V}}^{\top}{\mathbf{x}}$, where ${\mathbf{L}}^{u}={\mathbf{V}}\mathrm{diag}(\{\lambda_{k}\}){\mathbf{V}}^{\top}$, and $\mathbf{V}$$^{\top}$ is the graph Fourier transform (GFT).
+The $\ell_{2}$-norm graph Laplacian regularizer (GLR) (pang17), ${\mathbf{x}}^{\top}{\mathbf{L}}^{u}{\mathbf{x}}=\sum_{(i,j)\in{\mathcal{E}}^{u}}w^{u}_{i,j}(x_{i}-x_{j})^{2}$, is used to regularize an ill-posed signal restoration problem, like denoising, to bias low-frequency signal reconstruction (i.e., signals consistent with similarity graph ${\mathcal{G}}^{u}$) given observation $\mathbf{y}$}:
+${\mathbf{x}}^{*}$ $=\arg\min_{\mathbf{x}}\|{\mathbf{y}}-{\mathbf{x}}\|^{2}_{2}+\mu\,{\mathbf{x}}^{\top}{\mathbf{L}}^{u}{\mathbf{x}}=({\mathbf{I}}+\mu{\mathbf{L}}^{u})^{-1}{\mathbf{y}}$
+$={\mathbf{V}}\mathrm{diag}\left(\frac{1}{1+\mu\lambda_{1}},\ldots,\frac{1}{1+\mu\lambda_{N}}\right){\mathbf{V}}^{\top}{\mathbf{y}}.$ (1)
 $\mu$ is a bias-variance tradeoff parameter, and the low-pass filter response is $f(\lambda)=(1+\mu\lambda)^{-1}$.
@@
 The directed random-walk graph Laplacian matrix is ${\mathbf{L}}^{d}_{r}\triangleq{\mathbf{I}}-({\mathbf{D}}^{d})^{-1}{\mathbf{W}}^{d}$.
-Similar frequency notion cannot be simply defined for directed graphs via eigen-decomposition of directed graph Laplacian ${\mathbf{L}}^{d}$, because ${\mathbf{L}}^{d}$ is asymmetric, and thus the Spectral Theorem (Hawkins, 1975) does not apply.
-We circumvent this problem by designing new variational terms in Section 3.2.
+Similar frequency notion cannot be simply defined for directed graphs via eigen-decomposition of directed graph Laplacian ${\mathbf{L}}^{d}$, because ${\mathbf{L}}^{d}$ is asymmetric, and thus the Spectral Theorem (hawkins1975) does not apply.
+We circumvent this problem via new variational terms in Section 3.2.
 3 Optimization Formulation & Algorithm
@@
 3.2.1 $\ell_{2}$-norm Variational Term
-Interpreting ${\mathbf{W}}_{r}^{d}$ as a graph shift operator (GSO) (Chen et al., 2015) on ${\mathcal{G}}^{d}$, we first define an $\ell_{2}$-norm variational term [^4] A similar directed graph variational term was defined in Li et al.
-(2023) towards an objective for directed graph sampling.
+Interpreting ${\mathbf{W}}_{r}^{d}$ as a graph shift operator (GSO) (chen15) on ${\mathcal{G}}^{d}$, we first define an $\ell_{2}$-norm variational term [^6] A similar directed graph variational term was defined in li2024 towards an objective for directed graph sampling.
 We focus instead on directed graph signal restoration.
@@
 Importantly, we prove that our frequency definition using DGLR defaults to pure sinusoids as frequencies in the unweighted directed line graph case.
-Thus, our algorithm minimizing DGLR includes GSP schemes like Ramakrishna et al.
-(2020) that analyze signals along the time dimension using classical Fourier filters as special cases.
+Thus, our algorithm minimizing DGLR includes GSP schemes like ramakrishna2020 that analyze signals along the time dimension using classical Fourier filters as special cases.
 Theorem 3.1.
@@
 Example of a 3-node DAG (c), and a 3-node DAG with opposite directional edges (d).
-Remark: Using our symmetrized ${\mathcal{L}}^{d}_{r}$ does not mean directionality has been lost.
-For the 3-node DAG in Fig.
+Remark: Theorem 3.1 only gives a special case when the DGLR of a graph is equal to the GLR of its undirected base graph.
+In fact, using our symmetrized ${\mathcal{L}}^{d}_{r}$ does not mean directionality has been lost.
+As an example, for the 3-node DAG in Fig.
 2(c), the variational term is $\|{\mathbf{L}}^{d}_{r}{\mathbf{x}}\|^{2}_{2}=(x_{3}-\frac{1}{2}(x_{1}+x_{2}))^{2}$ and computes to 0 for ${\mathbf{x}}=[2~~0~~1]^{\top}$.
@@
 2(d) with opposite directions, the variational term is $\|{\mathbf{L}}^{d}_{r}{\mathbf{x}}\|^{2}_{2}=(x_{1}-x_{3})^{2}+(x_{2}-x_{3})^{2}$ and computes to 0 only if $\mathbf{x}$} is a constant vector.
+With our directed graph contruction with time windows as described in Section 3.1, since all nodes after the second time step has more than one predecessors, the DGLR of the undirected graph does not equal the GLR of its base graph.
 3.2.2 $\ell_{1}$-Norm Variational Term
@@
 $\|{\mathbf{x}}-{\mathbf{W}}^{d}_{r}{\mathbf{x}}\|_{1}=\|{\mathbf{L}}^{d}_{r}{\mathbf{x}}\|_{1}=\sum_{j\in\bar{{\mathcal{S}}}}|x_{j}-\sum_{i}w_{j,i}x_{i}|.$ (3)
-Unlike DGLR in eq. 2, DGTV is not symmetric and has no obvious frequency interpretation. Nonetheless, we provide a two-channel filterbank interpretation in the sequel.
+Unlike DGLR in Equation 2, DGTV is not symmetric and has no obvious frequency interpretation. Nonetheless, we provide a two-channel filterbank interpretation in the sequel.
 3.3 Optimization Formulation
@@
 Given defined undirected and directed edges, we write an objective for $\mathbf{x}$} containing a squared-error fidelity term and graph smoothness terms for the two graphs ${\mathcal{G}}^{u}$ and ${\mathcal{G}}^{d}$—GLR, DGLR, and DGTV:
-$\min_{\mathbf{x}}\|{\mathbf{y}}-{\mathbf{H}}{\mathbf{x}}\|^{2}_{2}+\mu_{u}{\mathbf{x}}^{\top}{\mathbf{L}}^{u}{\mathbf{x}}+\mu_{d,2}{\mathbf{x}}^{\top}{\mathcal{L}}_{r}^{d}{\mathbf{x}}+\mu_{d,1}\|{\mathbf{L}}^{d}_{r}{\mathbf{x}}\|_{1}$ (4)
+$\hskip-4.25006pt\min_{\mathbf{x}}\|{\mathbf{y}}\!-\!{\mathbf{H}}{\mathbf{x}}\|^{2}_{2}+\mu_{u}{\mathbf{x}}\!^{\top}{\mathbf{L}}\!^{u}{\mathbf{x}}+\mu_{d,2}{\mathbf{x}}\!^{\top}\!{\mathcal{L}}_{r}^{d}{\mathbf{x}}+\mu_{d,1}\!\|{\mathbf{L}}^{d}_{r}{\mathbf{x}}\|_{1},$ (4)
 where $\mu_{u},\mu_{d,2},\mu_{d,1}\in\mathbb{R}_{+}$ are weight parameters for the three regularization terms.
 Combination of $\ell_{2}$- and $\ell_{1}$-norm penalties—DGLR and DGTV for directed graph ${\mathcal{G}}^{d}$ in our case [^7] We can employ both $\ell_{2}$- and $\ell_{1}$-norm regularization terms for the undirected graph ${\mathcal{G}}^{u}$ as well.
-For simplicity, we focus on designing new regularization terms for directed graphs in this paper.
-—is called elastic net regularization in statistics (Zou & Hastie, 2005) with demonstrable improved robustness.
+For simplicity, we focus on designing new regularization terms for directed graphs here.
+—is called elastic net regularization in statistics (zou2005) with demonstrable improved robustness.
 Further, employing both regularization terms means that, after algorithm unrolling, we can adapt an appropriate mixture by learning weights $\mu_{d,2}$ and $\mu_{d,1}$ per neural layer.
-Convex minimization in eq.
-4 is composed of smooth $\ell_{2}$-norm terms and one non-smooth $\ell_{1}$-norm term.
-We pursue a divide-and-conquer approach and solve eq.
-4 via an ADMM framework (Boyd et al., 2011) by minimizing $\ell_{2}$- and $\ell_{1}$-norm terms in eq.
-4 alternately.
+Convex minimization in Equation 4 is composed of smooth $\ell_{2}$-norm terms and one non-smooth $\ell_{1}$-norm term.
+We pursue a divide-and-conquer approach and solve Equation 4 via an ADMM framework (boyd11) by minimizing $\ell_{2}$- and $\ell_{1}$-norm terms in Equation 4 alternately.
 3.4 ADMM Optimization Algorithm
-We first introduce auxiliary variable $\boldsymbol{\phi}$} and rewrite the optimization in eq. 4 as
-$\min_{{\mathbf{x}},{\boldsymbol{\phi}}}$ $\|{\mathbf{y}}-{\mathbf{H}}{\mathbf{x}}\|^{2}_{2}+\mu_{u}{\mathbf{x}}^{\top}{\mathbf{L}}^{u}{\mathbf{x}}+\mu_{d,2}{\mathbf{x}}^{\top}{\mathcal{L}}_{r}^{d}{\mathbf{x}}+\mu_{d,1}\|{\boldsymbol{\phi}}\|_{1},\qquad\mbox{s.t.}~~~{\boldsymbol{\phi}}={\mathbf{L}}^{d}_{r}{\mathbf{x}}.$ (5)
-Using the augmented Lagrangian method (Boyd & Vandenberghe, 2004), we rewrite eq. 5 in an unconstrained form:
-$\min_{{\mathbf{x}},{\boldsymbol{\phi}}}$ $\|{\mathbf{y}}-{\mathbf{H}}{\mathbf{x}}\|^{2}_{2}+\mu_{u}{\mathbf{x}}^{\top}{\mathbf{L}}^{u}{\mathbf{x}}+\mu_{d,2}{\mathbf{x}}^{\top}{\mathcal{L}}_{r}^{d}{\mathbf{x}}+\mu_{d,1}\|{\boldsymbol{\phi}}\|_{1}+{\bm{\gamma}}^{\top}({\boldsymbol{\phi}}-{\mathbf{L}}^{d}_{r}{\mathbf{x}})+\frac{\rho}{2}\|{\boldsymbol{\phi}}-{\mathbf{L}}^{d}_{r}{\mathbf{x}}\|^{2}_{2}$ (6)
-where ${\bm{\gamma}}\in\mathbb{R}^{N(T+S+1)}$ is a Lagrange multiplier vector, and $\rho\in\mathbb{R}_{+}$ is a non-negative ADMM parameter.
-We solve eq.
-6 iteratively by minimizing $\mathbf{x}$} and $\boldsymbol{\phi}$} in alternating steps till convergence.
+We first introduce auxiliary variable $\boldsymbol{\phi}$} and rewrite the optimization in Equation 4 as
+$\min_{{\mathbf{x}},{\boldsymbol{\phi}}}$ $~~\|{\mathbf{y}}-{\mathbf{H}}{\mathbf{x}}\|^{2}_{2}+\mu_{u}{\mathbf{x}}^{\top}{\mathbf{L}}^{u}{\mathbf{x}}+\mu_{d,2}{\mathbf{x}}^{\top}{\mathcal{L}}_{r}^{d}{\mathbf{x}}+\mu_{d,1}\|{\boldsymbol{\phi}}\|_{1},$
+s.t. $~~{\boldsymbol{\phi}}={\mathbf{L}}^{d}_{r}{\mathbf{x}}.$ (5)
+Using the augmented Lagrangian method (boyd04), we rewrite Equation 5 in an unconstrained form:
+$\min_{{\mathbf{x}},{\boldsymbol{\phi}}}$ $~\textstyle{\|{\mathbf{y}}-{\mathbf{H}}{\mathbf{x}}\|^{2}_{2}+\mu_{u}{\mathbf{x}}^{\top}{\mathbf{L}}^{u}{\mathbf{x}}+\mu_{d,2}{\mathbf{x}}^{\top}{\mathcal{L}}_{r}^{d}{\mathbf{x}}+\mu_{d,1}\|{\boldsymbol{\phi}}\|_{1}}$
+$~\textstyle{+{\boldsymbol{\gamma}}^{\top}({\boldsymbol{\phi}}-{\mathbf{L}}^{d}_{r}{\mathbf{x}})+\frac{\rho}{2}\|{\boldsymbol{\phi}}-{\mathbf{L}}^{d}_{r}{\mathbf{x}}\|^{2}_{2}}$ (6)
+where ${\boldsymbol{\gamma}}\in\mathbb{R}^{N(T+S+1)}$ is a Lagrange multiplier vector, and $\rho\in\mathbb{R}_{+}$ is a non-negative ADMM parameter.
+We solve Equation 6 iteratively by minimizing $\mathbf{x}$} and $\boldsymbol{\phi}$} in alternating steps till convergence.
 3.4.1 Minimizing ${\mathbf{x}}^{\tau+1}$
-Fixing ${\boldsymbol{\phi}}^{\tau}$ and ${\bm{\gamma}}^{\tau}$ at iteration $\tau$, optimization in eq.
-6 for ${\mathbf{x}}^{\tau+1}$ becomes an unconstrained convex quadratic objective.
+Fixing ${\boldsymbol{\phi}}^{\tau}$ and ${\boldsymbol{\gamma}}^{\tau}$ at iteration $\tau$, optimization in Equation 6 for ${\mathbf{x}}^{\tau+1}$ becomes an unconstrained convex quadratic objective.
 The solution is a system of linear equations:
-$\hskip-10.00002pt\left({\mathbf{H}}^{\top}{\mathbf{H}}+\mu_{u}{\mathbf{L}}^{u}+(\mu_{d,2}+\frac{\rho}{2}){\mathcal{L}}^{d}_{r}\right){\mathbf{x}}^{\tau+1}$ $=({\mathbf{L}}_{r}^{d})^{\top}(\frac{\rho}{2}\,{\boldsymbol{\phi}}^{\tau}+\frac{{\bm{\gamma}}^{\tau}}{2})+{\mathbf{H}}^{\top}{\mathbf{y}}.$ (7)
-Because coefficient matrix ${\mathbf{H}}^{\top}{\mathbf{H}}+\mu_{u}{\mathbf{L}}^{u}+(\mu_{d,2}+\frac{\rho}{2}){\mathcal{L}}^{d}_{r}$ is sparse, symmetric and PD, eq.
-7 can be solved in linear time via conjugate gradient (CG) (Shewchuk, 1994) without matrix inversion.
-Splitting $\ell_{2}$-norm Terms: Instead of solving eq.
-7 directly for ${\mathbf{x}}^{\tau+1}$ at iteration $\tau$, we introduce again auxiliary variables ${\mathbf{z}}_{u}$ and ${\mathbf{z}}_{d}$ in eq.
-5 for the $\ell_{2}$-norm GLR and DGLR terms respectively, and rewrite the optimization as
-$\min_{{\mathbf{x}},{\mathbf{z}}_{u},{\mathbf{z}}_{d}}$ $\|{\mathbf{y}}-{\mathbf{H}}{\mathbf{x}}\|^{2}_{2}+\mu_{u}{\mathbf{z}}_{u}^{\top}{\mathbf{L}}^{u}{\mathbf{z}}_{u}+\mu_{d,2}{\mathbf{z}}_{d}^{\top}{\mathcal{L}}_{r}^{d}{\mathbf{z}}_{d}+({\bm{\gamma}}^{\tau})^{\top}({\boldsymbol{\phi}}^{\tau}-{\mathbf{L}}^{d}_{r}{\mathbf{x}})+\frac{\rho}{2}\|{\boldsymbol{\phi}}^{\tau}-{\mathbf{L}}^{d}_{r}{\mathbf{x}}\|^{2}_{2}$ (8)
-s.t. $~{\mathbf{x}}={\mathbf{z}}_{u}={\mathbf{z}}_{d}.$
+$\hskip-10.00002pt\textstyle{\left({\mathbf{H}}^{\top}{\mathbf{H}}+\mu_{u}{\mathbf{L}}^{u}+(\mu_{d,2}+\frac{\rho}{2}){\mathcal{L}}^{d}_{r}\right){\mathbf{x}}^{\tau+1}}$ $=$
+$\textstyle{({\mathbf{L}}_{r}^{d})^{\top}(\frac{\rho}{2}\,{\boldsymbol{\phi}}^{\tau}+\frac{{\boldsymbol{\gamma}}^{\tau}}{2})}$ $\textstyle{+{\mathbf{H}}^{\top}{\mathbf{y}}}.$ (7)
+Because coefficient matrix ${\mathbf{H}}^{\top}{\mathbf{H}}+\mu_{u}{\mathbf{L}}^{u}+(\mu_{d,2}+\frac{\rho}{2}){\mathcal{L}}^{d}_{r}$ is sparse, symmetric and PD, Equation 7 can be solved in linear time via conjugate gradient (CG) (shewchuk94) without matrix inversion.
+Splitting $\ell_{2}$-norm Terms: Instead of solving Equation 7 directly for ${\mathbf{x}}^{\tau+1}$ at iteration $\tau$, we introduce again auxiliary variables ${\mathbf{z}}_{u}$ and ${\mathbf{z}}_{d}$ in Equation 5 for the $\ell_{2}$-norm GLR and DGLR terms respectively, and rewrite the optimization as
+$\min_{{\mathbf{x}},{\mathbf{z}}_{u},{\mathbf{z}}_{d}}$~~ $\textstyle{\|{\mathbf{y}}-{\mathbf{H}}{\mathbf{x}}\|^{2}_{2}+\mu_{u}{\mathbf{z}}_{u}^{\top}{\mathbf{L}}^{u}{\mathbf{z}}_{u}+\mu_{d,2}{\mathbf{z}}_{d}^{\top}{\mathcal{L}}_{r}^{d}{\mathbf{z}}_{d}}$
+$\textstyle{+({\boldsymbol{\gamma}}^{\tau})^{\top}({\boldsymbol{\phi}}^{\tau}-{\mathbf{L}}^{d}_{r}{\mathbf{x}})+\frac{\rho}{2}\|{\boldsymbol{\phi}}^{\tau}-{\mathbf{L}}^{d}_{r}{\mathbf{x}}\|^{2}_{2}}$ (8)
+s.t. $\textstyle{{\mathbf{x}}={\mathbf{z}}_{u}={\mathbf{z}}_{d}.}$
 Using again the augmented Lagrangian method, we rewrite the unconstrained version as
-$\min_{{\mathbf{x}},{\mathbf{z}}_{u},{\mathbf{z}}_{d}}$ $\|{\mathbf{y}}-{\mathbf{H}}{\mathbf{x}}\|^{2}_{2}+\mu_{u}{\mathbf{z}}_{u}^{\top}{\mathbf{L}}^{u}{\mathbf{z}}_{u}+\mu_{d,2}{\mathbf{z}}_{d}^{\top}{\mathcal{L}}_{r}^{d}{\mathbf{z}}_{d}+({\bm{\gamma}}^{\tau})^{\top}({\boldsymbol{\phi}}^{\tau}-{\mathbf{L}}^{d}_{r}{\mathbf{x}})+\frac{\rho}{2}\|{\boldsymbol{\phi}}^{\tau}-{\mathbf{L}}^{d}_{r}{\mathbf{x}}\|^{2}_{2}$ (9)
-$+({\bm{\gamma}}_{u}^{\tau})^{\top}({\mathbf{x}}-{\mathbf{z}}_{u})+\frac{\rho_{u}}{2}\|{\mathbf{x}}-{\mathbf{z}}_{u}\|^{2}_{2}+({\bm{\gamma}}_{d}^{\tau})^{\top}({\mathbf{x}}-{\mathbf{z}}_{d})+\frac{\rho_{d}}{2}\|{\mathbf{x}}-{\mathbf{z}}_{d}\|^{2}_{2}$
-where ${\bm{\gamma}}_{u},{\bm{\gamma}}_{d}\in\mathbb{R}^{N(T+S+1)}$ are multipliers, and $\rho_{u},\rho_{d}\in\mathbb{R}_{+}$ are ADMM parameters.
-Optimizing $\mathbf{x}$}, ${\mathbf{z}}_{u}$ and ${\mathbf{z}}_{d}$ in eq.
-9 in turn at iteration $\tau$ lead to three linear systems:
-$\hskip-4.2679pt{\left({\mathbf{H}}^{\top}{\mathbf{H}}+\frac{\rho}{2}{\mathcal{L}}^{d}_{r}+\frac{\rho_{u}+\rho_{d}}{2}{\mathbf{I}}\right)}{\mathbf{x}}^{\tau+1}$ $={({\mathbf{L}}^{d}_{r})^{\top}\!\left(\frac{{\bm{\gamma}}^{\tau}}{2}+\frac{\rho}{2}{\boldsymbol{\phi}}^{\tau}\right)-\frac{{\bm{\gamma}}_{u}^{\tau}}{2}+\frac{\rho_{u}}{2}{\mathbf{z}}_{u}^{\tau}-\frac{{\bm{\gamma}}_{d}^{\tau}}{2}+\frac{\rho_{d}}{2}{\mathbf{z}}_{d}^{\tau}+{\mathbf{H}}^{\top}{\mathbf{y}}}$ (10)
-${\left(\mu_{u}{\mathbf{L}}^{u}+\frac{\rho_{u}}{2}{\mathbf{I}}\right)}{\mathbf{z}}_{u}^{\tau+1}$ $={\frac{{\bm{\gamma}}_{u}^{\tau}}{2}+\frac{\rho_{u}}{2}{\mathbf{x}}^{\tau+1}}$ (11)
-${\left(\mu_{d,2}{\mathcal{L}}^{d}_{r}+\frac{\rho_{d}}{2}{\mathbf{I}}\right)}{\mathbf{z}}_{d}^{\tau+1}$ $={\frac{{\bm{\gamma}}_{d}^{\tau}}{2}+\frac{\rho_{d}}{2}{\mathbf{x}}^{\tau+1}}.$ (12)
-Thus, instead of solving eq.
-7 for ${\mathbf{x}}^{\tau+1}$ directly at iteration $\tau$, ${\mathbf{x}}^{\tau+1}$, ${\mathbf{z}}_{u}^{\tau+1}$ and ${\mathbf{z}}_{d}^{\tau+1}$ are optimized in turn via eq.
-10 through eq.
-12 using CG.
+$~~~{\min_{{\mathbf{x}},{\mathbf{z}}_{u},{\mathbf{z}}_{d}}}~\textstyle{\|{\mathbf{y}}-{\mathbf{H}}{\mathbf{x}}\|^{2}_{2}+\mu_{u}{\mathbf{z}}_{u}^{\top}{\mathbf{L}}^{u}{\mathbf{z}}_{u}+\mu_{d,2}{\mathbf{z}}_{d}^{\top}{\mathcal{L}}_{r}^{d}{\mathbf{z}}_{d}}$
+$\textstyle{+({\boldsymbol{\gamma}}^{\tau})^{\top}({\boldsymbol{\phi}}^{\tau}-{\mathbf{L}}^{d}_{r}{\mathbf{x}})+\frac{\rho}{2}\|{\boldsymbol{\phi}}^{\tau}-{\mathbf{L}}^{d}_{r}{\mathbf{x}}\|^{2}_{2}+({\boldsymbol{\gamma}}_{u}^{\tau})^{\top}({\mathbf{x}}-{\mathbf{z}}_{u})}$ (9)
+$\textstyle{+\frac{\rho_{u}}{2}\|{\mathbf{x}}-{\mathbf{z}}_{u}\|^{2}_{2}+({\boldsymbol{\gamma}}_{d}^{\tau})^{\top}({\mathbf{x}}-{\mathbf{z}}_{d})+\frac{\rho_{d}}{2}\|{\mathbf{x}}-{\mathbf{z}}_{d}\|^{2}_{2}}$
+where ${\boldsymbol{\gamma}}_{u},{\boldsymbol{\gamma}}_{d}\in\mathbb{R}^{N(T+S+1)}$ are multipliers, and $\rho_{u},\rho_{d}\in\mathbb{R}_{+}$ are ADMM parameters.
+Optimizing $\mathbf{x}$}, ${\mathbf{z}}_{u}$ and ${\mathbf{z}}_{d}$ in Equation 9 in turn at iteration $\tau$ lead to three linear systems:
+${\textstyle\scriptsize\left({\mathbf{H}}^{\top}{\mathbf{H}}+\frac{\rho}{2}{\mathcal{L}}^{d}_{r}+\frac{\rho_{u}+\rho_{d}}{2}{\mathbf{I}}\right)}{\mathbf{x}}^{\tau+1}$ $=\textstyle{({\mathbf{L}}^{d}_{r})^{\top}\!\left(\frac{{\boldsymbol{\gamma}}^{\tau}}{2}+\frac{\rho}{2}{\boldsymbol{\phi}}^{\tau}\right)}$
+$\textstyle{-\frac{{\boldsymbol{\gamma}}_{u}^{\tau}}{2}+\frac{\rho_{u}}{2}{\mathbf{z}}_{u}^{\tau}-}$ $\textstyle{\frac{{\boldsymbol{\gamma}}_{d}^{\tau}}{2}+\frac{\rho_{d}}{2}{\mathbf{z}}_{d}^{\tau}+{\mathbf{H}}^{\top}{\mathbf{y}}}$ (10)
+$\textstyle{\left(\mu_{u}{\mathbf{L}}^{u}+\frac{\rho_{u}}{2}{\mathbf{I}}\right){\mathbf{z}}_{u}^{\tau+1}}$ $=\textstyle{\frac{{\boldsymbol{\gamma}}_{u}^{\tau}}{2}+\frac{\rho_{u}}{2}{\mathbf{x}}^{\tau+1}}$ (11)
+$\textstyle{\left(\mu_{d,2}{\mathcal{L}}^{d}_{r}+\frac{\rho_{d}}{2}{\mathbf{I}}\right){\mathbf{z}}_{d}^{\tau+1}}$ $=\textstyle{\frac{{\boldsymbol{\gamma}}_{d}^{\tau}}{2}+\frac{\rho_{d}}{2}{\mathbf{x}}^{\tau+1}}.$ (12)
+Thus, instead of solving Equation 7 for ${\mathbf{x}}^{\tau+1}$ directly at iteration $\tau$, ${\mathbf{x}}^{\tau+1}$, ${\mathbf{z}}_{u}^{\tau+1}$ and ${\mathbf{z}}_{d}^{\tau+1}$ are optimized in turn via Equation 10 through Equation 12 using CG.
 This splitting induces more network parameters for data-driven learning after algorithm unrolling (see Section 4.1) towards better performance, and affords us spectral filter interpretations of the derived linear systems.
-Interpretation: To interpret solution ${\mathbf{z}}_{u}^{\tau+1}$ spectrally, we rewrite eq. 11 as
-$\hskip-2.84526pt{\mathbf{z}}_{u}^{\tau+1}=\left(\frac{2\mu_{u}}{\rho_{u}}{\mathbf{L}}^{u}+{\mathbf{I}}\right)^{-1}\left(\frac{{\bm{\gamma}}_{u}}{\rho_{u}}+{\mathbf{x}}^{\tau+1}\right)={\mathbf{V}}\mathrm{diag}\left(\left\{\frac{1}{1+\frac{2\mu_{u}}{\rho_{u}}\lambda_{k}}\right\}\right){\mathbf{V}}^{\top}\left(\frac{{\bm{\gamma}}_{u}}{\rho_{u}}+{\mathbf{x}}^{\tau+1}\right)$ (13)
+Interpretation: To interpret solution ${\mathbf{z}}_{u}^{\tau+1}$ spectrally, we rewrite Equation 11 as
+$\hskip-2.84526pt\textstyle{{\mathbf{z}}_{u}^{\tau+1}}$ $=\textstyle{\left(\frac{2\mu_{u}}{\rho_{u}}{\mathbf{L}}^{u}+{\mathbf{I}}\right)^{-1}\left(\frac{{\boldsymbol{\gamma}}_{u}}{\rho_{u}}+{\mathbf{x}}^{\tau+1}\right)}$ (13)
+$={\mathbf{V}}\textstyle{\mathrm{diag}\left(\left\{\frac{1}{1+\frac{2\mu_{u}}{\rho_{u}}\lambda_{k}}\right\}\right){\mathbf{V}}^{\top}\left(\frac{{\boldsymbol{\gamma}}_{u}}{\rho_{u}}+{\mathbf{x}}^{\tau+1}\right)}$
 where ${\mathbf{L}}^{u}={\mathbf{V}}\mathrm{diag}(\{\lambda_{k}\}){\mathbf{V}}^{\top}$ is the eigen-decomposition.
-Thus, using frequencies defined by eigen-pairs $\{(\lambda_{k},{\mathbf{v}}_{k})\}$ of ${\mathbf{L}}^{u}$, ${\mathbf{z}}_{u}^{\tau+1}$ is the low-pass filter output of ${\mathbf{x}}^{\tau+1}$ offset by $\frac{{\bm{\gamma}}_{u}}{\rho_{u}}$, with frequency response $f(\lambda)=(1+\frac{2\mu_{u}}{\rho_{u}}\lambda)^{-1}$.
-As done for ${\mathbf{z}}_{u}^{\tau+1}$, we can rewrite ${\mathbf{z}}_{d}^{\tau+1}$ in eq. 12 as
-${\mathbf{z}}_{d}^{\tau+1}$ $={\mathbf{U}}\mathrm{diag}\left(\left\{\frac{1}{1+\frac{2\mu_{d,2}}{\rho_{d}}\xi_{k}}\right\}\right){\mathbf{U}}^{\top}\left(\frac{{\bm{\gamma}}_{d}}{\rho_{d}}+{\mathbf{x}}^{\tau+1}\right)$ (14)
+Thus, using frequencies defined by eigen-pairs $\{(\lambda_{k},{\mathbf{v}}_{k})\}$ of ${\mathbf{L}}^{u}$, ${\mathbf{z}}_{u}^{\tau+1}$ is the low-pass filter output of ${\mathbf{x}}^{\tau+1}$ offset by $\frac{{\boldsymbol{\gamma}}_{u}}{\rho_{u}}$, with frequency response $f(\lambda)=(1+\frac{2\mu_{u}}{\rho_{u}}\lambda)^{-1}$.
+As done for ${\mathbf{z}}_{u}^{\tau+1}$, we can rewrite ${\mathbf{z}}_{d}^{\tau+1}$ in Equation 12 as
+$\hskip-10.00002pt\textstyle{{\mathbf{z}}_{d}^{\tau+1}}=\textstyle{{\mathbf{U}}\mathrm{diag}\left(\left\{\frac{1}{1+\frac{2\mu_{d,2}}{\rho_{d}}\xi_{k}}\right\}\right){\mathbf{U}}^{\top}\left(\frac{{\boldsymbol{\gamma}}_{d}}{\rho_{d}}+{\mathbf{x}}^{\tau+1}\right),}$ (14)
 where ${\mathcal{L}}^{d}_{r}={\mathbf{U}}\mathrm{diag}(\{\xi_{k}\}){\mathbf{U}}^{\top}$ is the eigen-decomposition.
-Analogously, ${\mathbf{z}}_{d}^{\tau+1}$ is the low-pass filter output of ${\mathbf{x}}^{\tau+1}$ offset by $\frac{{\bm{\gamma}}_{d}}{\rho_{d}}$, with frequency response $f(\xi)=(1+\frac{2\mu_{d,2}}{\rho_{d}}\xi)^{-1}$.
-To interpret solution ${\mathbf{x}}^{\tau+1}$ in eq. 10 as a low-pass filter output, see Appendix C.
+Analogously, ${\mathbf{z}}_{d}^{\tau+1}$ is the low-pass filter output of ${\mathbf{x}}^{\tau+1}$ offset by $\frac{{\boldsymbol{\gamma}}_{d}}{\rho_{d}}$, with frequency response $f(\xi)=(1+\frac{2\mu_{d,2}}{\rho_{d}}\xi)^{-1}$.
+To interpret solution ${\mathbf{x}}^{\tau+1}$ in Equation 10 as a low-pass filter output, see Appendix C.
 3.4.2 Minimizing ${\boldsymbol{\phi}}^{\tau+1}$
-Fixing ${\mathbf{x}}^{\tau+1}$ at iteration $\tau$, optimizing ${\boldsymbol{\phi}}^{\tau+1}$ in eq. 6 means
-${\boldsymbol{\phi}}^{\tau+1}=$ $\arg\min_{{\boldsymbol{\phi}}}g({\boldsymbol{\phi}})=\mu_{d,1}\|{\boldsymbol{\phi}}\|_{1}+({\bm{\gamma}}^{\tau})^{\top}({\boldsymbol{\phi}}-{\mathbf{L}}^{d}_{r}{\mathbf{x}}^{\tau+1})+\frac{\rho}{2}\|{\boldsymbol{\phi}}-{\mathbf{L}}^{d}_{r}{\mathbf{x}}^{\tau+1}\|^{2}_{2}.$ (15)
-We can compute eq. 15 element-wise as follows (see Appendix D for a derivation):
-$\delta$ $=({\mathbf{L}}^{d}_{r})_{i}{\mathbf{x}}^{\tau+1}-\rho^{-1}\gamma_{i}^{\tau}$ (16)
+Fixing ${\mathbf{x}}^{\tau+1}$ at iteration $\tau$, optimizing ${\boldsymbol{\phi}}^{\tau+1}$ in Equation 6 means
+$~\textstyle{{\boldsymbol{\phi}}^{\tau+1}=\arg\min_{{\boldsymbol{\phi}}}g({\boldsymbol{\phi}})}$ (15)
+$=$ $~\textstyle{\mu_{d,1}\|{\boldsymbol{\phi}}\|_{1}+({\boldsymbol{\gamma}}^{\tau})^{\top}({\boldsymbol{\phi}}-{\mathbf{L}}^{d}_{r}{\mathbf{x}}^{\tau+1})+\frac{\rho}{2}\|{\boldsymbol{\phi}}-{\mathbf{L}}^{d}_{r}{\mathbf{x}}^{\tau+1}\|^{2}_{2}}.$
+We can compute Equation 15 element-wise as follows (see Appendix D for a derivation):
+$\delta$ $=({\mathbf{L}}^{d}_{r})_{i}{\mathbf{x}}^{\tau+1}-\rho^{-1}\gamma_{i}^{\tau},$ (16)
 $\phi_{i}^{\tau+1}$ $=\text{sign}(\delta)\cdot\max(|\delta|-\rho^{-1}\mu_{d,1},0).$
-Interpretation: Soft-thresholding in eq.
-16 to compute $\phi_{i}^{\tau+1}$ attenuates the $i$-th entry of ${\mathbf{L}}^{d}_{r}{\mathbf{x}}^{\tau+1}$.
-Given that graph Laplacians—second difference matrices on graphs—are high-pass filters (Ortega et al., 2018), we interpret ${\mathbf{L}}^{d}_{r}$ as the high-pass channel of a two-channel filterbank (Vetterli & Kovacevic, 2013), where the corresponding low-pass channel is ${\mathbf{I}}-{\mathbf{L}}^{d}_{r}={\mathbf{W}}^{d}_{r}$.
-Using eq.
-16 for processing means that the low-pass channel ${\mathbf{W}}^{d}_{r}$ is unused and thus preserved.
+Interpretation: Soft-thresholding in Equation 16 to compute $\phi_{i}^{\tau+1}$ attenuates the $i$-th entry of ${\mathbf{L}}^{d}_{r}{\mathbf{x}}^{\tau+1}$.
+Given that graph Laplacians—second difference matrices on graphs—are high-pass filters (ortega18ieee), we interpret ${\mathbf{L}}^{d}_{r}$ as the high-pass channel of a two-channel filterbank (Vetterli2013), where the corresponding low-pass channel is ${\mathbf{I}}-{\mathbf{L}}^{d}_{r}={\mathbf{W}}^{d}_{r}$.
+Using Equation 16 for processing means that the low-pass channel ${\mathbf{W}}^{d}_{r}$ is unused and thus preserved.
 For example, the constant (low-frequency) signal $\mathbf{1}$} passes through the low-pass channel undisturbed, i.e., ${\mathbf{1}}={\mathbf{W}}^{d}_{r}{\mathbf{1}}$, while it is filtered out completely by the high-pass channel, i.e., ${\mathbf{0}}={\mathbf{L}}^{d}_{r}{\mathbf{1}}$.
-Thus, attenuation of the high-pass channel means eq.
-16 is a low-pass filter.
+Thus, attenuation of the high-pass channel means Equation 16 is a low-pass filter.
 3.4.3 Updating Lagrange Multipliers
-For objective eq. 6, we follow standard ADMM procedure to update Lagrange multiplier ${\bm{\gamma}}^{\tau+1}$ after each computation of ${\mathbf{x}}^{\tau+1}$ and ${\boldsymbol{\phi}}^{\tau+1}$:
-${\bm{\gamma}}^{\tau+1}$ $={\bm{\gamma}}^{\tau}+\rho({\boldsymbol{\phi}}^{\tau+1}-{\mathbf{L}}^{d}_{r}{\mathbf{x}}^{\tau+1}).$ (17)
-Splitting the $\ell_{2}$-norm terms means that additional multipliers ${\bm{\gamma}}_{u}^{\tau+1}$ and ${\bm{\gamma}}_{d}^{\tau+1}$ in objective eq.
-9 must also be updated after each computation of ${\mathbf{x}}^{\tau+1}$, ${\mathbf{z}}_{u}^{\tau+1}$, ${\mathbf{z}}_{d}^{\tau+1}$, and ${\boldsymbol{\phi}}^{\tau+1}$:
-${\bm{\gamma}}_{u}^{\tau+1}={\bm{\gamma}}_{u}^{\tau}+\rho_{u}({\mathbf{x}}^{\tau+1}-{\mathbf{z}}_{u}^{\tau+1}),~~~~~~~~~{\bm{\gamma}}_{d}^{\tau+1}={\bm{\gamma}}_{d}^{\tau}+\rho_{d}({\mathbf{x}}^{\tau+1}-{\mathbf{z}}_{d}^{\tau+1}).$ (18)
+For objective Equation 6, we follow standard ADMM procedure to update Lagrange multiplier ${\boldsymbol{\gamma}}^{\tau+1}$ after each computation of ${\mathbf{x}}^{\tau+1}$ and ${\boldsymbol{\phi}}^{\tau+1}$:
+${\boldsymbol{\gamma}}^{\tau+1}$ $={\boldsymbol{\gamma}}^{\tau}+\rho({\boldsymbol{\phi}}^{\tau+1}-{\mathbf{L}}^{d}_{r}{\mathbf{x}}^{\tau+1}).$ (17)
+Splitting the $\ell_{2}$-norm terms means that additional multipliers ${\boldsymbol{\gamma}}_{u}^{\tau+1}$ and ${\boldsymbol{\gamma}}_{d}^{\tau+1}$ in objective Equation 9 must also be updated after each computation of ${\mathbf{x}}^{\tau+1}$, ${\mathbf{z}}_{u}^{\tau+1}$, ${\mathbf{z}}_{d}^{\tau+1}$, and ${\boldsymbol{\phi}}^{\tau+1}$:
+$\begin{aligned} {\boldsymbol{\gamma}}_{u}^{\tau+1}&={\boldsymbol{\gamma}}_{u}^{\tau}+\rho_{u}({\mathbf{x}}^{\tau+1}-{\mathbf{z}}_{u}^{\tau+1}),\\ {\boldsymbol{\gamma}}_{d}^{\tau+1}&={\boldsymbol{\gamma}}_{d}^{\tau}+\rho_{d}({\mathbf{x}}^{\tau+1}-{\mathbf{z}}_{d}^{\tau+1}).\end{aligned}.$ (18)
 The ADMM loop is repeated till ${\mathbf{x}}^{\tau+1}$ and ${\boldsymbol{\phi}}^{\tau+1}$ converge.
@@
 We unroll our iterative ADMM algorithm to an interpretable neural net.
-The crux to our network is two periodically inserted graph learning modules for undirected and directed edge weight computation that are akin to the classical self-attention mechanism in transformers (Bahdanau et al., 2014).
+The crux to our network is two periodically inserted graph learning modules for undirected and directed edge weight computation that are akin to the classical self-attention mechanism in transformers (bahdanau14).
 4.1 ADMM Algorithm Unrolling
 We implement each iteration of our algorithm in Section 3.4 as a neural layer.
-Specifically, each ADMM iteration is implemented as four sub-layers in sequence: solving linear systems eq.
-10 for ${\mathbf{x}}^{\tau+1}$, eq.
-11 for ${\mathbf{z}}_{u}^{\tau+1}$, and eq.
-12 for ${\mathbf{z}}_{d}^{\tau+1}$ via CG, and computing ${\boldsymbol{\phi}}^{\tau+1}$ via eq.
-16.
-CG is itself an iterative descent algorithm, where parameters $\alpha$ and $\beta$ corresponding to step size and momentum can be tuned end-to-end (see Appendix E for detailed implementation).
+Specifically, each ADMM iteration is implemented as four sub-layers in sequence: solving linear systems Equation 10 for ${\mathbf{x}}^{\tau+1}$, Equation 11 for ${\mathbf{z}}_{u}^{\tau+1}$, and Equation 12 for ${\mathbf{z}}_{d}^{\tau+1}$ via CG, and computing ${\boldsymbol{\phi}}^{\tau+1}$ via Equation 16.
+CG is itself an iterative descent algorithm, where parameters $\alpha$ and $\beta$ corresponding to step size and momentum can be tuned end-to-end (see Appendix E.1 for detailed implementation).
 Weight parameters $\mu_{u},\mu_{d,2},\mu_{d,1}$ for prior terms and ADMM parameters $\rho,\rho_{u},\rho_{d}$ are also tuned per layer.
-Multipliers ${\bm{\gamma}}^{\tau+1},{\bm{\gamma}}_{u}^{\tau+1},{\bm{\gamma}}_{d}^{\tau+1}$ are then updated via eq.
-17 and eq.
-18 to complete one layer.
-Parameters learned during back-propagation for each ADMM block $b$ are denoted by ${\bm{\Theta}}_{b}$.
+Multipliers ${\boldsymbol{\gamma}}^{\tau+1},{\boldsymbol{\gamma}}_{u}^{\tau+1},{\boldsymbol{\gamma}}_{d}^{\tau+1}$ are then updated via Equation 17 and Equation 18 to complete one layer.
+Parameters learned during back-propagation for each ADMM block $b$ are denoted by ${\boldsymbol{\Theta}}_{b}$.
 See Fig.
-3 for an illustration.
+3 for an illustration, and Appendix E.2 for the detailed algorithms.
 4.2 Self-Attention Operator in Transformer
-We first review the basic self-attention mechanism: a scaled dot product following by a softmax operation (Bahdanau et al., 2014).
+We first review the basic self-attention mechanism: a scaled dot product following by a softmax operation (bahdanau14).
 Denote by ${\mathbf{x}}_{i}\in\mathbb{R}^{E}$ an embedding for token $i$.
@@
 Using softmax, non-negative attention weight $a_{i,j}$ is computed from $e(i,j)$’s as
-$a_{i,j}=\frac{\exp(e(i,j))}{\sum_{l=1}^{N}\exp(e(i,l))},~~~~~~e(i,j)=({\mathbf{Q}}{\mathbf{x}}_{j})^{\top}({\mathbf{K}}{\mathbf{x}}_{i}).$ (19)
+$\textstyle{a_{i,j}=\frac{\exp(e(i,j))}{\sum_{l=1}^{N}\exp(e(i,l))},~~e(i,j)=({\mathbf{Q}}{\mathbf{x}}_{j})^{\top}({\mathbf{K}}{\mathbf{x}}_{i})}.$ (19)
 Given attention weights $a_{i,j}$, output embedding ${\mathbf{y}}_{i}$ for token $i$ is computed as
-${\mathbf{y}}_{i}=\sum_{l=1}^{N}a_{i,l}{\mathbf{x}}_{l}{\mathbf{V}}$ (20)
+$\textstyle{{\mathbf{y}}_{i}=\sum_{l=1}^{N}a_{i,l}{\mathbf{x}}_{l}{\mathbf{V}}},$ (20)
 where ${\mathbf{V}}\in\mathbb{R}^{E\times E}$ is a value matrix.
@@
 Undirected Graph Learning: We learn an undirected graph ${\mathcal{G}}^{u}$ in a module $\texttt{UGL}_{b}$ at block $b$.
-For each node $i$, we compute a low-dimensional feature vector ${\mathbf{f}}_{i}^{u}=F^{u}({\mathbf{e}}_{i})\in\mathbb{R}^{K}$, where ${\mathbf{e}}_{i}\in\mathbb{R}^{E}$ is a vector of signal values, position embeddings in time and Laplacian embeddings of the physical graph as done in Feng & Tassiulas (2022), and $F^{u}(\cdot)$ is a chosen learned nonlinear feature function $F^{u}:\mathbb{R}^{E}\mapsto\mathbb{R}^{K}$, e.g., a shallow GNN (see Appendix F.3 for details).
-As done in Thuc et al.
-(2024), the feature distance between nodes $i$ and $j$ is computed as the Mahalanobis distance $\text{d}^{u}(i,j)$:
-$\text{d}^{u}(i,j)=({\mathbf{f}}_{i}^{u}-{\mathbf{f}}_{j}^{u})^{\top}{\mathbf{M}}({\mathbf{f}}_{i}^{u}-{\mathbf{f}}_{j}^{u})$ (21)
+For each node $i$, we compute a low-dimensional feature vector ${\mathbf{f}}_{i}^{u}=F^{u}({\mathbf{e}}_{i})\in\mathbb{R}^{K}$, where ${\mathbf{e}}_{i}\in\mathbb{R}^{E}$ is a vector of signal values, position embeddings in time and Laplacian embeddings of the physical graph as done in feng2022, and $F^{u}(\cdot)$ is a chosen learned nonlinear feature function $F^{u}:\mathbb{R}^{E}\mapsto\mathbb{R}^{K}$, e.g., a shallow GNN (see Appendix F.3 for details).
+As done in Do2024, the feature distance between nodes $i$ and $j$ is computed as the Mahalanobis distance $\text{d}^{u}(i,j)$:
+$\textstyle{\text{d}^{u}(i,j)=({\mathbf{f}}_{i}^{u}-{\mathbf{f}}_{j}^{u})^{\top}{\mathbf{M}}({\mathbf{f}}_{i}^{u}-{\mathbf{f}}_{j}^{u})},$ (21)
 where ${\mathbf{M}}\succeq 0$ is a symmetric PSD metric matrix of dimension $\mathbb{R}^{K\times K}$.
-Distance in eq.
-21 is non-negative and symmetric, i.e., $\text{d}^{u}(i,j)=\text{d}^{u}(j,i)$.
+Distance in Equation 21 is non-negative and symmetric, i.e., $\text{d}^{u}(i,j)=\text{d}^{u}(j,i)$.
 Weight $w^{u}_{i,j}$ of undirected edge $(i,j)\in{\mathcal{E}}^{u}$ is computed as
-$w^{u}_{i,j}=\frac{\exp\left(-\text{d}^{u}(i,j)\right)}{\sqrt{\sum_{l\in{\mathcal{N}}_{i}}\exp(-\text{d}^{u}(i,l))}\sqrt{\sum_{k\in{\mathcal{N}}_{j}}\exp(-\text{d}^{u}(k,j))}}$ (22)
+$\textstyle{w^{u}_{i,j}=\frac{\exp\left(-\text{d}^{u}(i,j)\right)}{\sqrt{\sum_{l\in{\mathcal{N}}_{i}}\exp(-\text{d}^{u}(i,l))}\sqrt{\sum_{k\in{\mathcal{N}}_{j}}\exp(-\text{d}^{u}(k,j))}}},$ (22)
 where the denominator summing over 1-hop neighborhood ${\mathcal{N}}_{i}$ is inserted for normalization.
-Remark: Interpreting distance $\text{d}^{u}(i,j)$ as e(i,j)-e(i,j), edge weights $w^{u}_{i,j}$’s in eq.
-22 are essentially attention weights $a_{i,j}$’s in eq.
-19, and thus the undirected graph learning module constitutes a self-attention mechanism, albeit requiring fewer parameters, since parameters for function $F^{u}(\cdot)$ and metric matrix $\mathbf{M}$} can be much smaller than large and dense query and key matrices, $\mathbf{Q}$} and $\mathbf{K}$}.
-Further, no value matrix $\mathbf{V}$} is required, since output signal ${\mathbf{x}}^{\tau+1}$ is computed through a set of low-pass filters—eq.
-10, eq.
-11, eq.
-12 and eq.
-16—derived from our optimization of objective eq.
-4.
+Remark: Interpreting distance $\text{d}^{u}(i,j)$ as e(i,j)-e(i,j), edge weights $w^{u}_{i,j}$’s in Equation 22 are essentially attention weights $a_{i,j}$’s in Equation 19, and thus the undirected graph learning module constitutes a self-attention mechanism, albeit requiring fewer parameters, since parameters for function $F^{u}(\cdot)$ and metric matrix $\mathbf{M}$} can be much smaller than large and dense query and key matrices, $\mathbf{Q}$} and $\mathbf{K}$}.
+Further, no value matrix $\mathbf{V}$} is required, since output signal ${\mathbf{x}}^{\tau+1}$ is computed through a set of low-pass filters—Equation 10, Equation 11, Equation 12 and Equation 16—derived from our optimization of objective Equation 4.
 Directed Graph Learning: Similarly, we learn a directed graph ${\mathcal{G}}^{d}$ in a module $\texttt{DGL}_{b}$ at block $b$.
 For each node $i$, we compute a feature vector ${\mathbf{f}}^{d}_{i}=F^{d}({\mathbf{e}}_{i})\in\mathbb{R}^{K}$ using feature function $F^{d}(\cdot)$.
-Like eq.
-21, the Mahalanobis distance $\text{d}^{d}(i,j)$ between nodes $i$ and $j$ is computed from ${\mathbf{f}}_{i}^{d}$ and ${\mathbf{f}}_{j}^{d}$ with a learned PSD metric $\mathbf{P}$}.
-Weight $w_{j,i}^{d}$ of directed edge $i,j]\in{\mathcal{E}}^{d}$ is then computed using exponentials and normalization, as done in eq.
-22 for an undirected edge.
-Parameters learned via back-propagation by the two graph learning modules at block bb—parameters of $F^{u}(\cdot)$, $F^{d}(\cdot)$, $\mathbf{M}$} and $\mathbf{P}$}—are denoted by ${\bm{\Phi}}_{b}$.
+Like Equation 21, the Mahalanobis distance $\text{d}^{d}(i,j)$ between nodes $i$ and $j$ is computed from ${\mathbf{f}}_{i}^{d}$ and ${\mathbf{f}}_{j}^{d}$ with a learned PSD metric $\mathbf{P}$}.
+Weight $w_{j,i}^{d}$ of directed edge $i,j]\in{\mathcal{E}}^{d}$ is then computed using exponentials and normalization as
+${w_{j,i}^{d}=\frac{\exp(-\text{d}^{d}(j,i)}{\sum_{(i,k)\in{\mathcal{E}}^{d}}\exp(-\text{d}^{d}(k,i))}.}$ (23)
+Parameters learned via back-propagation by the two graph learning modules at block bb—parameters of $F^{u}(\cdot)$, $F^{d}(\cdot)$, $\mathbf{M}$} and $\mathbf{P}$}—are denoted by ${\boldsymbol{\Phi}}_{b}$.
 5 Experiments
 5.1 Experimental Setup
-We evaluate our model’s performance on commonly used traffic speed dataset METR-LA (Li et al., 2017) and traffic flow dataset PEMS03 (Guo et al., 2022), each of which contains city-size traffic data with a sampling interval of 5 minutes in a local region, together with the real connections and distances (or travel cost) between sensors.
-We reduce the size of the original dataset by sub-sampling 1/3 of all data uniformly, and split it into training, validation, and test sets by 6:2:2.
+We evaluate our model’s performance on commonly used traffic speed dataset METR-LA (li2017diffusion) and traffic flow dataset PEMS03 (9346058), each of which contains city-size traffic data with a sampling interval of 5 minutes in a local region, together with the real connections and distances (or travel cost) between sensors.
+We reduce the size of the original dataset by sub-sampling 1/3 of all data uniformly [^8] With a drastically smaller parameter count, our model focuses on the efficiency in utilizing the observations and the effectiveness with small training sets.
+Uniform subsampling reduces the size of the training set while making full use of the entire observation.
+See a detailed analysis in Appendix G.4., and split it into training, validation, and test sets by 6:2:2.
 We predict the traffic speed/flow in the following 30 / 60 /120 minutes (6 / 12 /24 steps) from the previously observed 60 minutes (12 steps) for all datasets.
@@
 5.2 Experimental Results
-Table 1: Comparison of categorized baselines and model parameters for 60-minute forecasting on the PEMS03 dataset against ours.
-Category Selected Models Params #
-Model-based VAR (Reinsel, 2003) -
-[…]
-GATs GMAN (Zheng et al., 2020), ST-Wave (Fang et al., 2023) 210K, 883K
-Transformers PDFormer (Jiang et al., 2023), STAEformer (Liu et al., 2023) 531K, 1404K
-Adaptive-graph models Graph WaveNet (GWN) (Wu et al., 2019), AGCRN (Bai et al., 2020) 277K, 749K
-Simple linear models STID (Shao et al., 2022), SimpleTM (Chen et al., 2025) 123K, 540K
-Mixed-graph unrolling Ours 34K
-  • $\dagger$
-    STGCN predicts only one time step at a time. We train the model to predict only the next step, and predict the entire sequence recursively.
 Table 1: Comparison of RMSE / MAE / MAPE(%) metrics of our lightweight transformer to baseline models on 30 / 60-minute forecasting in PeMS03 and METR-LA datasets.
 We use boldface for the smallest error, color the 2nd and 3nd small error in blue, and underline the next 2 smallest errors.
-Dataset & PEMS03 (358 nodes, 547 edges) METR-LA (207 nodes, 1315 edges)
+Dataset & PEMS03 (358 nodes, 547 edges) METR-LA (207 nodes, 1,315 edges)
 Horizon 30 minutes 60 minutes 120 minutes 30 minutes 60 minutes 120 minutes
@@
 SimpleTM 23.75 / 15.13 / 15.59 25.56 / 15.97 / 15.49 30.58 / 18.73 / 18.18 8.76 / 4.12 / 7.63 11.96 / 5.49 / 9.65 15.44 / 7.64 / 12.27
-Ours 26.10 / 17.03 / 18.85 27.67 / 17.46 / 17.72 34.40 / 21.16 / 23.47 10.35 / 4.22 / 9.70 12.34 / 5.18 / 11.80 15.19 / 7.56 / 16.12
+Ours 25.05 / 15.85 / 16.49 26.96 / 16.58 / 18.07 34.06 / 20.23 / 23.86 10.06 / 4.05 / 9.23 12.17 / 5.27 / 11.78 15.19 / 7.14 / 16.44
+Table 2: Comparison of categorized baselines and model size for 60-minute forecasting on the PEMS03 dataset against ours.
+Category Selected Models Params #
+Model-based VAR (reinsel2003elements) -
+[…]
+STSGCN (song2020spatial) 3,496K
+GATs GMAN (zheng2020gman) 210K
+ST-Wave (fang2023spatio) 883K
+Transformers PDFormer (jiang2023pdformer) 531K
+STAEformer (liu2023spatio) 1,404K
+Adaptive-graph models Graph WaveNet (10.5555/3367243.3367303) 277K
+AGCRN (10.5555/3495724.3497218) 749K
+MLP-based models STID (10.1145/3511808.3557702) 123K
+SimpleTM (chen2025simpletm) 540K
+Mixed-graph unrolling Ours 38K
 We train each run of our model on an entire NVIDIA GeForce RTX 3090.
@@
 Table 1 shows the performance of our models and baselines in traffic forecasting, and the last column of Table 2 compares the number of parameters.
-We observe that our models reach comparable performance to most baseline models, while employ drastically fewer parameters (6.4% of transformer-based PDFormer).
+We observe that our models achieve comparable performance to most baseline models, achieving top-3 performance in at least one metric for 30- and 60-minute forecasting, while employing drastically fewer parameters (7.2% of transformer-based PDFormer).
 Among all the models, ST-Wave and SimpleTM generally achieve the best performance.
-However, ST-Wave relies on an application-specific usage of wavelet transforms, and SimpleTM employs linear models with no inductive bias, which is parameter-heavy (Battaglia et al., 2018); both models require even more parameters than PDFormer.
-Our unrolled model also generally outperforms the multi-attention model GMAN.
-Though both our model and GMAN leverage local graph neighborhoods when computing attention, we dramatically reduce parameters by replacing classical self-attention mechanisms in transformers with unrolled ADMM iterations with no observable degradation in performance.
+However, ST-Wave relies on an application-specific usage of wavelet transforms, and SimpleTM employs linear models with no inductive bias, which is parameter-heavy (47094) and lacks interpretability; both models require even more parameters than PDFormer.
 The largest GCN model, STSGCN, fails to achieve a satisfying performance for PEMS03, where the real connections are highly sparse.
 STAEformer, the largest transformer, also overfits heavily in the reduced PEMS03 datasets.
-In summary, none of the DL-based baseline schemes unambiguously achieves the best performance across the two datasets, while all require significantly more parameters than our models.
-In contrast, our model captures the complex and evolving spatial-temporal relationship accurately using a mixed graph, demonstrating consistent performance for all predicting horizons; notably, the performance gap to the best model decreases as the horizon grows.
-Further experiments on data efficiency are presented in Appendix G, and a comprehensive ablation study is provided in Appendix H.
+We also observe that our unrolled model also generally outperforms the multi-attention model GMAN.
+Though both our model and GMAN leverage local graph neighborhoods when computing attention, we dramatically reduce parameters by replacing classical self-attention mechanisms in transformers with unrolled ADMM iterations with no observable degradation in performance.
+In summary, no single DL-based baseline consistently outperforms the others across both datasets; furthermore, all baselines require significantly more parameters than our model.
+In contrast, our model achieves top-3 forecasting performance in two of three prediction windows for both datasets (top-5 in the remaining prediction window), in at least one metric.
+We plot the three metrics against parameter size (with computational cost coded in color) for 60-minute forecast on the PEMS03 datasets in Figure 4.
+The conventional baseline models with larger parameter counts generally exhibit lower prediction errors, showing an overall downward trend in the plot, except the outlier STAEFormer which overfits heavily.
+In contrast, our model lies distinctly apart from the main trend in the lower-left corner in each plot, delivering the accuracy of models over $10\times$ larger while using dramatically fewer parameters.
+Figure 4: The trade-off between the number of parameters and performance for 60-minute forecast on PEMS03 dataset.
+The color of each scatter point represents the inference computational cost (in MFLOPs) in Table 7 in Appendix G.3.
+Additional experiments are presented in Appendix G.
+5.3 Ablation Studies
+For ablation studies, we discuss the effectiveness of the new directed graph priors, DGLR and DGTV, and the benefit of directed-graph modeling of temporal relations.
+First, we discuss the effectiveness of the new directed graph priors, DGLR and DGTV, as well as the benefit of directed-graph modeling of temporal relations.
+We separately test the effects of DGLR and DGTV by removing the prior and unrolling the ADMM algorithm.
+We denote these models as “w/o DGTV” and “w/o DGLR”.
+To test the benefit of the directed-graph modeling, we change the temporal graph ${\mathcal{G}}^{d}$ into a fully undirected graph ${\mathcal{G}}^{n}$ while maintaining the connections, and replace the directed graph priors with a GLR prior ${\mathbf{x}}^{\top}{\mathbf{L}}^{n}{\mathbf{x}}$ with the same coefficient $\mu_{n}=\mu_{d,2}$.
+Detailed derivations of the ADMM algorithm for all the above ablation studies are in Appendix I.
+As is summarized in Table 3, removing the DGLR or DGTV term leads to noticeable drops in performance, which underscores their importance as directed graph priors.
+Moreover, our model outperforms the fully-undirected-graph-based model, demonstrating that sequential relationships—such as time—are better modeled as directed graphs.
+Additionally, we ran ablations of the splitting-term strategy by optimizing ${\mathbf{x}}^{\tau+1}$ using only Equation 7.
+However, the unrolled model crashed within the first few training epochs due to numerical instability, indicating that splitting the $\ell_{2}$ terms can stabilize the training process, since CG in solving Equation 7 converges much slower than in solving Equation 10-Equation 12.
+See Appendix H for further ablation studies in the number of nearest neighbors $k$, time window size $W$,and number of ADMM blocks and layers.
+Table 3: Performance comparison in RMSE / MAE / MAPE(%) of the ablation studies for mixed-graph modeling and directed graph priors in 60-minute forecasting with PEMS03 and METR-LA.
+Model PEMS03 METR-LA
+Undirected graph 29.00 / 18.21 / 19.99 12.53 / 5.37 / 11.97
+w/o DGTV 27.62 / 17.54 / 19.04 12.59 / 5.39 / 12.26
+w/o DGLR 30.53 / 18.69 / 19.06 12.45 / 5.34 / 11.98
+Ours (Mixed graph) 26.96 / 16.58 / 18.07 12.17 / 5.27 / 11.78
+5.4 Interpreting the Learned Graph via Eigenvector Centrality
+We analyze the relationship between the last undirected graph learned by the model and the observed traffic flows.
+For the best trained model on a 60-minute forecast on PEMS03, we calculate the learned weighted adjacency matrices ${\mathbf{W}}^{u}$ of the undirected graphs on the first 288 time steps (24 hours) in the test set.
+We calculate the normalized positive eigenvector corresponding to the largest eigenvalue of ${\mathbf{W}}^{u}$, known as the Perron vector, which is strictly positive by the Perron-Frobenius Theorem (johnson12).
+The Perron vector ${\mathbf{v}}\in\mathbb{R}^{N}$ reflects the eigenvector centrality (bonacich87) of the individual nodes: the “importance” of individual nodes, where a node is deemed important if it is connected to other important nodes, i.e., $\lambda v_{i}=\sum_{j}W^{u}_{i,j}v_{j}$ for $v_{j}>0$.
+(a) Neighborhood of nodes #27 and #350.
+(b) Neighborhood of node #171.
+(c) Neighborhood of node #328.
+Figure 5: The traffic patterns of the nodes in the neighborhood of the selected important nodes by Perron vectors. The traffic flows of selected nodes are bolded.
+Figure 6: The evolution of eigenvector centrality in the learned undirected graph in the last block and traffic flow in the first 24 hours in the test split of the PEMS03 dataset.
+Figure 6 shows the evolution of the top-10 Perron vector entries and the corresponding traffic flow.
+First, we observe that evident disturbances in the Perron vector entries only exist when there is a rapid lift in the traffic flows from near zero to hundreds, practically corresponding to the beginning of a new morning.
+Furthermore, the eigenvector centrality returns to the initial state once the traffic flow becomes stable, and there are no significant changes in the ranking of the eigenvector centrality, indicating that the node set with the largest Perron vector entries is generally consistent throughout the observation, which marks the critical congestion hubs.
+To further examine the role of the discovered important nodes in short-range interactions, we then select the four most important nodes #328, #27, #350 and #171 according to the elements in the Perron vectors, and plot the traffic flow of each node and the physical neighbors in selected windows of 30 and 60 steps in the daytime in Figure 5, where the selected important nodes are plotted with bolder lines (note that the neighborhood of nodes #27 and #350 are exactly the same set).
+In the red box in Figure 5(a), the spike in node #27 results in the negative spike in the neighboring nodes #274, #24 and #351.
+Further, the combined effect of this spike and the bump in node #350 results in the overshoot after the downward spike in node #274.
+The signal trends in the red box in Figure 5(b) clearly reflect the time delay of the traffic flow in the neighboring nodes except the “quiet sequences” to the selected node #171.
+Both figures 5(a) and 5(b) demonstrate clear spatiotemporal dependency patterns captured by the learned undirected graph, where high-centrality nodes play a dominant role in propagating local disturbances and temporal delays to their physical neighbors.
+Additionally, as shown in Figure 5(c), for the traffic patterns in the neighborhood of node #328, which has the largest Perron vector entry, two of its three neighboring nodes exhibit the 5th (#171) and 7th (#167) largest eigenvector centralities.
+This observation is consistent with the rationale that the most important node tends to be connected to other highly important nodes.
+Consequently, the traffic flow trends in the neighborhood of node #350 tend to be less affected by the central node, as its neighbors exhibit relatively high eigenvector centralities themselves.
+In conclusion, our model successfully discovers physically interpretable graph structures from data, and the learned node importance is highly consistent with actual traffic dynamics.
+Practically, our model can identify the key intersections that critically influence traffic dynamics and further guide traffic engineering in alleviating future traffic congestion.
 6 Conclusion
@@
 We will study extensions to signed distances and more complex graph modeling as future work.
+Impact Statement
+This paper presents work whose goal is to advance the field of Machine Learning. There are many potential societal consequences of our work, none which we feel must be specifically highlighted here.
 Appendix A Related Works
@@
 Given the importance of parameter reduction for increasingly large deep learning models, there exist various studies towards this common goal across application domains.
-For large language models (LLMs), smaller transformer models have been achieved via low-rank assumption (Hu et al., 2022) or parameter quantization (Frantar et al., 2023).
-For speech enhancement, Oostermeijer et al.
-(2021) proposes local causal self-attention, while Zhao & Madhu (2025) proposes parameter sharing across spectral and time dimensions.
-For low-light image enhancement, Brateanu et al.
-(2025) processes the Y-channel via convolution, pooling, and multi-head self-attention, separately from the U- and V-channels.
-For hyperspectral image classification, Sun et al.
-(2024) proposes a novel multiscale 3D-2D mixed CNN design.
-For cervical cancer detection, Pacal (2024) substitutes larger MBConv blocks in the MaxViT architecture with smaller ConvNeXtv2 blocks, and MLP blocks with GRN-based MLPs.
-Similar architecture miniaturization techniques have been proposed for edge devices (Mishra & Gupta, 2024) and resource-constrained environments (Liu et al., 2024).
+For large language models (LLMs), smaller transformer models have been achieved via low-rank assumption (Hu2022LoRA) or parameter quantization (Frantar2023GPTQ).
+For speech enhancement, oostermeijer2021lightweight proposes local causal self-attention, while zhao2025eusipco proposes parameter sharing across spectral and time dimensions.
+For low-light image enhancement, Brateanu2025 processes the Y-channel via convolution, pooling, and multi-head self-attention, separately from the U- and V-channels.
+For hyperspectral image classification, Sun2024 proposes a novel multiscale 3D-2D mixed CNN design.
+For cervical cancer detection, Pacal2024 substitutes larger MBConv blocks in the MaxViT architecture with smaller ConvNeXtv2 blocks, and MLP blocks with GRN-based MLPs.
+Similar architecture miniaturization techniques have been proposed for edge devices (Mishra2024) and resource-constrained environments (Liu2024).
 For traffic prediction, there are practical use cases where parameter reduction is important.
 For example, drivers in a city predict local traffic conditions (in an area-specific neighborhood graph) on their memory-constrained mobile devices every five minutes to compute optimal routes to their destinations, based on updated congestion information shared in their local neighborhoods, instead of sending individual requests to the cloud and overwhelming the central server.
-Thuc et al.
-(2024) achieves a lightweight model for image interpolation in a more principled manner: first design a linear-time optimization algorithm minimizing a chosen undirected graph smoothness prior (such as graph Laplacian regularizer (GLR) (Pang & Cheung, 2017) or graph total variation (GTV) (Bai et al., 2019)), then unroll its iterations into neural layers to compose a “white-box” transformer-like neural net for data-driven parameter learning.
+Do2024 achieves a lightweight model for image interpolation in a more principled manner: first design a linear-time optimization algorithm minimizing a chosen undirected graph smoothness prior (such as graph Laplacian regularizer (GLR) (pang17) or graph total variation (GTV) (bai19)), then unroll its iterations into neural layers to compose a “white-box” transformer-like neural net for data-driven parameter learning.
 This algorithm unrolling approach results in a lightweight and mathematically interpretable network.
-Our current work differs from Thuc et al.
-(2024) in that we model the more complex node-to-node relationships in spatial-temporal data using a mixed graph: i) an undirected graph ${\mathcal{G}}^{u}$ to capture spatial correlations among geographically near stations, and ii) a directed graph ${\mathcal{G}}^{d}$ to capture sequential relationships along the temporal dimension.
+Our current work differs from Do2024 in that we model the more complex node-to-node relationships in spatial-temporal data using a mixed graph: i) an undirected graph ${\mathcal{G}}^{u}$ to capture spatial correlations among geographically near stations, and ii) a directed graph ${\mathcal{G}}^{d}$ to capture sequential relationships along the temporal dimension.
 In particular, the notion of “frequencies” for directed graphs is not well understood in the graph signal processing (GSP) community.
-We are the first to define this notion and promote low-pass signal reconstruction via our designed $\ell_{2}$-norm directed graph Laplacian regularizer (DGLR) variational term in eq.
-2 and $\ell_{1}$-norm directed graph total variation (DGTV) variational term in eq.
-3, both derived from the graph shift operator (GSO) (Chen et al., 2015).
+We are the first to define this notion and promote low-pass signal reconstruction via our designed $\ell_{2}$-norm directed graph Laplacian regularizer (DGLR) variational term in Equation 2 and $\ell_{1}$-norm directed graph total variation (DGTV) variational term in Equation 3, both derived from the graph shift operator (GSO) (chen15).
 To the best of our knowledge, we are the first define and promote smooth signals on a directed graph in a data-driven transformer setting.
@@
 Traffic forecasting problems contain spatial and temporal correlations between stations and time steps; thus, Graph Neural Network (GNN) is naturally employed in early research.
-Early research, such as STGCN (Yu et al., 2018), STSGCN (Song et al., 2020) and GGRNN (Ruiz et al., 2020), employs Graph Convolutional Networks (GCNs) to capture spatial relations and uses 1D convolutions and RNNs to process temporal relationships.
-ASTGNN (Ta et al., 2022), Graph WaveNet (Wu et al., 2019) and AGCRN (Bai et al., 2019) explored adaptive graph settings to capture potential non-local signal relations over the graph, trading in interpretability for improved performance.
-Graph Attention Networks (GATs) learn graph weights from embedded signals to enable flexible and data-targeted graph learning, typical examples include GMAN (Zheng et al., 2020) and ST-Wave (Fang et al., 2023).
+Early research, such as STGCN (Yu_2018), STSGCN (song2020spatial) and GGRNN (ruiz2020gated), employs Graph Convolutional Networks (GCNs) to capture spatial relations and uses 1D convolutions and RNNs to process temporal relationships.
+ASTGNN (ta2022adaptive), Graph WaveNet (10.5555/3367243.3367303) and AGCRN (bai19) explored adaptive graph settings to capture potential non-local signal relations over the graph, trading in interpretability for improved performance.
+Graph Attention Networks (GATs) learn graph weights from embedded signals to enable flexible and data-targeted graph learning, typical examples include GMAN (zheng2020gman) and ST-Wave (fang2023spatio).
 All GNNs above focus only on the spatial correlations, while the product graph (PG) is an intuitive model that describes local spatial-temporal relations at the same time.
-Einizade et al.
-(2024) presents a continuous product graph neural network for spatial-temporal forecasting, which achieves superior performance to the previous most-popular graph methods.
+einizade2024continuous presents a continuous product graph neural network for spatial-temporal forecasting, which achieves superior performance to the previous most-popular graph methods.
 Transformers extend the powerful self-attention mechanism to all spatial-temporal data points, which enhances model performance with an extra-large number of parameters.
-Popular transformer-based models for traffic forecasting employ two transformers to process spatial and temporal dimensions separately (Xu et al., 2020; Chen et al., 2022; Huo et al., 2023), while state-of-the-art transformers introduce adaptive local attentions (Liu et al., 2023), time-delay (Jiang et al., 2023), or pre-training (Gao et al., 2024).
+Popular transformer-based models for traffic forecasting employ two transformers to process spatial and temporal dimensions separately (xu2020spatial; chen2022bidirectional; huo2023hierarchical), while state-of-the-art transformers introduce adaptive local attentions (liu2023spatio), time-delay (jiang2023pdformer), or pre-training (gao2024spatial).
 In recent years, simple linear models have been proposed to challenge the effectiveness and efficiency of heavy transformers.
-DLinear (Zeng et al., 2023) simply separates the input signals as trends and residuals, and employs 2 simple linear layers shared across nodes for prediction, but beats some of the transformer baselines in long-term forecasting.
-Other representative simple linear models include STID (Shao et al., 2022) and SimpleTM Chen et al.
-(2025).
+DLinear (Zeng2022AreTE) simply separates the input signals as trends and residuals, and employs 2 simple linear layers shared across nodes for prediction, but beats some of the transformer baselines in long-term forecasting.
+Other representative simple linear models include STID (10.1145/3511808.3557702) and SimpleTM (chen2025simpletm).
 As shown in Table 2, we select two most representative and well-performing models from each category as baselines for comparison.
+A.3 Studies on Directed Graph Laplacians
+In the DGSP literature, frequencies of a general graph (directed or undirected) were first defined in Sandryhaila2014 via Jordan decomposition of the adjacency matrix, which is known to be numerically unstable.
+In Shafipour2019, a set of orthornormal vectors (deemed directed graph frequencies) are computed via a minimization of spectral dispersion in a constrained quadratic programming formulation in eq.
+(8), which is computationally expensive.
+In Marques2020, a graph shift operator (GSO) $\mathbf{S}$ is assumed to be diagonalizable into $\mathbf{S}=\mathbf{V}\text{diag}(\{\lambda_{k}\})\mathbf{V}^{-1}$, where the generally non-orthogonal eigenvectors in $\mathbf{V}$ are deemed frequency modes and ordered using a total variation measure in eq.
+(2), given eigenvalues $\lambda_{k}$’s can be complex-valued.
+However, asymmetric adjacency and graph Laplacian matrices are not always diagonalizable via the Spectral Theorem.
+Seifert2021 proposes to augment a directed graph with extra edges to destroy nontrivial Jordan blocks, so that the resulting adjacency matrix becomes diagonalizable.
+Adding edges, however, fundamentally changes the graph representation of the original pairwise relationships.
+Chen2023Graph proposes to use singular value decomposition (SVD) on the directed Laplacian to derive frequencies for directed graphs, which is computationally expensive.
+Kwak2024 proposes to decompose the asymmetric Laplacian matrix into a unitary part (capturing rotation or cycles) and a positive semi-definite part (capturing diffusion-like behavior) also using SVD and polar decomposition, thus suffering the same computation burden.
+None of these related works proposed variational terms to define directed graph frequencies, which in turn can be simply added to an objective to promote signal smoothness during optimization without first computing frequency components, as we have done.
 Appendix B Proof of Theorem 3.1
@@
 ${\mathbf{W}}^{d}_{r}=\left[\begin{array}[]{cccc}1&0&0&0\\ 1&0&0&0\\ 0&1&0&0\\ 0&0&1&0\end{array}\right],~{\mathbf{L}}^{d}_{r}=\left[\begin{array}[]{cccc}0&0&0&0\\ -1&1&0&0\\ 0&-1&1&0\\ 0&0&-1&1\end{array}\right],{\mathcal{L}}^{d}_{r}=\left[\begin{array}[]{cccc}1&-1&0&0\\ -1&2&-1&0\\ 0&-1&2&-1\\ 0&0&-1&1\end{array}\right].$ (54)
-Appendix C Interpreting Linear System eq. 10
+Appendix C Interpreting Linear System Equation 10
 For notation simplicity, we first define
-${\mathbf{r}}^{\tau}\triangleq({\mathbf{L}}^{d}_{r})^{\top}(\frac{{\bm{\gamma}}^{\tau}}{2}+\frac{\rho}{2}{\boldsymbol{\phi}}^{\tau})-\frac{{\bm{\gamma}}_{u}^{\tau}}{2}+\frac{\rho_{u}}{2}{\mathbf{z}}_{u}^{\tau}-\frac{{\bm{\gamma}}_{d}^{\tau}}{2}+\frac{\rho_{d}}{2}{\mathbf{z}}_{d}^{\tau}.$ (54)
-Solution ${\mathbf{x}}^{\tau+1}$ to linear system eq. 10 can now be written as
+${\mathbf{r}}^{\tau}\triangleq\small({\mathbf{L}}^{d}_{r})^{\top}(\frac{{\boldsymbol{\gamma}}^{\tau}}{2}+\frac{\rho}{2}{\boldsymbol{\phi}}^{\tau})-\frac{{\boldsymbol{\gamma}}_{u}^{\tau}}{2}+\frac{\rho_{u}}{2}{\mathbf{z}}_{u}^{\tau}-\frac{{\boldsymbol{\gamma}}_{d}^{\tau}}{2}+\frac{\rho_{d}}{2}{\mathbf{z}}_{d}^{\tau}.$ (55)
+Solution ${\mathbf{x}}^{\tau+1}$ to linear system Equation 10 can now be written as
 ${\mathbf{x}}^{\tau+1}$ $=\left({\mathbf{H}}^{\top}{\mathbf{H}}+\frac{\rho}{2}{\mathcal{L}}^{d}_{r}+\frac{\rho_{u}+\rho_{d}}{2}{\mathbf{I}}\right)^{-1}({\mathbf{r}}^{\tau}+{\mathbf{H}}^{\top}{\mathbf{y}}).$ (56)
@@
 which demonstrates that ${\mathbf{x}}^{\tau+1}$ is a low-pass filter output of up-sampled ${\mathbf{H}}^{\top}{\mathbf{y}}$ (with bias ${\mathbf{r}}^{\tau}$).
-Appendix D Deriving Solution to Optimization eq. 15
-The optimization for $\boldsymbol{\phi}$} in objective eq. 15 can be performed entry-by-entry. Specifically, for the $i$-th entry $\phi_{i}$:
+Appendix D Deriving Solution to Optimization Equation 15
+The optimization for $\boldsymbol{\phi}$} in objective Equation 15 can be performed entry-by-entry. Specifically, for the $i$-th entry $\phi_{i}$:
 $\min_{\phi_{i}}~~g(\phi_{i})=\mu_{d,1}|\phi_{i}|+\gamma_{i}^{\tau}\left(\phi_{i}-({\mathbf{L}}^{d}_{r})_{i}{\mathbf{x}}^{\tau+1}\right)+\frac{\rho}{2}\left(\phi_{i}-({\mathbf{L}}^{d}_{r})_{i}{\mathbf{x}}^{\tau+1}\right)^{2}$ (58)
@@
 $\phi_{i}^{*}$ $=\text{sign}(\delta)\cdot\max(|\delta|-\rho^{-1}\mu_{d,1},\,0)$ (61)
-Alternative solution with subgradients: Denote $g_{1}({\boldsymbol{\phi}})=\mu_{d,1}\|{\boldsymbol{\phi}}\|_{1},g_{2}({\boldsymbol{\phi}})=({\bm{\gamma}}^{\tau})^{\top}({\boldsymbol{\phi}}-{\mathbf{L}}^{d}_{r}{\mathbf{x}}^{\tau+1})+\frac{\rho}{2}\|{\boldsymbol{\phi}}-{\mathbf{L}}^{d}_{r}{\mathbf{x}}^{\tau+1}\|_{2}^{2}$, the subgradient of $g({\boldsymbol{\phi}})$ is
-$\partial g({\boldsymbol{\phi}})=\{{\mathbf{g}}+\nabla g_{2}({\boldsymbol{\phi}})|{\mathbf{g}}\in\partial g_{1}({\boldsymbol{\phi}})\}=\{\mu_{d,1}\sigma+{\bm{\gamma}}^{\tau}+\rho({\boldsymbol{\phi}}-{\mathbf{L}}^{d}_{r}{\mathbf{x}}^{\tau+1})|\sigma\in\partial(\|{\boldsymbol{\phi}}\|_{1})\},$ (61)
+Alternative solution with subgradients: Denote $g_{1}({\boldsymbol{\phi}})=\mu_{d,1}\|{\boldsymbol{\phi}}\|_{1},g_{2}({\boldsymbol{\phi}})=({\boldsymbol{\gamma}}^{\tau})^{\top}({\boldsymbol{\phi}}-{\mathbf{L}}^{d}_{r}{\mathbf{x}}^{\tau+1})+\frac{\rho}{2}\|{\boldsymbol{\phi}}-{\mathbf{L}}^{d}_{r}{\mathbf{x}}^{\tau+1}\|_{2}^{2}$, the subgradient of $g({\boldsymbol{\phi}})$ is
+$\partial g({\boldsymbol{\phi}})=\{{\mathbf{g}}+\nabla g_{2}({\boldsymbol{\phi}})|{\mathbf{g}}\in\partial g_{1}({\boldsymbol{\phi}})\}=\{\mu_{d,1}\sigma+{\boldsymbol{\gamma}}^{\tau}+\rho({\boldsymbol{\phi}}-{\mathbf{L}}^{d}_{r}{\mathbf{x}}^{\tau+1})|\sigma\in\partial(\|{\boldsymbol{\phi}}\|_{1})\},$ (62)
 where
@@
 (65)
-Thus the solution of eq. 15 is
+Thus the solution of Equation 15 is
 ${\boldsymbol{\phi}}^{\tau+1}={\boldsymbol{\phi}}^{*}=\text{soft}_{\frac{\mu_{d,1}}{\rho}}\left({\mathbf{L}}^{d}_{r}{\mathbf{x}}-\frac{\gamma_{i}^{\tau}}{\rho}\right)=\text{sign}\left({\mathbf{L}}^{d}_{r}{\mathbf{x}}-\frac{\gamma_{i}^{\tau}}{\rho}\right)\max\left(\left|{\mathbf{L}}^{d}_{r}{\mathbf{x}}-\frac{\gamma_{i}^{\tau}}{\rho}\right|-\frac{\mu_{d,1}}{\rho},0\right).$ (66)
-Appendix E Parameterizing Conjugated Gradient Method in Solving Linear Equations
+Appendix E Algorithms for the Unrolled Network
+E.1 Parameterizing Conjugated Gradient Method in Solving Linear Equations
 A linear system ${\mathbf{A}}{\mathbf{x}}={\mathbf{b}}$ where $\mathbf{A}$} is a PD matrix can be solved iteratively by the conjugated gradient (CG) method, as is described in Algorithm 1.
@@
 Note that $\alpha,\beta\geq 0$ in every iteration in the original algorithm, and should be kept non-negative in the unrolled version.
-Algorithm 1 Conjugate Gradient (CG) Method to Solve ${\mathbf{A}}{\mathbf{x}}={\mathbf{b}}$
+Algorithm 1 Conjugate Gradient (CG) Method to Solve ${\mathbf{A}}{\mathbf{x}}={\mathbf{b}}$ (Unrolled version uCG($\mathbf{A}$}, $\mathbf{b}$}, ${\mathbf{x}}_{0}$))
 1:PSD matrix ${\mathbf{A}}\in\mathbb{R}^{n\times n}$, vector ${\mathbf{b}}\in\mathbb{R}^{n}$, initial guess ${\mathbf{x}}_{0}$
@@
 6:while $\|{\mathbf{A}}{\mathbf{x}}_{k}-{\mathbf{b}}\|>\epsilon$ do
-7:  $\boxed{\alpha_{k}\leftarrow\dfrac{{\mathbf{r}}_{k}^{\top}{\mathbf{r}}_{k}}{{\mathbf{p}}_{k}^{\top}{\mathbf{A}}{\mathbf{p}}_{k}}}$
+7:  $\boxed{\alpha_{k}\leftarrow\dfrac{{\mathbf{r}}_{k}^{\top}{\mathbf{r}}_{k}}{{\mathbf{p}}_{k}^{\top}{\mathbf{A}}{\mathbf{p}}_{k}}}$ $\triangleright$ Learnable in unrolling
 8:  ${\mathbf{x}}_{k+1}\leftarrow{\mathbf{x}}_{k}+\alpha_{k}{\mathbf{p}}_{k}$
 9:  ${\mathbf{r}}_{k+1}\leftarrow{\mathbf{r}}_{k}-\alpha_{k}{\mathbf{A}}{\mathbf{p}}_{k}$ $\triangleright$ Dual residual
-10:  $\boxed{\beta_{k}\leftarrow\dfrac{{\mathbf{r}}_{k+1}^{\top}{\mathbf{r}}_{k+1}}{{\mathbf{r}}_{k}^{\top}{\mathbf{r}}_{k}}}$
+10:  $\boxed{\beta_{k}\leftarrow\dfrac{{\mathbf{r}}_{k+1}^{\top}{\mathbf{r}}_{k+1}}{{\mathbf{r}}_{k}^{\top}{\mathbf{r}}_{k}}}\triangleright$ Learnable in unrolling
 11:  ${\mathbf{p}}_{k+1}\leftarrow{\mathbf{r}}_{k+1}+\beta_{k}{\mathbf{p}}_{k}$
@@
 14:return ${\mathbf{x}}_{k+1}$
+E.2 Detailed Algorithms and Scalability Analysis for the Proposed Model
+Algorithms 2 detailedly explained the operations in ADMM blocks corresponding to Figure 3(b) and (c).
+An overall algorithm of the unrolled network corresponding to Figure 3(a) is shown in Algorithm 3.
+All matrix multiplication is computed locally to ensure low computational and memory cost with
+$({\mathbf{L}}^{u}{\mathbf{x}})_{j}=$ $D_{j,j}^{u}x_{j}-\sum_{i\in{\mathcal{N}}_{j}}W_{i.j}x_{j},\quad({\mathbf{L}}^{u}_{r}{\mathbf{x}})_{j}=x_{j}-\sum_{i\in{\mathcal{N}}_{i}}\frac{W_{i,j}}{\sqrt{D_{i,i}^{u}D_{j,j}^{u}}}x_{i},$ (67)
+$({\mathbf{L}}^{d}_{r}{\mathbf{x}})_{j}=$ $x_{j}-\sum_{(i,j)\in{\mathcal{E}}^{d}}\frac{W_{j,i}^{d}}{D_{j,j}^{d}}x_{i},\quad(({\mathbf{L}}^{d}_{r})^{\top}{\mathbf{x}})_{j}=x_{j}-\sum_{(j,i)\in{\mathcal{E}}^{d}}\frac{W_{i,j}^{d}}{D_{i,i}^{d}}x_{i},\quad{\mathcal{L}}^{d}_{r}{\mathbf{x}}=({\mathbf{L}}^{d}_{r})^{\top}{\mathbf{L}}^{d}_{r}{\mathbf{x}}.$ (68)
+Algorithm 2 Forward Propagation of the Unrolled ADMM_Block
+1:Initial signal output of the last ADMM block ${\mathbf{x}}\in\mathbb{R}^{T+S+1}$, undirected graph Laplacian matrix ${\mathbf{L}}^{u}$, directed graph Laplacian matrix ${\mathbf{L}}^{d}_{r}$.
+2:Network output of the unrolled ADMM block with $M$ layers.
+[…]
+4:Select real observations with mask: ${\mathbf{y}}\leftarrow{\mathbf{H}}{\mathbf{x}},{\mathbf{y}}\in\mathbb{R}^{T+1}$.
+5:for $\tau$ in $0:M$ do $\triangleright$ $M$ ADMM layers
+6:  Initialize ${\boldsymbol{\phi}}\leftarrow{\mathbf{L}}^{d}_{r}{\mathbf{x}}$
+7:  ${\mathbf{x}}\leftarrow\texttt{uCG}({\mathbf{H}}^{\top}{\mathbf{H}}+\frac{\rho^{\tau}}{2}{\mathcal{L}}^{d}_{r}+\frac{\rho_{u}^{\tau}+\rho_{d}^{\tau}}{2}{\mathbf{I}},({\mathbf{L}}^{d}_{r})^{\top}\!(\frac{{\boldsymbol{\gamma}}}{2}+\frac{\rho}{2}{\boldsymbol{\phi}})-\frac{{\boldsymbol{\gamma}}_{u}}{2}+\frac{\rho_{u}^{\tau}}{2}{\mathbf{z}}_{u}-\frac{{\boldsymbol{\gamma}}_{d}}{2}+\frac{\rho_{d}^{\tau}}{2}{\mathbf{z}}_{d}+{\mathbf{H}}^{\top}{\mathbf{y}},{\mathbf{x}})$
+8:  ${\mathbf{z}}_{u}\leftarrow\texttt{uCG}(\mu_{u}^{\tau}{\mathbf{L}}^{u}+\frac{\rho_{u}^{\tau}}{2}{\mathbf{I}},\frac{{\boldsymbol{\gamma}}_{u}}{2}+\frac{\rho_{u}^{\tau}}{2}{\mathbf{x}},{\mathbf{z}}_{u})$
+9:  ${\mathbf{z}}_{d}\leftarrow\texttt{uCG}(\mu_{d,2}^{\tau}{\mathcal{L}}^{d}_{r}+\frac{\rho_{d}^{\tau}}{2}{\mathbf{I}},\frac{{\boldsymbol{\gamma}}_{d}}{2}+\frac{\rho_{d}^{\tau}}{2}{\mathbf{x}},{\mathbf{z}}_{d})\triangleright$ Minimize $\mathbf{x}$} with Equations 10, 11 and 12
+10:  ${\boldsymbol{\phi}}\leftarrow\text{soft}_{\mu_{d,1}^{\tau}/\rho^{\tau}}({\mathbf{L}}^{d}_{r}{\mathbf{x}}-{\boldsymbol{\gamma}}/\rho^{\tau})$ $\triangleright$ Minimize $\phi$ with Equation 16
+11:  ${\boldsymbol{\gamma}}\leftarrow{\boldsymbol{\gamma}}+\rho^{\tau}({\boldsymbol{\phi}}^{\tau+1}-{\mathbf{L}}^{d}_{r}{\mathbf{x}}^{\tau+1})$
+12:  ${\boldsymbol{\gamma}}_{u}\leftarrow{\boldsymbol{\gamma}}_{u}+\rho_{u}^{\tau}({\mathbf{x}}-{\mathbf{z}}_{u})$
+13:  ${\boldsymbol{\gamma}}_{d}\leftarrow{\boldsymbol{\gamma}}_{d}+\rho_{d}^{\tau}({\mathbf{x}}-{\mathbf{z}}_{d})$ $\triangleright$ Update multipliers with Equations 17 and 18
+14:end for
+15:return $\mathbf{x}$}
+Algorithm 3 Overall Algorithm of the Proposed Unrolled Network
+1:An observation from the training set ${\mathbf{y}}\in\mathbb{R}^{T+1}$.
+2:The reconstructed signal ${\mathbf{x}}\in\mathbb{R}^{T+S+1}$ containing both the obseration and prediction.
+[…]
+[…]
+[…]
+[…]
+7:for $i$ in $0:B$ do $\triangleright$ $B$ ADMM blocks and graph learning modules
+8:  ${\mathbf{e}}_{i}\leftarrow[{\mathbf{x}}_{i};{\mathbf{e}}_{\text{st},i}],{\mathbf{f}}_{i}\leftarrow F^{u}({\mathbf{e}}_{i})$ for $i=1,\dots,N$ $\triangleright$ Feature extraction (Appendix F.3.2)
+9:  ${\mathbf{L}}^{u(h)}\leftarrow\texttt{UGL}_{h}({\mathbf{f}}_{1},\dots,{\mathbf{f}}_{N}),\quad{\mathbf{L}}^{d(h)}_{r}\leftarrow\texttt{DGL}_{h}({\mathbf{f}}_{1},\dots,{\mathbf{f}}_{N}),h=1,2,\dots,H$
+[…]
+11:  ${\mathbf{x}}_{\text{new}}^{(h)}\leftarrow\texttt{ADMM\_Block}_{h}({\mathbf{x}},{\mathbf{L}}^{u(h)},{\mathbf{L}}^{d(h)}_{r}),h=1,2,\dots,H$
+12:  ${\mathbf{x}}_{\text{new}}=\sum_{h=1}^{H}a_{h}{\mathbf{x}}_{\text{new}}^{(h)}$ $\triangleright$ Learnable merge of multi-head results
+13:  ${\mathbf{x}}\leftarrow p_{i}{\mathbf{x}}_{\text{new}}+(1-p_{i}){\mathbf{x}}$ $\triangleright$ Residual connections with learnable parameter $p_{i}$
+14:end for
+15:return $\mathbf{x}$}
+E.3 Scability Analysis
+our model scales linearly as long as the city map is sparse.
+The conjugated gradient algorithm is detailed in Appendix E, which involves only sparse matrix-vector multiplications in each iteration.
+Specifically, each matrix multiplication involves graph Laplacians describing pairwise spatial and temporal similarities in local neighborhoods.
+Thus, as long as the temporal window $W$ and number of neighbors $k$ do not change, the computation cost in unrolled CG grows only linearly with the increase of total time steps and number of nodes.
+Our graph learning modules also calculate and save the edge weights locally in each node’s neighborhood, as is demonstrated in Equation 68.
+Taken together, with fixed $W$ and $k$, the size of our model, computation time, and memory cost all grow linearly with the total time steps and network size.
 Appendix F Model Setup
@@
 Those parameters are trained every ADMM layer.
-We initialize the CGD parameters ($\bm{\alpha}$}, $\bm{\beta}$})s as 0.08, and tune them for every CGD iteration.
-We clamp each element of the CGD step size $\bm{\alpha}$} to [0, 0.8] for stability.
-Each element of the momentum term coefficient $\bm{\beta}$} is clamped to be non-negative.
+We initialize the CGD parameters ($\boldsymbol{\alpha}$}, $\boldsymbol{\beta}$})s as 0.08, and tune them for every CGD iteration.
+We clamp each element of the CGD step size $\boldsymbol{\alpha}$} to [0, 0.8] for stability.
+Each element of the momentum term coefficient $\boldsymbol{\beta}$} is clamped to be non-negative.
 F.2 Graph Learning Modules Setup
@@
 F.3 Feature Extractors Setup
-Input Laplacian embeddings Position embeddings Concatenate embedded input
+Input Spatial embeddings Temporal embeddings Concatenate embedded input
 (a) Input embedding layer.
@@
 F.3.1 Input Embeddings
-We generate non-parametric spatial and temporal embeddings for each node in our mixed spatio-temporal graph, see Figure 4(a).
-For spatial embeddings, we first generate the adjacency matrix ${\mathbf{W}}^{s}$ for the real-world connection of monitor stations using the physical distances $\text{d}^{S}(i,j)$ with
-$w_{i,j}^{S}=w_{j,i}^{S}=\exp\left[-\frac{(\text{d}^{S}(i,j))^{2}}{50\times\text{Var}(\text{d}^{S})}\right],$ (66)
-where $\text{Var}(\text{d}^{S})$ denotes the variance of the physical distances.
-We use Laplacian embedding supposed by Dwivedi & Bresson (2020), which uses the eigenvectors of the $k^{S}$-smallest non-zero eigenvalues of the Laplacian matrix ${\mathbf{L}}^{s}={\mathbf{D}}^{s}-{\mathbf{W}}^{s}$ to generate $k$-dimensional embeddings.
-Our experiments set $k^{S}=5$.
-For temporal embeddings, we use the sine-cosine position embeddings (Vaswani et al., 2018):
+We generate low-parametric spatial and temporal embeddings for each node in our mixed spatio-temporal graph, see Figure 7(a).
+For spatial embeddings, we embed each monitoring station with a 5-dimensional learnable vector, and broadcast the embeddings to every time step in the product graph.
+For temporal embeddings, we follow feng2022 to use real-time-stamp, time-in-day and day-in-week embeddings.
+To achieve better fitting capacity while keeping the model size small, we set only the time-in-day and day-in-week embeddings learnable, and use fixed embeddings for real time stamps.
+We use a fixed 10-dimensional sine-cosine position embeddings (tensor2tensor) for each time stamp:
 $e[t,2i]=$ $\sin(t/10000^{i})$
-$e[t,2i+1]=$ $\cos(t/10000^{i}),i=0,1,\dots,\left\lfloor\frac{k^{T}}{2}\right\rfloor-1.$ (67)
-We follow Feng & Tassiulas (2022) to use real-time-stamp, time-in-day and day-in-week embeddings of dimensions 10, 6, 4.
+$e[t,2i+1]=$ $\cos(t/10000^{i}),i=0,1,2,3,4,$ (69)
+and set dimensions of learnable time-in-day and day-in-week embeddings as 6 and 4. All embeddings are shared across all graph learning modules.
 We directly concatenate these spatial and temporal embedding to our signals ${\mathbf{x}}^{t}_{j}$ for node $j$ at time $t$ as ${\mathbf{e}}_{j}^{t}=[{\mathbf{x}}^{t}_{j};{\mathbf{e}}^{S}_{j};{\mathbf{e}}^{T}_{t}]$ for feature extractors, where ${\mathbf{e}}^{S}_{j}$ denote the spatial Laplacian embeddings of node $j$ of the physical graph, and ${\mathbf{e}}^{T}_{t}$ denotes the temporal position embeddings of time step $t$.
 F.3.2 Feature Extraction from embeddings
-We use a variant of GraphSAGE (Hamilton et al., 2017) to aggregate spatial features and introduce Temporal History Aggregation (THA) to aggregate spatial features.
+We use a variant of GraphSAGE (hamilton2017inductive) to aggregate spatial features and introduce Temporal History Aggregation (THA) to aggregate spatial features.
 The GraphSAGE module aggregates the inputs of each node’s $k$-NNs with a simple and uniformed linear layer to generate $K$-dimension spatial features for each node of each time step.
 The THA module aggregates the embedded inputs of past $W$ time steps for each node with a uniformed linear layer, where $W$ is the size of the time window.
-For both GraphSAGE and THA modules, we pad zeros to inputs as placeholders, and use Swish (Ramachandran et al., 2017) as acitvation functions with $\beta=0.8$.
+For both GraphSAGE and THA modules, we pad zeros to inputs as placeholders, and use Swish (ramachandran2017swish) as acitvation functions with $\beta=0.8$.
 F.4 Constructing Graphs ${\mathbf{W}}^{u,0},{\mathbf{W}}^{d,0}$ for the First ADMM Block
@@
 Table 4: Training setup of graph construction and batch size for all traffic datasets.
-Dataset PeMS03 PeMS04 METR-LA
+Dataset PeMS03 PeMS08 METR-LA
 $k$ 4 6 6
 $W$ 6 6 6
-Batch size 12 12 16
-Appendix G Training Performance w.r.t. Training Data Cost
+Batch size 12 16 16
+Appendix G Further Experiments
+G.1 Additional Experimental Results
+Table 5 shows the performance comparison between our model and baselines in 30-/60-minute forecast on the PEMS08 (9346058) dataset which contains 170 nodes and 295 edges.
+Results show that our model also ranks top-3 in at least one metric of all experiments, and even gives the smallest MAPE among all models.
+Table 5: Experimental results on additional dataset PEMS08 for 30- and 60-minute forecast.The indication of bolded, highlighted and underlined entries is the same as Table 1.
+Horizons 30 minutes (6 steps) 60 minutes (12 steps)
+Model RMSE MAE MAPE (%) RMSE MAE MAPE (%)
+VAR 26.81 17.74 12.69 28.59 18.99 13.52
+STGCN 36.58 24.39 16.83 50.61 33.01 20.57
+STSGCN 26.73 17.43 11.48 30.22 19.74 12.90
+GMAN 24.84 15.73 10.87 26.22 16.83 13.64
+ST-Wave 26.85 17.19 11.08 27.28 16.85 11.77
+PDFormer 22.92 13.91 9.17 27.36 17.61 12.90
+STAEFormer 26.54 16.62 10.88 32.68 20.78 14.86
+Graph WaveNet 30.95 21.65 13.56 40.24 29.12 17.79
+AGCRN 34.31 19.91 10.88 38.09 22.78 12.75
+SITD 27.40 18.11 12.10 33.12 22.66 20.08
+SimpleTM 24.44 15.07 9.09 32.48 20.26 12.34
+Ours 25.92 16.22 10.17 27.78 17.22 11.11
+G.2 Experiments on the Entire Datasets
+To further strengthen our proposed method, we ran the 60-minute forecast experiment on the PEMS03 dataset without sampling with 70 epochs of training, and the results are shown in Table 6.
+Our unrolled network still ranks top-3 in at least one of the metrics for each task, indicating the consistency of our performance with subsampling.
+Table 6: Performance comparison of our model to the baselines on a 60-minute forecast for the entire PEMS03 dataset.
+The indication of bolded, highlighted and underlined entries is the same as Table 1.
+Dataset PEMS03 METR-LA
+Model RMSE MAE MAPE (%) RMSE MAE MAPE (%)
+VAR 30.53 18.30 19.70 12.57 7.05 13.48
+STGCN 36.77 21.81 19.37 12.27 5.88 11.79
+STSGCN 28.17 17.46 16.86 12.73 5.11 10.66
+PDFormer 25.05 14.78 15.53 12.32 4.77 10.20
+GMAN 28.44 17.04 22.82 13.75 6.44 12.72
+GWN 28.68 17.08 16.49 13.82 6.41 11.92
+AGCRN 27.65 15.77 15.02 12.18 4.74 10.50
+STID 26.80 15.14 16.45 12.00 4.63 10.14
+SimpleTM 24.51 15.32 15.55 12.86 5.67 9.89
+Ours 26.71 16.36 17.18 12.22 5.35 11.87
+G.3 Inference-time Computational and Memory Cost
+We compute the inference-time computational and memory cost on 60-minute forecasting for the PEMS03 dataset.
+We calculate the total floating-point operations (FLOPs) during inference as the computational cost with a batch size of 1, and evaluate the peak allocated memory by CUDA for each batch of size 32 as the memory cost.
+Results in Table 7 show that our unrolled network has the least computation cost among all the baselines except AGCRN and STID, which are generally consistently outperformed by our model.
+Our model’s inference-time memory cost is only larger than SimpleTM and the two baselines mentioned above.
+In conclusion, our lightweight transformer reduces the computational cost to only 4.9% of transformer-based PDFormer and 11.2% of the GAT-based GMAN with reduced or at least comparable memory cost.
+Table 7: Comparison of computational cost (GFLOPs) for each data point and inference memory cost (GB) for a batch of size 32 in 60-minute forecasting for the PEMS03 dataset.
+Model Ours STGCN GMAN ST-Wave PDFormer
+Computation (GFLOPS) 0.087 0.013 $\times 12^{\dagger}$ 0.777 4.496 1.771
+[…]
+Model STAEFormer GWN AGCRN STID SimpleTM
+[…]
+[…]
+    The computational and memory cost for STSGCN is not included due to the limitation of applying tools such as thop in the public code.
+  • $\dagger$
+    The $\times 12$ notation in the table means that the model predicts the outputs recursively.
+G.4 Training Performance w.r.t. Training Data Cost
 Due to the parameter efficiency of our model, our unrolling model is also data-efficient under the same observation length.
@@
 Model Stride = 3 (original) Stride = 5 Stride = 10
-Ours 27.67 / 17.46 / 17.72 30.42 / 18.60 / 20.05 30.65 / 19.02 / 19.67
+Ours 26.96 / 16.58 / 18.07 28.98 / 17.73 / 18.63 29.63 / 18.58 / 19.76
 VAR 30.54 / 18.31 / 19.61 30.68 / 18.37 / 19.74 31.25 / 18.71 / 21.22
@@
 AGCRN 29.90 / 16.76 / 15.32 39.25 / 20.60 / 16.95 62.58 / 109.27 / 30.72
-Appendix H Ablation Studies
-H.1 Ablation Studies for $\ell_{2}$- and $\ell_{1}$- norms in Direct Graph Modeling
-First, we discuss the effectiveness of the new directed graph priors, DGLR and DGTV, as well as the benefit of directed-graph modeling of temporal relations.
-We separately test the effects of DGLR and DGTV by removing the prior and unrolling the ADMM algorithm.
-We denote these models as “w/o DGTV” and “w/o DGLR”.
-To test the benefit of the directed-graph modeling, we change the temporal graph ${\mathcal{G}}^{d}$ into a fully undirected graph ${\mathcal{G}}^{n}$ while maintaining the connections, and replace the directed graph priors with a GLR prior ${\mathbf{x}}^{\top}{\mathbf{L}}^{n}{\mathbf{x}}$ with the same coefficient $\mu_{n}=\mu_{d,2}$.
-We denote the undirected-time unrolling model as “UT”.
-Detailed derivations of the ADMM algorithm for all the above ablation studies are in Appendix I.
-Table 5: Performance comparison in RMSE / MAE / MAPE(%) of the ablation studies of directed-graph modeling and directed graph priors in 60-minute forecasting with PEMS03 and METR-LA.
-Dataset Mixed-Graph UT w/o DGTV w/o DGLR
-PEMS03 27.67 / 17.46 / 17.72 27.52 / 17.87 / 18.82 27.78 / 17.85 / 17.90 30.89 / 20.02 / 21.10
-METR-LA 12.34 / 5.18 / 11.80 12.51 / 5.42 / 12.11 12.36 / 5.40 / 12.31 12.41 / 5.35 / 12.20
-As is summarized in Table 5, removing the DGLR or DGTV term leads to noticeable drops in performance, which underscores their importance as directed graph priors.
-Moreover, our model outperforms the fully-undirected-graph-based model, demonstrating that sequential relationships—such as time—are better modeled as directed graphs.
-H.2 Sensitivity of $k$, $W$ in Graph Construction
+Appendix H Further Ablation Studies for Hyperparameter Sensitivity
+H.1 Sensitivity of $k$, $W$ in Graph Construction
 We change the number of spatial nearest neighbors $k=4,6,8$ and time window size $W=4,6,8$ and present the results for PEMS03 and METR-LA dataset in Table 9.
@@
 $k$ $W$ RMSE MAE MAPE(%) $k$ $W$ RMSE MAE MAPE(%)
- 4*  6* 27.67 17.46 17.72  6*  6* 12.37 5.27 11.91
-4 4 29.54 17.09 17.92 4 6 12.43 5.24 11.87
-4 8 27.45 17.25 18.35 8 6 12.61 5.30 11.92
-6 6 28.98 17.33 18.23 6 4 12.44 5.40 11.85
-8 6 45.37 18.15 20.59 6 8 12.44 5.35 11.96
-H.3 Sensitivity of Number of Blocks and Layers in the Unrolling Model
+ 4*  6* 26.96 16.58 18.07  6*  6* 12.17 5.27 11.78
+4 4 27.23 16.80 17.73 4 6 12.28 5.24 11.86
+4 8 27.30 17.34 19.34 8 6 12.45 5.26 11.91
+6 6 29.57 18.00 19.43 6 4 12.38 5.23 11.84
+8 6 28.40 17.99 18.48 6 8 12.52 5.42 12.29
+H.2 Sensitivity of Number of Blocks and Layers in the Unrolling Model
 We vary the number of blocks and ADMM layers for PEMS03 dataset and present the results in Table 10.
@@
 RMSE MAE MAPE(%) RMSE MAE MAPE(%)
- 5*  25* 27.67 17.46 17.72 12.34 5.18 11.80
-5 20 27.95 17.32 18.02 12.53 5.35 12.15
-5 30 27.86 17.73 18.80 12.54 5.31 11.94
-4 25 27.77 17.63 18.61 12.45 5.33 12.00
-6 25 27.84 17.54 18.61 12.37 5.27 11.90
-H.4 Ablation Study for the Splitting-$\ell_{2}$-term Strategy
-We also ran ablations of the splitting-term strategy by optimizing ${\mathbf{x}}^{\tau+1}$ using only eq.
-7.
-CG in solving eq.
-7 converges much slower than in solving eq.
-10-eq.
-12, and the unrolled model can become numerically unstable in training, indicating that splitting the $\ell_{2}$ terms can stabilize the training process.
+ 5*  25* 26.96 16.58 18.07 12.17 5.27 11.78
+5 20 27.52 17.36 18.97 12.51 5.30 11.99
+5 30 27.25 17.27 18.77 12.51 5.29 11.96
+4 25 27.84 17.23 17.96 12.43 5.32 11.93
+6 25 26.97 16.92 17.97 12.46 5.25 11.80
 Appendix I Optimization Formulation and ADMM Algorithm in Ablation Studies
 I.1 Optimization & Algorithm without DGTV Term
-By removing the DGTV term $\|{\mathbf{L}}^{d}_{r}{\mathbf{x}}\|_{1}$ in eq.
-4, and introducing the auxiliary variables ${\mathbf{z}}_{u},{\mathbf{z}}_{d}$ as eq.
-9, we rewrite the unconstrained version of the optimization by augmented Lagrangian method as
-$\min_{{\mathbf{x}},{\mathbf{z}}_{u},{\mathbf{z}}_{d}}$ $\|{\mathbf{y}}-{\mathbf{H}}{\mathbf{x}}\|_{2}^{2}+\mu_{u}{\mathbf{z}}_{u}^{\top}{\mathbf{L}}^{u}{\mathbf{z}}_{u}+\mu_{d,2}{\mathbf{z}}_{d}^{\top}{\mathcal{L}}^{d}_{r}{\mathbf{z}}_{d}+({\bm{\gamma}}_{u}^{\tau})^{\top}({\mathbf{x}}-{\mathbf{z}}_{u})+\frac{\rho_{u}}{2}\|{\mathbf{x}}-{\mathbf{z}}_{u}\|_{2}^{2}$
-$+({\bm{\gamma}}_{d}^{\tau})^{\top}({\mathbf{x}}-{\mathbf{z}}_{d})+\frac{\rho_{d}}{2}\|{\mathbf{x}}-{\mathbf{z}}_{d}\|_{2}^{2}.$ (68)
+By removing the DGTV term $\|{\mathbf{L}}^{d}_{r}{\mathbf{x}}\|_{1}$ in Equation 4, and introducing the auxiliary variables ${\mathbf{z}}_{u},{\mathbf{z}}_{d}$ as Equation 9, we rewrite the unconstrained version of the optimization by augmented Lagrangian method as
+$\min_{{\mathbf{x}},{\mathbf{z}}_{u},{\mathbf{z}}_{d}}$ $\|{\mathbf{y}}-{\mathbf{H}}{\mathbf{x}}\|_{2}^{2}+\mu_{u}{\mathbf{z}}_{u}^{\top}{\mathbf{L}}^{u}{\mathbf{z}}_{u}+\mu_{d,2}{\mathbf{z}}_{d}^{\top}{\mathcal{L}}^{d}_{r}{\mathbf{z}}_{d}+({\boldsymbol{\gamma}}_{u}^{\tau})^{\top}({\mathbf{x}}-{\mathbf{z}}_{u})+\frac{\rho_{u}}{2}\|{\mathbf{x}}-{\mathbf{z}}_{u}\|_{2}^{2}$
+$+({\boldsymbol{\gamma}}_{d}^{\tau})^{\top}({\mathbf{x}}-{\mathbf{z}}_{d})+\frac{\rho_{d}}{2}\|{\mathbf{x}}-{\mathbf{z}}_{d}\|_{2}^{2}.$ (70)
 We optimize ${\mathbf{x}},{\mathbf{z}}_{u}$ and ${\mathbf{z}}_{d}$ in turn at iteration $\tau$ as follows:
-$\left({\mathbf{H}}^{\top}{\mathbf{H}}+\frac{\rho_{u}+\rho_{d}}{2}{\mathbf{I}}\right){\mathbf{x}}^{\tau+1}=$ ${\mathbf{H}}^{\top}{\mathbf{y}}-\frac{{\bm{\gamma}}_{u}^{\tau}+{\bm{\gamma}}_{d}^{\tau}}{2}+\frac{\rho_{u}}{2}{\mathbf{z}}_{u}^{\tau}+\frac{\rho_{d}}{2}{\mathbf{z}}_{d}^{\tau},$ (69)
-$\left(\mu_{u}{\mathbf{L}}^{u}+\frac{\rho_{u}}{2}{\mathbf{I}}\right){\mathbf{z}}_{u}^{\tau+1}=$ $\frac{{\bm{\gamma}}_{u}^{\tau}}{2}+\frac{\rho_{u}}{2}{\mathbf{x}}^{\tau+1},$ (70)
-$\left(\mu_{d,2}{\mathcal{L}}^{d}_{r}+\frac{\rho_{d}}{2}{\mathbf{I}}\right){\mathbf{z}}_{d}^{\tau+1}=$ $\frac{{\bm{\gamma}}_{d}^{\tau}}{2}+\frac{\rho_{d}}{2}{\mathbf{x}}^{\tau+1}.$ (71)
-The updating of ${\bm{\gamma}}_{u}^{\tau},{\bm{\gamma}}_{d}^{\tau}$ follows eq. 18.
+$\left({\mathbf{H}}^{\top}{\mathbf{H}}+\frac{\rho_{u}+\rho_{d}}{2}{\mathbf{I}}\right){\mathbf{x}}^{\tau+1}=$ ${\mathbf{H}}^{\top}{\mathbf{y}}-\frac{{\boldsymbol{\gamma}}_{u}^{\tau}+{\boldsymbol{\gamma}}_{d}^{\tau}}{2}+\frac{\rho_{u}}{2}{\mathbf{z}}_{u}^{\tau}+\frac{\rho_{d}}{2}{\mathbf{z}}_{d}^{\tau},$ (71)
+$\left(\mu_{u}{\mathbf{L}}^{u}+\frac{\rho_{u}}{2}{\mathbf{I}}\right){\mathbf{z}}_{u}^{\tau+1}=$ $\frac{{\boldsymbol{\gamma}}_{u}^{\tau}}{2}+\frac{\rho_{u}}{2}{\mathbf{x}}^{\tau+1},$ (72)
+$\left(\mu_{d,2}{\mathcal{L}}^{d}_{r}+\frac{\rho_{d}}{2}{\mathbf{I}}\right){\mathbf{z}}_{d}^{\tau+1}=$ $\frac{{\boldsymbol{\gamma}}_{d}^{\tau}}{2}+\frac{\rho_{d}}{2}{\mathbf{x}}^{\tau+1}.$ (73)
+The updating of ${\boldsymbol{\gamma}}_{u}^{\tau},{\boldsymbol{\gamma}}_{d}^{\tau}$ follows Equation 18.
 I.2 Optimization & Algorithm without DGLR Term
-By removing the DGLR term ${\mathbf{x}}{\mathcal{L}}^{d}_{r}{\mathbf{x}}$ in eq.
-4 and introduce only the auxiliary variables $\boldsymbol{\phi}$} in eq.
-5 and ${\mathbf{z}}_{u}$ in eq.
-9, we rewrite the unconstrained version of the optimization of ${\mathbf{x}}^{\tau+1}$ by augmented Lagrangian method as
-$\min_{x,{\mathbf{z}}_{u}}\|{\mathbf{y}}-{\mathbf{H}}{\mathbf{x}}\|_{2}^{2}+\mu_{u}{\mathbf{z}}_{u}^{\top}{\mathbf{L}}^{u}{\mathbf{z}}_{u}+({\bm{\gamma}}^{\tau})^{\top}({\boldsymbol{\phi}}^{\tau}-{\mathbf{L}}^{d}_{r}{\mathbf{x}})+\frac{\rho}{2}\|{\boldsymbol{\phi}}^{\tau}-{\mathbf{L}}^{d}_{r}{\mathbf{x}}\|_{2}^{2}+({\bm{\gamma}}_{u}^{\tau})^{\top}({\mathbf{x}}-{\mathbf{z}}_{u})+\frac{\rho_{u}}{2}\|{\mathbf{x}}-{\mathbf{z}}_{u}\|_{2}^{2}$ (72)
+By removing the DGLR term ${\mathbf{x}}{\mathcal{L}}^{d}_{r}{\mathbf{x}}$ in Equation 4 and introduce only the auxiliary variables $\boldsymbol{\phi}$} in Equation 5 and ${\mathbf{z}}_{u}$ in Equation 9, we rewrite the unconstrained version of the optimization of ${\mathbf{x}}^{\tau+1}$ by augmented Lagrangian method as
+$\min_{x,{\mathbf{z}}_{u}}\|{\mathbf{y}}-{\mathbf{H}}{\mathbf{x}}\|_{2}^{2}+\mu_{u}{\mathbf{z}}_{u}^{\top}{\mathbf{L}}^{u}{\mathbf{z}}_{u}+({\boldsymbol{\gamma}}^{\tau})^{\top}({\boldsymbol{\phi}}^{\tau}-{\mathbf{L}}^{d}_{r}{\mathbf{x}})+\frac{\rho}{2}\|{\boldsymbol{\phi}}^{\tau}-{\mathbf{L}}^{d}_{r}{\mathbf{x}}\|_{2}^{2}+({\boldsymbol{\gamma}}_{u}^{\tau})^{\top}({\mathbf{x}}-{\mathbf{z}}_{u})+\frac{\rho_{u}}{2}\|{\mathbf{x}}-{\mathbf{z}}_{u}\|_{2}^{2}$ (74)
 We optimize $\mathbf{x}$} and ${\mathbf{z}}_{u}$ in turn at iteration $\tau$ as follows:
-$\left({\mathbf{H}}^{\top}{\mathbf{H}}+\frac{\rho}{2}{\mathcal{L}}^{d}_{r}+\frac{\rho_{u}}{2}{\mathbf{I}}\right){\mathbf{x}}^{\tau+1}$ $={\mathbf{H}}^{\top}{\mathbf{y}}+({\mathbf{L}}^{d}_{r})^{\top}(\frac{\gamma^{\tau}}{2}+\frac{\rho}{2}{\boldsymbol{\phi}}^{\tau})-\frac{{\bm{\gamma}}_{u}^{\tau}}{2}+\frac{\rho_{u}}{2}{\mathbf{z}}_{u}^{\tau}$ (73)
-$\left(\mu_{u}{\mathbf{L}}^{u}+\frac{\rho_{u}}{2}{\mathbf{I}}\right){\mathbf{z}}_{u}^{\tau+1}$ $=\frac{{\bm{\gamma}}_{u}^{\tau}}{2}+\frac{\rho_{u}}{2}{\mathbf{x}}^{\tau+1}$ (74)
-The optimization of ${\boldsymbol{\phi}}^{\tau+1}$ is the same as eq. 15, and the solution is the same as eq. 16. The updating of ${\bm{\gamma}},{\bm{\gamma}}_{u}$ follows eq. 17 and eq. 18.
+$\left({\mathbf{H}}^{\top}{\mathbf{H}}+\frac{\rho}{2}{\mathcal{L}}^{d}_{r}+\frac{\rho_{u}}{2}{\mathbf{I}}\right){\mathbf{x}}^{\tau+1}$ $={\mathbf{H}}^{\top}{\mathbf{y}}+({\mathbf{L}}^{d}_{r})^{\top}(\frac{\gamma^{\tau}}{2}+\frac{\rho}{2}{\boldsymbol{\phi}}^{\tau})-\frac{{\boldsymbol{\gamma}}_{u}^{\tau}}{2}+\frac{\rho_{u}}{2}{\mathbf{z}}_{u}^{\tau}$ (75)
+$\left(\mu_{u}{\mathbf{L}}^{u}+\frac{\rho_{u}}{2}{\mathbf{I}}\right){\mathbf{z}}_{u}^{\tau+1}$ $=\frac{{\boldsymbol{\gamma}}_{u}^{\tau}}{2}+\frac{\rho_{u}}{2}{\mathbf{x}}^{\tau+1}$ (76)
+The optimization of ${\boldsymbol{\phi}}^{\tau+1}$ is the same as Equation 15, and the solution is the same as Equation 16.
+The updating of ${\boldsymbol{\gamma}},{\boldsymbol{\gamma}}_{u}$ follows Equation 17 and Equation 18.
 I.3 Algorithm Unrolling with an Undirected Temporal Graph
@@
 By a similar approach which introduces auxiliary variables ${\mathbf{z}}_{u}={\mathbf{x}},{\mathbf{z}}_{n}={\mathbf{x}}$, we optimize ${\mathbf{x}},{\mathbf{z}}_{u},{\mathbf{z}}_{n}$ as follows:
-$\left({\mathbf{H}}^{\top}{\mathbf{H}}+\frac{\rho_{u}+\rho_{n}}{2}{\mathbf{I}}\right){\mathbf{x}}^{\tau+1}=$ ${\mathbf{H}}^{\top}{\mathbf{y}}-\frac{{\bm{\gamma}}_{u}^{\tau}+{\bm{\gamma}}_{n}^{\tau}}{2}+\frac{\rho_{u}}{2}{\mathbf{z}}_{u}^{\tau}+\frac{\rho_{n}}{2}{\mathbf{z}}_{n}^{\tau},$ (76)
-$\left(\mu_{u}{\mathbf{L}}^{u}+\frac{\rho_{u}}{2}{\mathbf{I}}\right){\mathbf{z}}_{u}^{\tau+1}=$ $\frac{{\bm{\gamma}}_{u}^{\tau}}{2}+\frac{\rho_{u}}{2}{\mathbf{x}}^{\tau+1},$ (77)
-$\left(\mu_{n}{\mathbf{L}}^{n}+\frac{\rho_{n}}{2}{\mathbf{I}}\right){\mathbf{z}}_{n}^{\tau+1}=$ $\frac{{\bm{\gamma}}_{n}^{\tau}}{2}+\frac{\rho_{n}}{2}{\mathbf{x}}^{\tau+1}.$ (78)
-The updating of ${\bm{\gamma}}_{u}^{\tau},{\bm{\gamma}}_{n}^{\tau}$ is also similar to eq. 18 as follows:
-${\bm{\gamma}}_{u}^{\tau+1}={\bm{\gamma}}_{u}^{\tau}+\rho_{u}({\mathbf{x}}^{\tau+1}-{\mathbf{z}}_{u}^{\tau+1}),\quad{\bm{\gamma}}_{n}^{\tau+1}={\bm{\gamma}}_{n}^{\tau}+\rho_{n}({\mathbf{x}}^{\tau+1}-{\mathbf{z}}_{n}^{\tau+1}).$ (79)
+$\left({\mathbf{H}}^{\top}{\mathbf{H}}+\frac{\rho_{u}+\rho_{n}}{2}{\mathbf{I}}\right){\mathbf{x}}^{\tau+1}=$ ${\mathbf{H}}^{\top}{\mathbf{y}}-\frac{{\boldsymbol{\gamma}}_{u}^{\tau}+{\boldsymbol{\gamma}}_{n}^{\tau}}{2}+\frac{\rho_{u}}{2}{\mathbf{z}}_{u}^{\tau}+\frac{\rho_{n}}{2}{\mathbf{z}}_{n}^{\tau},$ (78)
+$\left(\mu_{u}{\mathbf{L}}^{u}+\frac{\rho_{u}}{2}{\mathbf{I}}\right){\mathbf{z}}_{u}^{\tau+1}=$ $\frac{{\boldsymbol{\gamma}}_{u}^{\tau}}{2}+\frac{\rho_{u}}{2}{\mathbf{x}}^{\tau+1},$ (79)
+$\left(\mu_{n}{\mathbf{L}}^{n}+\frac{\rho_{n}}{2}{\mathbf{I}}\right){\mathbf{z}}_{n}^{\tau+1}=$ $\frac{{\boldsymbol{\gamma}}_{n}^{\tau}}{2}+\frac{\rho_{n}}{2}{\mathbf{x}}^{\tau+1}.$ (80)
+The updating of ${\boldsymbol{\gamma}}_{u}^{\tau},{\boldsymbol{\gamma}}_{n}^{\tau}$ is also similar to Equation 18 as follows:
+${\boldsymbol{\gamma}}_{u}^{\tau+1}={\boldsymbol{\gamma}}_{u}^{\tau}+\rho_{u}({\mathbf{x}}^{\tau+1}-{\mathbf{z}}_{u}^{\tau+1}),\quad{\boldsymbol{\gamma}}_{n}^{\tau+1}={\boldsymbol{\gamma}}_{n}^{\tau}+\rho_{n}({\mathbf{x}}^{\tau+1}-{\mathbf{z}}_{n}^{\tau+1}).$ (81)
 The undirected temporal graph ${\mathcal{G}}^{n}$ is learned by an UGL described in Section 4.3, and is initialized similar to the spatial UGL described in Appendix F.2.
-The initialization of $\mu_{n},\rho_{n}$ and the CGD parameters in solving eq.
-78 follows the setup in Appendix F.1.
+The initialization of $\mu_{n},\rho_{n}$ and the CGD parameters in solving Equation 80 follows the setup in Appendix F.1.
 Appendix J Error Bar for Our Model in Table 1
@@
 Table 11: $2\sigma$ error bar of the performance metrics of our model in 60-minute traffic forecast in all three datasets.
-Dataset PEMS03 METR-LA
-$2\sigma_{\text{RMSE}}$ 0.05 0.04
-$2\sigma_{\text{MAE}}$ 0.02 0.02
-$2\sigma_{\text{MAPE(\%)}}$ 0.03 0.04
+Dataset PEMS03 METR-LA PEMS08
+$2\sigma_{\text{RMSE}}$ 0.05 0.04 0.04
+$2\sigma_{\text{MAE}}$ 0.02 0.02 0.02
+$2\sigma_{\text{MAPE(\%)}}$ 0.03 0.04 0.04